### v1

# A Comprehensive Graph Pooling Benchmark: Effectiveness, Robustness and Generalizability

Abstract
Graph pooling has gained attention for its ability to obtain effective node and graph representations for various downstream tasks. Despite the recent surge in graph pooling approaches, there is a lack of standardized experimental settings and fair benchmarks to evaluate their performance. To address this issue, we have constructed a comprehensive benchmark that includes 15 graph pooling methods and 21 different graph datasets. This benchmark systematically assesses the performance of graph pooling methods in three dimensions, i.e., effectiveness, robustness, and generalizability. We first evaluate the performance of these graph pooling approaches across different tasks including graph classification, graph regression and node classification. Then, we investigate their performance under potential noise attacks and out-of-distribution shifts in real-world scenarios. We also involve detailed efficiency analysis and parameter analysis. Extensive experiments validate the strong capability and applicability of graph pooling approaches in various scenarios, which can provide valuable insights and guidance for deep geometric learning research. The source code of our benchmark is available at https://github.com/goose315/Graph_Pooling_Benchmark.
1 Introduction
Recently, graph neural networks (GNNs) have garnered significant attention due to their remarkable ability to process graph-structured data across various domains [1, 2] including social networks [3, 4, 5], rumor detection [6, 7], biological networks [8, 9], recommender systems [10] and community detection [11, 12]. Graph pooling approaches play a crucial role in GNNs by enabling the hierarchical reduction of graph representations, which is essential for capturing multi-scale structures and long-range dependencies [13, 14, 15]. They can preserve crucial topological semantics and relationships, which have shown effective for tasks including graph classification, node clustering, and graph generation [13, 16, 17, 2]. In addition, by aggregating nodes and edges, graph pooling can also simplify large-scale graphs, facilitating the application of GNNs in real-world problems [18, 19, 20, 21, 22]. Therefore, understanding and enhancing graph pooling approaches is the key to increasing GNN performance across various domains, driving the progress of deep geometric learning.
In literature, existing graph pooling approaches [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] can be roughly divided into two categories [38, 13], i.e., sparse pooling [23, 24, 25, 26, 27, 28, 29, 30, 31] and dense pooling approaches [32, 33, 34, 35, 36, 37] based on the number of nodes after pooling. Sparse pooling approaches generally keep the number of nodes as the constant cardinality, i.e., $O(1)$ while dense pooling approaches typically have the number of nodes proportional to the number of total nodes after pooling [39, 38]. Even though graph pooling research is becoming increasingly popular, there is still no standardized benchmark that allows for an impartial and consistent comparison of various graph pooling methods. Furthermore, due to the diversity and complexity of graph datasets, numerous experimental settings have been used in previous studies, such as varied proportions of training data and train/validation/test splits [6, 32, 15, 40]. As a result, a comprehensive and publicly available benchmark of graph pooling approaches is highly expected that can facilitate the evaluation and comparison of different approaches, ensuring the reproducibility of results and further advancing the area of graph machine learning.
Towards this end, we present a comprehensive graph pooling benchmark, which includes 15 graph pooling methods and 21 datasets across different graph machine learning problems. In particular, we extensively investigate graph pooling approaches across three key dimensions, i.e., effectiveness, robustness, and generalizability. To begin, we provide a fair and thorough effectiveness comparison of existing graph pooling approaches across graph classification, graph regression and node classification. Then, we evaluate the robustness of graph pooling approaches under both noise attacks on graph structures and node attributes. In addition, we study the generalizability of different approaches under out-of-distribution shifts from both size and density levels. Finally, we include efficiency comparison, parameter analysis and backbone analysis for completeness.
From extensive experiments, we have four observations as follows: (1) Dense pooling approaches generally outperform pooling on graph classification and graph regression. (2) Different graph pooling approaches have a limited impact on node classification. (3) Feature masking influences the performance of graph pooling more seriously compared with structure perturbation and dense pooling is more robust to noise attack. (4) Most of graph pooling approaches suffer from serious performance decrement from distribution shifts.
The main contributions of this paper are as follows:
  • Comprehensive Benchmark. We present the first comprehensive graph pooling benchmark, which incorporates 15 state-of-the-art graph pooling approaches and 21 diverse datasets across graph classification, graph regression and node classification.
  • Extensive Analysis. To investigate the pros and cons of graph pooling approaches, we thoroughly evaluate current approaches from three perspectives, i.e., effectiveness, robustness, and generalizability, which can serve as guidance for researchers in different applications.
  • Open-source Material. We have made our benchmark of all these graph pooling approaches publicly available and reproducible, and believe our benchmark can benefit researchers in both graph machine learning and interdisciplinary fields.
Figure 1: Overview of our hierarchical backbone for graph classification and graph regression.
2 Preliminaries
Notations. Consider a graph $G$ characterized by a vertex set $V$ and an edge set $E$. The features associated with each vertex are represented by the matrix $\bm{X}\in\mathbb{R}^{|V|\times d}$, where $|V|$ denotes the number of vertices, and $d$ signifies the dimensionality of the attribute vectors. The adjacency relationships within the graph are encapsulated by the adjacency matrix $\bm{A}\in\{0,1\}^{|V|\times|V|}$, where an entry $\bm{A}[i,j]=1$ indicates the presence of an edge between vertex $v_{i}$ and vertex $v_{j}$; otherwise, $\bm{A}[i,j]=0$.
Graph Pooling [39, 41, 38]. The aim of graph pooling is to reduce the spatial size of feature maps while preserving essential semantics, which thereby decreases computational complexity and memory usage. In this work, we focus on hierarchical pooling approaches [19]. Let POOL denote a graph pooling function which maps $G$ to a graph $G^{\prime}=(V^{\prime},E^{\prime})$ with the reduced size:
${G}^{\prime}=\texttt{POOL}({G}),$ (1)
where $|V^{\prime}|<|V|$. the process has two main principal components, i.e., reduction, and connection [41]. In particular, the reduction component aims to generate pooled nodes and their attributes in ${G}^{\prime}$ while the connection. The connection component computes the edges $E^{\prime}$ among the $V^{\prime}$ nodes.
Figure 2: Overview of the graph U-Net framework for node classification.
Graph Classification and Regression [23, 42, 41]. The two primary graph-level tasks are graph regression and graph classification. Here, a graph dataset $\mathcal{G}$ is provided as a set of graph-label pairs $(G_{i},y_{i})$, where $y_{i}$ denotes the label for graph $G_{i}$. The objective is to train a powerful discriminative model $f$ that predicts the correct label $y_{i}$ given an input graph $G_{i}$. In graph classification, $y_{i}$ are categorical labels ${1,\cdots,K}$ with $K$ as the number of classes, while in graph regression, $y_{i}$ are continuous values. A well-trained graph classification model should output labels that closely match the true labels, and similarly, a graph regression model should predict values that are nearly identical to the ground truth values. In these tasks, graph pooling always accompanies graph convolutional operators. In formulation, the basic updating rule is written as follows:
$\bm{H}^{(l+1)}=\sigma(\tilde{\bm{D}}^{-\frac{1}{2}}\tilde{\bm{A}}\tilde{D}^{-\frac{1}{2}}\bm{H}^{(l)}\bm{W}^{(l)}),$ (2)
where $\bm{H}^{(l)}$ denotes the node feature matrix at layer $l$, $\bm{W}^{(l)}$ denotes the weight matrix at the crossponding layer, $\tilde{\bm{A}}=\bm{A}+\bm{I}$ is the adjacency matrix $\bm{A}$ plus the identity matrix $\bm{I}$, $\tilde{\bm{D}}$ is the degree matrix of $\tilde{\bm{A}}$, and $\sigma$ is a nonlinear activation function [43]. The pooling layers can be formulated as:
$\bm{H}^{(pool)}=\texttt{POOL}(\bm{H}^{(L)}),$ (3)
where $\bm{H}^{(pool)}$ is the node feature matrix after pooling We iteratively conduct graph convolution and graph pooling operators and adopt a readout function to output the graph representation for downstream tasks. The overview of the basic hierarchical backbone can be found in Figure 1.
Node Classification [43, 44, 45]. The aim of node classification is to assign semantic labels to nodes in a graph according to their attributes and relationships with different nodes. Each dataset involves a graph $G$, consisting of nodes $v_{i}$ and their corresponding labels $y_{i}$. $|V|$ is divided into a labeled set $V^{l}$ and a unlabeled set $V^{u}$. We are required to train a graph neural network model that can predict the missing labels of nodes in $V^{u}$ using the attributes of other nodes. As shown in Figure 2, U-Net framework [46] is widely used to incorporate pooling operations for node classification. In the encoder part, U-Net progressively applies pooling and graph convolution to downsample the graphs and extract multi-scale features. The decoder part of U-Net utilizes upsampling and graph convolution to gradually upsample the low-resolution feature maps back to the original graph size. Residual connections are employed to directly transfer the feature maps from the encoder to the decoder, facilitating the preservation of fine-grained semantics during upsampling [46, 47, 48].
3 Graph Pooling Benchmark
In this section, we present an overview of our graph pooling benchmark, which includes extensive state-of-the-art approaches, graph datasets, and evaluation protocols.
3.1 Graph Pooling Approaches
Our benchmark contains 15 state-of-the-art graph pooling approaches, which can be divided into two categories, i.e., sparse pooling and dense pooling based on the number of nodes after pooling. Table 1 summarizes the experimental details in their corresponding papers.
Sparse Pooling. Sparse pooling approaches typically maintain the number of nodes as a constant cardinality $O(1)$ after graph pooling [41, 38]. These approaches usually utilize a score function to evaluate the importance of each node and keep the nodes with high scores and are easy to implement using sparse matrix calculation. Therefore, they require significantly fewer computational resources, especially when it comes to large-scale graphs. In our benchmark, we select 9 sparse pooling approaches including TopKPool [23], SAGPool [24], ASAPool [25], PANPool [26], COPool [27], CGIPool [28], KMISPool [29], GSAPool [30], and HGPSLPool [31].
Dense Pooling. Dense pooling approaches usually have the number of nodes after pooling proportional to the original node numbers, i.e., $O(|V|)$. These approaches typically adopt graph clustering to learn a new coarsened graph, which can provide a comprehensive semantic view by summarizing similar nodes. One limitation of these approaches is the potential high complexity. In this work, we select 6 dense pooling approaches including AsymCheegerCutPool [32], DiffPool [33], MincutPool [34], DMoNPool [35], HoscPool [36], and JustBalancePool [37].
3.2 Datasets
To systematically evaluate different graph pooling methods, we integrate 21 datasets from different domains for three types of tasks. For graph classification, we select eight publicly available datasets from TUDataset [49], including three molecules datasets, i.e., BZR [50], NCI1 [51], and NCI109 [51], four bioinformatics datasets, i.e., PROTEINS, PROTEINS_full [52], D&D [53], and ENZYMES [54], one social network dataset, i.e., IMDB-MULTI [55], and one synthetic dataset, i.e., COLORS-3 [56]. For graph regression, we choose five datasets from previous works [8] including QM8, BACE, ESOL, FreeSolv, and Lipophilicity. For node classification, we utilize three citation network datasets, i.e., Cora, Citeseer, Pubmed [57] and three website networks, i.e., Cornell, Texas, and Wisconsin [58] and the GitHub dataset [59]. More information can be found in Appendix.
Table 1: Overview of experimental details of graph pooling research. These papers utilize different settings, which validates the necessity of building a comprehensive and fair graph pooling benchmark.
Methods Datasets Tasks
Sparse Pooling
TopKPool NIPS’19 MNIST, COLLAB, PROTEINS, D&D Graph Classification
SAGPool ICML’19 D&D, PROTEINS, NCI1, NCI109, FRANKENSTEIN Graph Classification
ASAPool AAAI’20 D&D, PROTEINS, NCI1, NCI109, FRANKENSTEIN Graph Classification
PANPool NIPS’20 PROTEINS, PROTEINS_FULL, NCI1, AIDS, MUTAGENCITY Graph Classification
COPool ECMLPKDD’22 BZR, AIDS, NCI1, NCI109, PROTEINS, QM7, IMDB-M Graph Classification, Graph Regression
CGIPool SIGIR’22 NCI1, NCI109, Mutagenicity, IMDB-B, IMDB-M, COLLAB PROTEINS Graph Classification
KMISPool AAAI’23 D&D, REDDIT-B, REDDIT-5K, REDDIT-12K, Github Graph Classification, Node Classification
GSAPool WWW’20 D&D, NCI1, NCI109, Mutagencity Graph Classification
HGPSLPool Arxiv’19 D&D, PROTEINS, NCI1, NCI109, ENZYMES, Mutagencity Graph Classification
Dense Pooling
AsymCheegerCutPool ICML’23 Cora, Citeseer, Pubmed, DBLP Node Classification
DiffPool NIPS’18 D&D, PROTEINS, COLLAB, ENZYMES, REDDIT-MULTI Graph Classification
MincutPool ICML’20 D&D, PROTEINS, COLLAB, REDDIT-B, Mutagenicity, QM9 Graph Classification, Graph Regression
DMoNPool JMLR’23 Cora, Citeseer Pubmed, Coauthor Node Classification
HoscPool CIKM’22 Cora, Citeseer Pubmed, Coauthor, DBLP, Email-EU Node Classification
JustBalancePool Arxiv’22 Cora, Citeseer, Pubmed, DBLP Node Classification
3.3 Evaluation Protocols
Our benchmark evaluation encompasses three key aspects, i.e., effectiveness, robustness, and generalizability. Firstly, we conduct a performance comparison of graph pooling approaches across three tasks including graph classification, graph regression, and node classification. For graph and node classification tasks, we employ accuracy as the evaluation metric. For graph regression, we use root mean square error (RMSE) for ESOL, FreeSolv, and Lipophilicity [8]. Following previous research [40], we use the area under the receiver operating characteristic (AUROC) curve to evaluate BACE, and Mean Absolute Error (MAE) for QM8. Secondly, our benchmark evaluates the robustness of graph pooling approaches across two views, i.e., structural robustness and feature robustness [60]. In particular, we add and drop edges of graphs to study structural robustness and mask node features to investigate feature robustness. Thirdly, we employ size-based and density-based distribution shifts to evaluate the generalizability of different pooling methods under real-world scenarios [61]. In addition to these three views, we conduct a further analysis of these graph pooling approaches including the comparison of efficiency, and different backbone parameter choices.
4 Experiment
Table 2: Results of graph classification for different graph pooling methods. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods PROTEINS NCI1 NCI109 ENZYMES BZR IMDB-M COLORS-3 Avg. Rank
Sparse Pooling
TopKPool 72.74$\pm$2.36 69.08$\pm$2.43 67.65$\pm$2.11 38.22$\pm$7.59 83.23$\pm$2.41 41.21$\pm$0.34 43.88$\pm$5.39 62.02 12.50
SAGPool 75.71$\pm$1.90 71.77$\pm$2.31 71.29$\pm$1.98 38.89$\pm$5.35 82.90$\pm$2.19 45.43$\pm$2.08 45.27$\pm$2.89 64.33 8.50
ASAPool 73.45$\pm$2.18 74.81$\pm$1.73 71.97$\pm$0.62 39.33$\pm$8.68 86.45$\pm$2.99 45.72$\pm$1.63 42.79$\pm$2.02 65.29 6.83
PANPool 73.69$\pm$1.38 74.75$\pm$1.45 70.03$\pm$1.01 39.56$\pm$2.86 85.16$\pm$1.21 38.29$\pm$0.31 32.00$\pm$3.68 63.58 8.83
COPool 73.69$\pm$4.54 71.25$\pm$1.24 69.90$\pm$3.45 42.89$\pm$3.49 80.00$\pm$3.47 48.68$\pm$0.75 38.44$\pm$3.56 64.40 8.33
CGIPool 72.26$\pm$1.79 76.24$\pm$2.04 71.81$\pm$1.41 40.00$\pm$5.26 84.84$\pm$2.81 50.36$\pm$0.38 37.78$\pm$5.35 65.92 7.50
KMISPool 72.98$\pm$1.10 79.06$\pm$0.82 72.23$\pm$0.67 20.00$\pm$5.71 87.10$\pm$3.38 40.37$\pm$0.38 20.67$\pm$3.76 61.96 8.17
GSAPool 72.50$\pm$1.90 69.40$\pm$1.55 68.65$\pm$1.81 35.78$\pm$3.74 83.87$\pm$0.00 41.35$\pm$1.28 38.44$\pm$4.19 61.93 12.17
HGPSLPool 73.45$\pm$2.49 75.24$\pm$0.89 68.74$\pm$0.45 43.56$\pm$2.93 88.06$\pm$1.64 39.45$\pm$0.28 41.33$\pm$1.91 64.75 7.83
Dense Pooling
AsymCheegerCutPool 76.59$\pm$1.47 74.75$\pm$0.12 72.13$\pm$0.69 43.78$\pm$5.52 84.52$\pm$2.41 49.12$\pm$0.43 OOT 66.82 5.00
DiffPool 73.05$\pm$1.98 78.35$\pm$0.85 77.00$\pm$1.81 46.00$\pm$3.62 75.16$\pm$1.64 48.48$\pm$0.33 OOT 66.34 6.00
MincutPool 73.41$\pm$2.74 77.60$\pm$1.51 76.58$\pm$1.45 42.89$\pm$2.69 81.61$\pm$3.90 42.44$\pm$0.20 OOT 65.76 6.50
DMoNPool 76.71$\pm$1.18 71.67$\pm$0.58 72.32$\pm$1.48 48.89$\pm$3.78 75.81$\pm$1.77 41.14$\pm$0.34 OOT 64.42 7.50
HoscPool 73.17$\pm$2.00 77.44$\pm$1.79 73.94$\pm$1.53 46.67$\pm$3.78 78.39$\pm$3.16 42.34$\pm$0.58 OOT 65.33 6.67
JustBalancePool 73.17$\pm$2.00 77.63$\pm$1.85 73.94$\pm$1.53 43.78$\pm$3.76 78.39$\pm$3.16 43.08$\pm$0.42 OOT 65.00 6.33
4.1 Experimental Settings
All graph pooling methods in our benchmark are implemented by PyTorch [62]. Graph convolutional networks serve as the default encoders for all algorithms unless otherwise specified. The experimental setup includes a Linux server equipped with NVIDIA L4 and NVIDIA A100 GPUs, as well as an Intel Xeon Gold 6354 CPU. The software stack comprises PyTorch version 1.11.0, PyTorch-geometric version 2.1.0 [63], and Python version 3.9.16.
4.2 Effectiveness Analysis
Performance on Graph Classification. To begin, we investigate the performance of different graph pooling approaches on graph classification. The results of compared approaches on seven popular datasets are recorded in Table 2. From the results, we have the following observations. Firstly, in general, dense pooling methods outperform sparse pooling methods. However, dense pooling methods encounter memory bottlenecks when processing large datasets with tens of thousands of graphs, such as COLORS-3. Secondly, two dense pooling approaches, AsymCheegerCutPool and DiffPool are among the top out of all the compared methods, achieving optimal or near-optimal results on five datasets. The potential reasons include AsymCheegerCutPool minimizing the graph total variation of the cluster assignments, and DiffPool improving the quality of membership assignments by using a link prediction auxiliary objective [33, 32]. Thirdly, among the sparse pooling methods, ASAPool, CGIPool, and HGPSLPool demonstrate the best overall performance. CGIPool achieves the best results on IMDB-M, while HGPSLPool achieves the best results on BZR.
Table 3: Results of graph regression for different pooling methods.
Methods QM8 BACE ESOL FreeSolv Lipophilicity Avg. Rank
Sparse Pooling
TopKPool 0.026 $\pm$ 0.000 0.79 $\pm$ 0.01 1.14 $\pm$ 0.06 1.91 $\pm$ 0.30 0.99 $\pm$ 0.09 1.35 10.67
SAGPool 0.024 $\pm$ 0.000 0.78 $\pm$ 0.01 1.49 $\pm$ 0.28 1.87 $\pm$ 0.06 1.00 $\pm$ 0.09 1.45 12.33
ASAPool 0.025 $\pm$ 0.002 0.77 $\pm$ 0.04 1.06 $\pm$ 0.13 2.52 $\pm$ 0.38 0.95 $\pm$ 0.09 1.51 10.67
PANPool 0.024 $\pm$ 0.000 0.75 $\pm$ 0.02 1.31 $\pm$ 0.05 3.42 $\pm$ 0.32 1.11 $\pm$ 0.02 1.95 14.67
COPool 0.024 $\pm$ 0.002 0.81 $\pm$ 1.56 1.21$\pm$ 0.10 2.53 $\pm$ 0.26 0.94 $\pm$ 0.03 1.56 11.67
CGIPool 0.020 $\pm$ 0.001 0.83 $\pm$ 0.01 1.09 $\pm$ 0.19 1.78 $\pm$ 0.17 0.91 $\pm$ 0.05 1.26 8.00
KMISPool 0.018 $\pm$ 0.001 0.81 $\pm$ 0.01 0.80 $\pm$ 0.03 1.56 $\pm$ 0.65 0.77 $\pm$ 0.01 1.04 5.67
GSAPool 0.025 $\pm$ 0.001 0.81 $\pm$ 0.02 1.22 $\pm$ 0.10 1.95 $\pm$ 0.32 0.99 $\pm$ 0.08 1.39 11.67
HGPSLPool 0.019 $\pm$ 0.000 0.81 $\pm$ 0.02 1.02 $\pm$ 0.14 2.00 $\pm$ 0.31 0.93 $\pm$ 0.03 1.32 9.00
Dense Pooling
AsymCheegerCutPool 0.031 $\pm$ 0.005 0.84 $\pm$ 0.01 0.99 $\pm$ 0.12 2.00 $\pm$ 0.18 0.95 $\pm$ 0.11 1.31 9.33
DiffPool 0.037 $\pm$ 0.010 0.84 $\pm$ 0.02 0.81 $\pm$ 0.05 1.20 $\pm$ 0.09 0.73 $\pm$ 0.03 0.91 4.67
MincutPool 0.020 $\pm$ 0.001 0.85 $\pm$ 0.02 0.76 $\pm$ 0.02 1.19 $\pm$ 0.18 0.73 $\pm$ 0.02 0.89 3.33
DMoNPool 0.021 $\pm$ 0.001 0.85 $\pm$ 0.02 0.68 $\pm$ 0.02 1.16 $\pm$ 0.15 0.69 $\pm$ 0.02 0.84 1.33
HoscPool 0.019 $\pm$ 0.002 0.84 $\pm$ 0.01 0.76 $\pm$ 0.02 1.14 $\pm$ 0.13 0.72 $\pm$ 0.02 0.87 2.33
JustBalancePool 0.022 $\pm$ 0.004 0.85 $\pm$ 0.02 0.74 $\pm$ 0.03 1.26 $\pm$ 0.16 0.70 $\pm$ 0.01 0.90 3.00
Performance on Graph Regression. We further explore the performance of different pooling methods through graph-level regression tasks. As shown in Table 3, we can observe the following: Firstly, dense pooling methods significantly outperform sparse pooling methods. DMoNPool, HoscPool, and JustBalancePool are the best-performing pooling methods, with DMoNPool achieving optimal or near-optimal results on four datasets, and JustBalancePool achieving optimal or near-optimal results on three datasets. Secondly, KMISPool performs well among sparse pooling methods, surpassing AsymCheegerCutPool in dense pooling methods.
Performance on Node Classification. Table 4 presents the performance of various sparse pooling methods in node classification tasks. We can observe that: Firstly, different pooling approaches have consistent performance across these datasets. The potential reason is that graph convolution is the core of semantics exploration on the node level. Secondly, TopKPool demonstrates the best performance among the methods used, potentially because it can adapt to the irregularity of graph data and extract information from the k most important nodes [23]. GSAPool achieves the best or second-best accuracy in 5 out of 7 datasets, as previously mentioned studies have indicated. Before discarding less important nodes, GSAPool aggregates the features of nodes, ensuring that the pooled nodes contain sufficient effective graph information [30]. This aggregation method ensures that the selected nodes not only contain their own features but also integrate information from neighboring nodes. Thirdly, the scalability of ASAPool, PANPool, and HGPSLPool still requires improvement, as they are unable to complete training on larger datasets such as GitHub.
Table 4: Results of node classification for different pooling methods.
Methods Cora Citeseer Pubmed Cornell Texas Wisconsin Github Avg. Rank
TopKPool 89.67$\pm$0.18 77.58$\pm$0.13 87.90$\pm$0.15 79.99$\pm$0.83 82.95$\pm$1.66 82.23$\pm$1.13 87.12$\pm$0.05 82.48 2.80
SAGPool 89.57$\pm$0.18 77.37$\pm$0.12 87.85$\pm$0.16 80.10$\pm$1.58 82.41$\pm$1.06 81.99$\pm$1.07 86.87$\pm$0.04 82.29 4.60
ASAPool 89.70$\pm$0.14 77.57$\pm$0.24 87.59$\pm$0.15 78.89$\pm$1.27 82.52$\pm$0.58 82.55$\pm$0.96 OOM 82.25 3.80
PANPool 89.36$\pm$0.26 77.45$\pm$0.20 87.60$\pm$0.09 79.33$\pm$0.86 81.63$\pm$0.74 81.99$\pm$1.51 OOM 81.95 6.00
COPool 89.29$\pm$0.25 77.36$\pm$0.23 87.52$\pm$0.23 78.25$\pm$0.65 82.97$\pm$1.26 82.23$\pm$0.66 86.25$\pm$0.45 82.02 5.80
CGIPool 89.03$\pm$0.43 77.31$\pm$0.26 87.76$\pm$0.21 79.88$\pm$2.14 81.54$\pm$1.14 81.19$\pm$1.11 84.09$\pm$1.31 81.79 7.80
KMISPool 89.75$\pm$0.17 77.69$\pm$0.27 87.80$\pm$0.06 79.12$\pm$0.91 82.08$\pm$0.98 81.12$\pm$1.19 87.09$\pm$0.04 81.95 4.80
GSAPool 89.61$\pm$0.16 77.61$\pm$0.22 87.89$\pm$0.04 80.01$\pm$0.56 83.50$\pm$1.23 81.75$\pm$0.68 87.19$\pm$0.04 82.50 3.20
HGPSLPool 89.63$\pm$0.24 77.50$\pm$0.18 OOM 79.24$\pm$0.76 82.08$\pm$0.63 81.67$\pm$1.14 OOM 82.02 5.60
Table 5: Results of graph classification under random noise attack for different pooling methods.
Dataset Ptb Method TopKPool SAGPool ASAPool CoPool KMISPool DiffPool MincutPool JustBalancePool
PROTEINS RAW 72.74$\pm$2.36 75.71$\pm$1.90 73.45$\pm$2.18 73.69$\pm$4.54 72.98$\pm$1.10 73.05$\pm$1.98 73.41$\pm$2.74 73.17$\pm$2.00
ADD 66.71$\pm$1.99 70.49$\pm$2.06 64.15$\pm$1.41 63.54$\pm$1.98 59.88$\pm$2.26 71.59$\pm$1.42 69.63$\pm$1.98 70.12$\pm$0.67
DROP 72.50$\pm$2.21 69.09$\pm$1.44 68.17$\pm$1.89 68.41$\pm$2.45 63.72$\pm$4.11 70.37$\pm$3.06 68.84$\pm$4.58 73.05$\pm$1.51
MASK 65.49$\pm$4.24 62.87$\pm$2.89 63.35$\pm$2.89 64.88$\pm$4.71 61.34$\pm$0.56 69.63$\pm$3.91 68.72$\pm$2.33 69.39$\pm$3.39
NCI1 RAW 69.08$\pm$2.43 71.77$\pm$2.31 74.81$\pm$1.73 71.25$\pm$1.24 79.06$\pm$0.82 78.35$\pm$0.85 77.60$\pm$1.51 77.63$\pm$1.85
ADD 62.14$\pm$2.19 65.71$\pm$1.03 63.76$\pm$2.18 63.37$\pm$1.11 62.85$\pm$0.54 70.47$\pm$2.08 66.52$\pm$1.94 70.11$\pm$1.32
DROP 72.51$\pm$0.30 71.11$\pm$0.79 72.99$\pm$0.81 72.46$\pm$0.64 74.66$\pm$0.27 69.66$\pm$0.95 72.65$\pm$0.56 72.63$\pm$0.50
MASK 56.40$\pm$3.65 64.76$\pm$1.61 66.09$\pm$0.76 64.08$\pm$1.37 65.09$\pm$0.91 66.84$\pm$0.95 63.82$\pm$3.06 67.52$\pm$2.48
NCI109 RAW 67.65$\pm$2.11 71.29$\pm$1.9 71.97$\pm$0.62 69.90$\pm$3.45 72.23$\pm$0.67 77.00$\pm$1.81 76.58$\pm$1.45 73.94$\pm$1.53
ADD 66.81$\pm$1.44 56.23$\pm$1.63 60.32$\pm$2.06 63.87$\pm$2.39 62.48$\pm$1.14 64.32$\pm$1.60 65.52$\pm$2.24 69.00$\pm$1.84
DROP 61.00$\pm$1.52 64.39$\pm$1.45 63.39$\pm$1.55 62.06$\pm$1.90 61.29$\pm$1.39 64.26$\pm$2.58 62.42$\pm$1.66 67.48$\pm$1.66
MASK 59.45$\pm$2.88 62.81$\pm$1.61 65.16$\pm$0.94 61.94$\pm$1.27 64.35$\pm$0.72 64.00$\pm$1.74 63.23$\pm$0.95 65.77$\pm$1.96
ENZYMES RAW 38.22$\pm$7.59 38.89$\pm$5.35 39.33$\pm$8.68 42.89$\pm$3.49 20.00$\pm$5.71 46.00$\pm$3.62 42.89$\pm$2.69 43.78$\pm$3.76
ADD 40.44$\pm$2.29 37.56$\pm$0.83 20.67$\pm$1.81 38.89$\pm$2.22 35.56$\pm$4.27 35.33$\pm$5.32 30.67$\pm$3.11 33.33$\pm$3.14
DROP 44.00$\pm$2.29 39.33$\pm$3.11 24.44$\pm$0.60 42.00$\pm$1.47 14.22$\pm$1.78 30.22$\pm$3.81 24.44$\pm$4.61 31.56$\pm$5.56
MASK 35.78$\pm$4.06 33.33$\pm$1.86 16.44$\pm$3.47 30.67$\pm$3.56 16.67$\pm$0.00 23.11$\pm$2.57 22.00$\pm$6.02 23.11$\pm$3.68
BZR RAW 83.23$\pm$2.41 82.90$\pm$2.19 86.45$\pm$2.99 80.00$\pm$3.47 87.10$\pm$3.38 75.16$\pm$1.64 81.61$\pm$3.90 78.39$\pm$3.16
ADD 75.81$\pm$1.77 80.97$\pm$1.21 88.39$\pm$1.88 74.52$\pm$4.38 69.35$\pm$3.38 80.32$\pm$1.58 81.94$\pm$2.37 82.26$\pm$2.04
DROP 76.45$\pm$0.79 71.61$\pm$3.32 74.19$\pm$0.00 74.19$\pm$1.44 79.35$\pm$0.65 72.90$\pm$1.21 76.45$\pm$2.99 76.77$\pm$2.19
MASK 70.00$\pm$2.99 78.39$\pm$3.16 64.52$\pm$8.22 78.06$\pm$3.32 71.94$\pm$0.79 80.32$\pm$2.96 78.71$\pm$4.00 81.29$\pm$3.62
Figure 3: Performance of different approaches w.r.t. different rates of random noise.
4.3 Robustness Analysis
The compared performance for three types of random noise on eight graph pooling methods on the PROTEINS, NCI1, NCI109, ENZYMES, and BZR datasets are shown in Table LABEL:noise. With a probability of 50%, edges of the graph are randomly removed or added, and node features are randomly masked with the same likelihood. From the results, we have the following observations. Firstly, masking node features would degrade the performance the most, which demonstrates that node attributes are still the most crucial for graph representation learning. Secondly, the performance ranking of different methods is consistent under different noise attacks but not consistent across different datasets. This indicates that we can choose the optimal graph pooling approaches by validating the performance under any type of noise attack in practice. Thirdly, overall, dense pooling methods exhibit stronger resistance to random noise attacks compared to sparse pooling methods. For instance, JustBalancePool achieves optimal or near-optimal results on three datasets, and DiffPool also performs exceptionally well on most datasets. Notably, sparse pooling methods significantly outperform dense pooling methods on the multi-classification dataset ENZYMES, as evidenced by the performance of TopKPool and COPool. This suggests that the choice of pooling method can be tailored based on the nature of the classification task.
As depicted in Figure 3, the model’s performance generally declines as the noise intensity increases. It is observed that at the same level of noise, the impact on accuracy is more pronounced on NCI109, while it is relatively minor on ENZYMES. The accuracy degradation due to edge addition noise is less than that caused by node masking noise and edge deletion noise. Among the three types of noise, although the accuracy of nearly all methods decreases amidst fluctuations, JustbalancePool and DiffPool exhibit the strongest robustness. For the NCI1 dataset, MincutPool shows significant variance in accuracy performance.
Table 6: Results of graph classification under distribution shifts. Size and density denote two popular types of shifts across training and test datasets.
Method D&D PROTEINS NCI1 NCI109
Size Density Size Density Size Density Size Density
Sparse Pooling
TopKPool 56.27$\pm$8.82 51.36$\pm$4.27 49.29$\pm$6.74 43.69$\pm$6.26 28.59$\pm$7.73 45.88$\pm$5.69 24.18$\pm$2.05 53.00$\pm$5.00
SAGPool 50.51$\pm$9.52 52.97$\pm$2.67 39.11$\pm$6.90 49.19$\pm$6.30 27.86$\pm$4.15 52.95$\pm$3.06 25.89$\pm$2.91 56.91$\pm$2.59
ASAPool 62.88$\pm$6.87 52.88$\pm$5.99 56.07$\pm$11.66 47.39$\pm$7.12 23.93$\pm$2.22 67.02$\pm$2.25 23.47$\pm$1.48 63.54$\pm$1.70
PANPool 56.1$\pm$5.73 51.86$\pm$5.88 43.39$\pm$14.91 33.69$\pm$15.52 20.24$\pm$0.57 55.46$\pm$2.18 23.56$\pm$2.64 55.18$\pm$2.77
COPool 55.93$\pm$9.94 49.49$\pm$4.88 49.29$\pm$8.21 52.34$\pm$6.70 24.47$\pm$1.34 53.63$\pm$3.40 26.98$\pm$2.92 52.98$\pm$2.62
CGIPool 63.56$\pm$3.83 52.37$\pm$3.81 49.29$\pm$17.31 44.95$\pm$11.85 28.83$\pm$5.88 58.61$\pm$0.88 27.55$\pm$2.45 55.78$\pm$3.12
KMISPool 75.08$\pm$4.37 57.37$\pm$3.68 84.11$\pm$4.70 61.35$\pm$10.65 69.61$\pm$3.09 61.75$\pm$2.76 65.37$\pm$9.00 62.62$\pm$1.67
GSAPool 53.56$\pm$9.34 50.00$\pm$4.30 42.68$\pm$7.54 44.32$\pm$9.66 28.5$\pm$2.51 53.03$\pm$3.76 26.46$\pm$6.55 54.68$\pm$1.57
HGPSLPool 54.41$\pm$5.71 52.97$\pm$2.31 50.71$\pm$6.32 53.42$\pm$3.22 28.01$\pm$0 56.44$\pm$2.16 23.94$\pm$2.18 58.24$\pm$1.39
Dense Pooling
AsymCheegerCutPool OOM OOM 75.71$\pm$3.85 67.21$\pm$6.76 23.83$\pm$5.14 72.18$\pm$0.90 55.25$\pm$24.14 70.35$\pm$0.46
DiffPool 59.83$\pm$7.63 48.77$\pm$14.98 70.00$\pm$5.43 76.04$\pm$5.21 30.73$\pm$6.32 71.35$\pm$1.51 35.82$\pm$21.10 71.43$\pm$1.54
MincutPool 63.39$\pm$10.32 60.17$\pm$22.57 65.00$\pm$9.01 63.15$\pm$8.51 21.21$\pm$2.50 66.12$\pm$4.96 24.70$\pm$3.32 70.04$\pm$1.58
DMoNPool 60.17$\pm$4.12 39.72$\pm$31.15 66.79$\pm$2.22 64.59$\pm$6.78 22.67$\pm$3.09 69.65$\pm$1.72 41.05$\pm$21.66 68.83$\pm$1.49
HoscPool 59.83$\pm$5.53 71.23$\pm$11.4 67.86$\pm$6.63 70.99$\pm$4.19 33.64$\pm$11.34 69.60$\pm$1.77 29.88$\pm$6.60 69.96$\pm$1.75
JustBalancePool 59.83$\pm$5.53 71.23$\pm$11.4 69.64$\pm$3.04 69.10$\pm$3.22 25.39$\pm$3.27 69.60$\pm$1.77 38.48$\pm$19.09 68.63$\pm$1.69
Figure 4: Performance vs Training time and space of different approaches.
4.4 Generalizability Analysis
Table 6 presents the performance of all 15 graph pooling methods under out-of-distribution shifts. For the datasets D&D, PROTEINS, NCI1, and NCI109, two types of distribution shifts are implemented. The first type is based on the number of nodes, where the smallest 50% of graphs by node count are used as the training set, and the largest 20% as the test set, with the remainder serving as the validation set [64, 65]. Following the same criteria, the second type of out-of-distribution shifts are generated based on graph density [65]. For further details, please refer to the Appendix.
From Table 6, we have the following observations. Firstly, the majority of graph pooling approaches suffer from serious performance decreases when it comes to out-of-distribution shifts. For example, most of the approaches have the accuracy below 35% when it comes to size-level distribution shifts. Secondly, for the size-based out-of-distribution shift, KMISPool exhibits superior performance. The underlying reason is its theoretical guarantee of path length distortion boundaries, as well as its ability to preserve key topological features in the coarsened graph. This endows it with enhanced performance when facing size-based out-of-distribution shifts [29]. Thirdly, for the density-based out-of-distribution shift, DiffPool demonstrates optimal performance. The potential reason is its stronger capability to preserve the information of the original graph [33, 41].
4.5 Further Analysis
Efficiency Comparison. In this part, we conduct an efficiency analysis of graph pooling methods on the ENZYMES and NCI1 datasets. We calculate the time of the algorithms by measuring the duration needed to complete 200 epochs of training with the 256 batch size. For space efficiency, we compute the GPU memory utilization during the training process. As shown in Figure 4, dense pooling methods such as DiffPool, MincutPool, and JustBalancePool have significantly higher time and space costs. Notably, although KMISPool performs poorly on the multi-classification dataset ENZYMES, it has one of the lowest time and space costs across all three datasets. Conversely, COPool has the highest GPU memory usage among the eight selected methods.
Backbone Analysis. Figure 5 presents the performance of four pooling methods based on GCNConv [66], TAGConv [67], SAGEConv [68], and GraphConv [69] on four datasets, NCI1, NCI109, PROTEINS, and PROTEINS_full. We observe that TAGConv achieves the best results across all datasets and all four methods except for PROTEINS, while SAGEConv consistently performs the worst on NCI1, NCI109, and PROTEINS_full datasets.
Figure 5: Performance w.r.t. graph convolution backbones for different pooling methods.
Figure 6: Performance w.r.t. different pooling ratios for four pooling methods.
Parameter Analysis. Figure 6 shows the performance of four pooling methods on different datasets. From the results, we observe that as the pooling rate increases from 0.1 to 0.9, the performance would increase before saturation in most datasets. The performance is relatively stable when the ratio is around $0.7$. However, the performance on PROTEINS is limited and keeps declining when the ratio is over $0.3$. The potential reason is the lack of node attributes in PROTEINS, which validates the significance of node attributes in graph classification.
5 Conclusion
In this paper, we construct the first graph pooling benchmark that includes 15 state-of-the-art approaches and 21 different graph datasets across graph classification, graph regression and node classification. This benchmark systematically analyzes the effectiveness, robustness, and generalizability of graph pooling methods. We also make our benchmark publically available to advance the fields of graph machine learning and applications. One limitation of our benchmark is the lack of more complicated settings under label scarcity. In future works, we would extend our graph pooling benchmark to more realistic settings such as semi-supervised learning and few-shot learning.
Appendix A Related Work
A.1 Graph Classification and Graph Regression
The characteristics of a graph are typically described by the properties of its substructures and the interactions between them [70]. Unlike node classification tasks, graph classification or graph regression tasks require attention to the global information of the graph [41, 39]. This necessitates a graph pooling mechanism capable of extracting such global information [38]. Currently, extensive research has been conducted on applying graph pooling techniques to graph classification and regression tasks. These methods can be broadly categorized into two main streams: flat pooling and hierarchical pooling [38, 39].
Flat pooling, as the name suggests, generally involves globally aggregating the features of all nodes in the graph without considering the graph’s structure [20]. It directly combines information from all nodes into a single global vector representation. The advantage of this approach lies in its simplicity and efficiency, which is often sufficiently effective for many application scenarios [39]. Classical flat pooling methods include Set2set, SortPool, DeepSet, and DAGCN.
Set2set’s strength is its use of the chain rule to effectively represent the joint probability of sequences, naturally accommodating variable-sized inputs and/or outputs in sequence form. This method is particularly suitable for scenarios requiring the handling of variable-length inputs [71]. SortPool’s advantage is that it sorts graph vertices in a consistent order, enabling the training of traditional neural networks on graphs. This sorting mechanism allows SortPool to better preserve graph structure information [72]. DeepSet frames the problem within the setting of set learning, encompassing a simple yet powerful readout layer formula. This formula can encode or approximate any continuous permutation-invariant function over a set [73]. DAGCN uses an attention-based graph convolution layer to automatically learn the importance of neighbors at different hops, followed by a self-attention pooling layer that generalizes graph representations from various aspects of the matrix graph embedding, thereby retaining as much of each node’s features and inter-node topological structure as possible [74].
In addition to these classic methods, some more novel flat pooling approaches have been proposed. For instance, GMT is a global pooling layer based on multi-head attention that captures node interactions according to structural dependencies between nodes, while satisfying injectivity and permutation invariance [75]. It also boasts high memory utilization efficiency. QSGCNN extracts multi-scale vertex features from the perspective of quantum information propagation between grid vertices in each graph, integrating graph representation and learning in quantum space graph convolution layers, which do not alter the original spatial positions of vertices [76]. GraphTrans employs Transformer-based self-attention to learn long-range pairwise relationships, effectively addressing optimization instability [77]. DMLAP adapts to both local and global structural information in the graph, featuring an attention pooling layer for each message passing step and computing the final graph representation through unified layer-by-layer graph representationy [78].
Hierarchical pooling, in contrast to flat pooling, incrementally simplifies graph information layer by layer [79, 41]. For a three-layer graph convolutional network (GCN), there are typically two hierarchical pooling layers interleaved between the three convolutional layers. This gradual simplification allows hierarchical pooling to better preserve the hierarchical structure and local information of the graph, thus extracting meaningful global representations at higher levels [79, 38]. This method has significant advantages when dealing with complex graph structures, as it can better capture the intricate relationships between nodes and substructures through multi-level simplification and aggregation [33, 39].
Hierarchical pooling methods for graph classification and regression have been extensively developed, with notable methods including TopKPool, EdgePool, SAGPool, AttPool, ASAPool, PANPool, MVPool, and LiftPool. TopKPool, as one of the most classical hierarchical pooling methods, selects the most important nodes by learning projection scores, thereby retaining crucial information while reducing the graph’s size. This method is applicable to various graph structures, including large-scale and sparse graphs [23]. EdgePool learns local and sparse pooling transformations, which can be integrated into existing architectures without necessitating any changes to the training process [80]. SAGPool is a self-attention-based graph pooling method that leverages the self-attention mechanism of graph convolution, allowing the pooling method to simultaneously consider node features and graph topology [24]. AttPool adaptively selects nodes significant to graph representation and generates hierarchical features by aggregating attention-weighted information from nodes, excelling in learning hierarchical representations of graph embeddings [81]. ASAPool utilizes a novel self-attention network and an improved GNN formula to capture the importance of each node within the given graph. It also learns sparse soft clustering assignments for each layer’s nodes, effectively pooling subgraphs to form a pooled graph [25]. PANPool extends the Laplacian graph to a new transition matrix, which can be customized for different graph data with varying sizes and structures [26]. MVPool is based on a multi-view graph pooling operator that uses attention mechanisms to facilitate cooperation among different perspectives to produce robust node rankings. The pooling operation then adaptively selects a subset of nodes to form an induced subgraph based on the ranking list [82]. LiftPool enhances hierarchical graph representations by maximizing the retention of local structural information within the graph pool. It introduces an additional graph lifting stage before graph coarsening to preserve the local information of the removed nodes and decouples node removal from feature reduction processes [83].
Despite the ongoing innovations in graph pooling for graph classification and regression, several challenges persist. Firstly, the time complexity and space complexity of algorithms remain critical issues [41]. Secondly, most methods are designed for standard graphs, whereas real-world datasets include many other types of graphs with different characteristics and structures [39]. Thirdly, the interpretability of graph pooling requires improvement. Currently, many graph pooling methods cannot effectively separate noise information from the input graph, and their performance lacks robustness when facing attacks on graph topology or features [38].
In conclusion, while hierarchical pooling methods have significantly advanced graph classification and regression tasks, addressing these challenges is crucial for further progress. Improving the efficiency, applicability, and interpretability of these methods will enhance their utility in diverse and complex real-world scenarios.
A.2 Node Classification
Graph node classification aims to classify individual nodes within a graph by predicting their labels based on node attributes and their relationships [84]. For example, in a social network, one might aim to predict each user’s political inclination, while in a protein-protein interaction network, the goal might be to predict each protein’s functional role [85].
Flat pooling methods are less commonly applied to node classification tasks. Among those that do exist, FusionPool is a representative method. FusionPool is designed to address the high parameter count and computational complexity inherent in higher-order graph convolutional networks (GCNs). It achieves this by non-linearly integrating information from both low-order and high-order GCNs [86]. FusionPool combines neighborhood information matrices from different orders, effectively balancing the complexity and richness of the learned node representations.
Hierarchical Pooling Methods for Node Classification Several hierarchical pooling methods have been specifically designed for node classification [41]. These methods aim to capture and preserve the graph’s local and global structural information while simplifying the graph. Representative methods include SEPool, GRAHIES, MVPool, VIPool, TopKPool, DHT, and EdgeCut. SEPool addresses the local structure damage and suboptimal issues inherent in hierarchical pooling methods. It employs the concept of structural entropy and designs a global optimization algorithm that generates a clustering assignment matrix for pooling in a single step, without the need for layer-specific compression ratios [87]. GRAHIES captures the inherent hierarchical topological features of many real-world graphs by merging hierarchical node embeddings. It adaptively learns the multi-level hierarchical structure of the input graph and combines graph representations from different hierarchical levels to capture the intrinsic global hierarchical structure of the original graph [88]. VIPool creates multi-scale graphs in a trainable manner and introduces a novel feature crossing layer that enables cross-scale feature exchange. It selects the most informative subset of vertices based on a neural estimation of mutual information between vertex features and their neighborhoods [89]. VIPool integrates intermediate features between two scales to achieve mutual enhancement. DHT transforms the edges of a graph into nodes of a hypergraph. This dual hypergraph structure allows the application of node representation message-passing techniques to edges. After obtaining edge representations from the hypergraph, it clusters or drops edges to derive an overall graph-level edge representation [90]. EdgeCut generates different versions of the input graph at various scales. It computes edge scores corresponding to the importance of edges during GNN information propagation, addressing the issue of nodes having unordered and irregular neighborhoods [91].
Appendix B Details of Selected Pooling Methods
TopKPool Firstly, a score is computed for each node. These scores can be obtained through a combination of node features and a learnable parameter vector. Based on the computed scores, the top k nodes with the highest scores are selected. Once the top k nodes are determined, the subgraph consisting of these nodes and their associated edges is extracted [23]. This subgraph preserves the local structure of the original graph while significantly reducing the number of nodes. The primary purpose of this step is to simplify the graph structure, making subsequent computations more efficient, and to embed the subgraph into a lower-dimensional space [23].
SAGPool Firstly, a self-attention mechanism is employed to assess the importance of nodes. Based on the importance scores of the nodes, a subset of nodes is selected for pooling. Typically, the top k nodes with the highest importance scores are chosen, and the selected nodes along with their corresponding edges form a new subgraph [24]. This subgraph preserves the critical structural information of the original graph while reducing its size. Finally, feature aggregation is performed on the selected nodes to obtain the final output [24].
ASAPool The ASAPool aims to enhance the aggregation effect of feature maps by introducing a spatial attention mechanism. This selectively focuses on important features, thereby strengthening the model’s expressive power. For the input features, convolutional operations are applied to the feature maps to extract local features [25]. The convolutional kernel size is typically the same as the pooling window size. Then, a lightweight attention module is used to compute the attention weights for each spatial position. This can be implemented using convolutional layers or fully-connected layers to generate the weight matrix [25]. Finally, the attention weight matrix is applied to the input feature map, yielding the weighted feature map [25].
PANPool includes a convolution operation that involves paths linking each message sender and receiver, with these paths having learnable weights depending on their length, corresponding to a maximum entropy random walk. It extends the graph Laplacian to what is referred to as the Maximum Entropy Transition (MET) matrix, whose diagonal entries are directly related to subgraph centrality, thereby providing a natural adaptive pooling mechanism [26].
COPool combines pooled representations learned from both node and edge views. Through cross-view interactions, the edge view pooling and node view pooling mutually enhance each other to learn more informative graph-level representations [27].
CGIPool combines pooled representations learned from both node and edge views. Through cross-view interactions, the edge view pooling and node view pooling mutually enhance each other to learn more informative graph-level representations [28].
KMISPool enhances the reduction sampling mechanism by applying downsampling to graph data. This graph pooling method corresponds to a controllable isometric coarsening mechanism in regular data structures [29].
GSAPool simultaneously considers both the structural and feature information of the graph. It aggregates node feature information before discarding unimportant nodes; thus, the selected nodes contain information from neighboring nodes, which can enhance the utilization of features from the unselected nodes [30].
HGPSLPool integrates graph pooling and structure learning into a unified module to generate hierarchical representations of the graph. Through the graph pooling operation, it adaptively selects a subset of nodes to form the induced subgraph for subsequent layers [31].
AsymCheegerCutPool computes cluster assignments by optimizing a tighter relaxation of the minimum cut based on Graph Total Variation (GTV). These cluster assignments can be directly used for vertex clustering or to implement graph pooling [32].
DiffPool is a differentiable graph pooling module capable of generating hierarchical representations of a graph. It can be integrated end-to-end with various graph neural network architectures. DiffPool learns differentiable soft cluster assignments for nodes at each layer of a deep GNN, mapping nodes to a set of clusters, which then form the coarsened input for the next GNN layer [33].
MincutPool formulates a continuous relaxation of the normalized minCUT problem and trains a GNN to compute cluster assignments that minimize this objective. It does not require spectral decomposition and learns a clustering function that can be quickly evaluated on out-of-sample graphs [34].
DMoNPool is an unsupervised pooling method inspired by the modularity measure of clustering quality, which is typically used to recover cluster structures [35].
HoscPool is a cluster-based graph pooling operator that hierarchically captures higher-order information, resulting in richer graph representations. It end-to-end learns a probabilistic clustering assignment matrix by minimizing a relaxed formulation of motif spectral clustering in the objective function, which is then extended to the pooling operator [36].
JustBalancePool optimizes an unsupervised loss composed of two terms. The first term ensures that connected nodes are assigned to the same cluster, while the second is a balancing term that prevents degenerate solutions by encouraging samples to be assigned to only one cluster and clusters to have similar sizes [37].
Appendix C Additional Experimental Details
C.1 Graph Classification
The graph classification model utilized in this study shares an identical structure across 15 different pooling methods. Primarily, the model comprises three SAGEConv layers with ReLU activation functions and two pooling layers, followed by a global average pooling layer. Both the hidden and output channels are set to 64. Additionally, the embedding output from the global average pooling layer is processed through a linear layer with ReLU activation and dimensions (64, 32), and subsequently through another linear layer without any activation function and dimensions (32, number of classes). The final output is obtained by applying a softmax function to the embedding output. Due to constraints in computational resources and time, hyperparameter tuning was not performed. All models employed the Adam optimizer with a learning rate set to 0.001 and were trained for 200 epochs using the negative log-likelihood loss function. The data was uniformly split into training, validation, and test sets with ratios of 70%, 15%, and 15%, respectively. Each experiment was run five times with different random seeds.
C.2 Graph Regression
For graph regression, we adopted the backbone network inspired by MESPool [92]. Primarily, the model comprises three GINConv layers with ReLU activation functions and BatchNorm, along with two pooling layers, followed by a global average pooling layer. Both the hidden and output channels are set to 64. Additionally, the embedding output from the global average pooling layer is processed through a linear layer with ReLU activation and dimensions (64, 32). Due to constraints in computational resources and time, hyperparameter tuning was not performed. All models employed the Adam optimizer with a learning rate set to 0.001 and were trained for 200 epochs using the negative log-likelihood loss function. The data was uniformly split into training, validation, and test sets with ratios of 80%, 10%, and 10%, respectively. Each experiment was run five times with different random seeds.
C.3 Node Classification
For node classification, we utilize a U-net architecture, which we divide into a downsampling convolutional part and an upsampling convolutional part [93]. The downsampling convolutional section includes two SAGEConv layers with ReLU activation functions, with pooling applied between these layers. In the upsampling convolutional section, we use the indices saved during pooling for upsampling, restoring features to their pre-pooling size. The upsampled features are then fused with the corresponding residual features from the downsampling path, either through summation or concatenation. Finally, the fused features are processed and activated through a SAGEConv layer.
Due to constraints in computational resources and time, we do not perform hyperparameter tuning. All models employ the Adam optimizer with a learning rate set to 0.001 and are trained for 200 epochs using cross-entropy loss. All data are processed using a 5-fold cross-validation and are run on five different seeds.
Appendix D Detailed Description of Datasets
D.1 Graph Classification
Table A.1 provides descriptive statistics of the selected datasets, revealing that our chosen datasets encompass graph data of varying scales and features. This diversity establishes a robust foundation for benchmarking. The following are detailed descriptions of these datasets:
PROTEINS is a collection of graphs representing protein structures where nodes denote secondary structure elements (SSEs) and edges indicate neighborhood relationships between these SSEs. The primary aim of this dataset is to facilitate the classification of proteins into different structural classes based on their amino acid sequences and structural characteristics. Each graph in the dataset is labeled according to its protein class, and the dataset encompasses a diverse range of protein structures, making it a valuable resource for studying the application of graph-based learning methods in bioinformatics and structural biology [52].
PROTEINS_full is an extended version of the PROTEINS dataset, consisting of protein structure graphs for the task of graph classification. Each graph represents a protein, where nodes correspond to secondary structure elements (SSEs) such as alpha helices and beta sheets, and edges represent neighborhood relationships between these SSEs based on spatial proximity or sequential adjacency [52].
NCI1 is a collection of chemical compound graphs derived from the National Cancer Institute (NCI) database, specifically used for the task of graph classification. Each graph in the dataset represents a chemical compound, where nodes correspond to atoms and edges represent the bonds between them. The primary objective of this dataset is to facilitate the classification of compounds based on their ability to inhibit or interact with cancer cell lines, thus aiding in drug discovery and cancer research. The dataset is labeled with two classes, indicating the compounds’ biological activity against a specific cancer cell line [51].
NCI109 is a collection of chemical compound graphs, derived from the National Cancer Institute (NCI) database, aimed at graph classification tasks. Each graph in the dataset represents a chemical compound, with nodes corresponding to atoms and edges representing the bonds between them. The dataset is specifically labeled to reflect the compounds’ bioactivity against cancer cell lines, similar to its counterpart NCI1. NCI109 includes two classes, distinguishing between compounds based on their ability to inhibit or interact with a particular cancer cell line [51].
ENZYMES is a collection of protein graphs specifically designed for the classification of enzymes into one of six EC (Enzyme Commission) top-level classes. Each graph represents a protein, where nodes denote secondary structure elements (SSEs) and edges indicate spatial or sequential neighborhood relationships between these SSEs. The dataset includes attributes for each node, reflecting the physicochemical properties of the amino acids they represent [54].
BZR is a collection of molecular graphs representing chemical compounds, specifically designed for the task of graph classification. Each graph in the dataset corresponds to a chemical compound where nodes represent atoms and edges denote the chemical bonds between them. The BZR dataset is focused on classifying compounds based on their biological activity related to binding with benzodiazepine receptors, which are significant in pharmacology for their role in the effects of various drugs on the central nervous system. Each graph is labeled to indicate whether the compound binds to the benzodiazepine receptor, making this dataset valuable for studying the interaction between chemical compounds and biological targets [50].
IMDB-M is a collection of social network graphs derived from the Internet Movie Database (IMDB). Each graph represents a collaboration network from movies, where nodes correspond to actors or actresses, and edges indicate that the two actors appeared in the same movie. The dataset is specifically designed for the task of multi-class graph classification, with labels indicating the genre of the movie (e.g., Action, Comedy, Drama). This dataset comprises three classes, reflecting the movie genres, and aims to facilitate the classification of social network structures based on their topological features [55].
COLORS-3 is a collection of artificially generated graphs designed for the task of graph classification. Each graph in the dataset represents a structure where nodes are assigned one of three possible colors. The dataset’s primary challenge is to classify these graphs based on their color distribution patterns and structural properties. COLORS-3 provides a controlled environment to test and benchmark graph-based learning algorithms, particularly focusing on the ability to distinguish graphs with different color configurations [56].
D&D is a collection of protein structure graphs designed for the task of graph classification. Each graph in this dataset represents a protein, with nodes corresponding to amino acids and edges representing the spatial or sequential proximity between these amino acids. The primary objective of the D&D dataset is to classify proteins into one of two categories: enzymes or non-enzymes [53].
Table A.1: Summary statistics of datasets for graph classification.
Datasets Graphs Classes Avg. Nodes Avg. Edges Node Attr.
PROTEINS 1,113 2 39.06 72.82 + (1)
PROTEINS_full 1,113 2 39.06 72.82 + (29)
NCI1 4,110 2 29.87 32.30 -
NCI109 4,127 2 29.68 32.13 -
ENZYMES 600 6 32.63 62.14 + (18)
BZR 405 2 35.75 38.36 + (3)
IMDB-M 1,500 3 13.00 65.94 -
COLORS-3 10,500 11 61.31 91.03 + (4)
D&D 1,178 2 284.32 715.66 -
D.2 Graph Regression
Table A.2 provides an overview of the selected datasets in terms of their tasks, compounds and their features, recommended splits, and metrics. A more detailed description is provided below.
QM8 is a benchmark dataset in computational chemistry, designed to facilitate the development and evaluation of machine learning models for quantum mechanical property prediction. It contains approximately 21,786 molecular structures, each characterized by their calculated properties using quantum chemistry methods, specifically focusing on electronic spectra [8, 94].
BACE is a collection of biochemical data used to evaluate computational methods for drug discovery. The dataset includes a total of 1,522 compounds, each annotated with their binding affinities, as well as molecular descriptors and fingerprints to facilitate the development and assessment of machine learning modelsa [8, 95].
ESOL is a prominent resource in cheminformatics, designed for evaluating machine learning models on the prediction of aqueous solubility of small molecules. The dataset, derived from the work of Delaney, encompasses a diverse range of chemical compounds with experimentally determined solubility values expressed in logS, where S is the solubility in mols per liter. It includes 1128 compounds, serving as a benchmark for solubility prediction tasks [96, 8].
FreeSolv is a collection of experimental and calculated hydration free energies for small molecules in aqueous solution. It comprises data for a wide range of organic molecules, providing both experimental values and calculated predictions based on molecular dynamics simulations [97, 8].
Lipophilicity is primarily utilized for studying and evaluating molecular lipophilicity. This dataset comprises 4,200 compounds sourced from the ChEMBL database, with experimentally measured partition coefficient (logD) values that reflect the distribution behavior of compounds in a water-octanol system [98, 8].
Table A.2: Details of datasets for graph regression.
Datasets Data type Tasks Compounds Split Metrics
QM8 SMILES, 3D coordinates 12 21,786 Random MAE
BACE SMILES 1 1,522 Scaffold AUROC
ESOL SMILES 1 1,128 Random RMSE
FreeSolv SMILES 1 643 Random RMSE
Lipophilicity SMILES 1 4,200 Random RMSE
D.3 Node Classification
Table A.3 presents descriptive statistics of the seven datasets used for node classification. It is evident that there is a significant variance in the scale of the selected datasets, each possessing distinct characteristics. Further background information and details about these datasets are provided below.
Cora is a benchmark commonly utilized in the evaluation of graph-based machine learning algorithms, comprises a collection of 2,708 scientific publications classified into seven distinct categories. These categories encompass a wide range of research topics in the field of machine learning. Each publication in the dataset is represented as a node in a citation network, where edges indicate citation relationships between papers [57].
CiteSeer is a widely used citation network dataset. It comprises scientific publications categorized into six classes, with each publication represented by a 3,327-dimensional binary vector indicating the presence or absence of specific words. Additionally, the dataset includes 4,552 citation links among these publications, forming a directed graph where each node represents a paper, and each edge denotes a citation from one paper to another [57].
PubMed consists of scientific publications from the PubMed database, categorized into three classes based on their Medical Subject Headings (MeSH) terms. Each publication is represented as a node, and the citation relationships between the publications form the edges of the graph. Additionally, each node is characterized by a sparse bag-of-words feature vector derived from the content of the corresponding publication [57].
Table A.3: Summary statistics of datasets for node classification.
Datasets Number of Nodes Number of Edges Number of Features Number of Classes
Cora 2,708 10,556 1,433 7
CiteSeer 3,327 9,104 3,703 6
PubMed 19,717 88,648 500 3
Cornell 183 298 1,703 5
Texas 183 325 1,703 5
Wisconsin 251 515 1,703 5
Github 37,700 578,006 0 2
Cornell, Texas, and Wisconsin originate from the citation networks of their respective universities and consist of nodes representing web pages and edges representing hyperlinks between these pages. Each node is labeled with a class, typically corresponding to the topic of the web page, allowing for tasks such as node classification and link prediction. The datasets vary in size, with Cornell, Texas, and Wisconsin having 183, 183, and 251 nodes respectively, and are characterized by relatively small and sparse networks compared to larger citation datasets [58].
Github includes rich node attributes representing the diverse features of developers, such as their interests, skills, and contributions to various repositories. Additionally, the edges within the network capture the interactions and collaborations among developers, creating a multi-faceted graph structure [59].
D.4 Out-of-distribution shifts
Size shifts For the selected datasets, including D&D, PROTEINS, NCI1, and NCI109, we utilized the data provided by the authors of size-invariant-GNNs [64]. In this setup, the graphs with the smallest 50% of nodes are used as the training set, those with the largest 20% of nodes are used as the test set, and the remaining graphs were used as the validation set. The data can be downloaded from https://www.dropbox.com/s/38eg3twe4dd1hbt/data.zip.
Density shifts For the selected datasets, we divide the datasets based on graph density: the 50% of graphs with the lowest density are used as the training set, the 20% with the highest density are used as the test set, and the remaining graphs are used as the validation set. After applying density shifts, the following densities are observed: for D&D, the training set density is 0.0274, the validation set density is 0.0567, and the test set density is 0.1245; for PROTEINS, the training set density is 0.1709, the validation set density is 0.4183, and the test set density is 1.0752; for NCI1, the training set density is 0.1229, the validation set density is 0.1976, and the test set density is 0.2922; for NCI109, the training set density is 0.1248, the validation set density is 0.2000, and the test set density is 0.2916.
Appendix E Additional Experiments
Table A.4 presents the results of fixing the backbone as GraphConv and setting the pooling ratios of the sparse pooling method to 0.1, 0.3, 0.5, 0.7, and 0.9, respectively. The pooling ratios for dense pooling remain unchanged. As shown in Table A.4, with the increase in pooling ratio, accuracy exhibits a fluctuating upward trend. Among the sparse pooling methods, HGPSLPool and KMISPool consistently demonstrate superior performance. However, changes in pooling ratio also result in varying stability across different methods. Furthermore, compared to binary classification tasks, the accuracy improvement with increasing pooling ratio is more pronounced in multi-class classification tasks. This suggests that researchers should consider using larger pooling ratios when conducting studies on multi-class classification.
Table A.5 presents the results of using GCNConv as the backbone model while varying the pooling ratio. It can be observed that the overall accuracy of GCNConv is lower than that of GraphConv. Generally, as the pooling ratio increases, classification accuracy improves, with dense pooling methods proving to be more powerful than sparse pooling methods, albeit at the cost of increased computation time. In GCNConv, among the sparse pooling methods, CGIPool and KMISPool show the best performance. Additionally, the impact of pooling ratio on multi-class classification accuracy remains significantly greater than its impact on binary classification accuracy.
Table A.6 presents the results of varying the pooling ratio (0.1, 0.3, 0.5, 0.7, 0.9) in node classification tasks using GraphConv as the backbone model. Overall, the change in pooling ratio has a minimal impact on the accuracy of node classification tasks. For most methods, accuracy does not improve with an increased pooling ratio but rather fluctuates. The two best-performing methods overall are TopKPool and GSAPool. Additionally, higher pooling ratios result in increased computational costs for some methods, such as HGPSLPool and ASAPool. This suggests that when performing node classification, graph pooling methods should be used more cautiously, and excessive concern over pooling ratios may be unnecessary.
Table A.7 presents the results of node classification with GCNConv as the backbone model, varying the pooling ratio from 0.1, 0.3, 0.5, 0.7, to 0.9. It can be observed that, except for HGPSLPool, the performance of most methods with GCNConv is inferior to that with GraphConv. The performance of different pooling methods varies across different pooling ratios, making it difficult to identify a single pooling method that consistently outperforms others at all pooling ratios. Nevertheless, CGIPool and KMISPool still relatively demonstrate the best overall performance. Additionally, increasing the pooling ratio does not consistently improve node classification tasks. Lastly, GCNConv has higher memory consumption compared to GraphConv, particularly for ASAPool, PANPool, COPool, and HGPSLPool.
In summary, based on the aforementioned experiments, GraphConv is recommended as the backbone model over GCNConv for both graph classification and node classification tasks due to its higher accuracy and lower computational resource consumption. For graph classification tasks, selecting a larger pooling ratio generally improves results, especially for multi-class classification tasks. In contrast, for node classification tasks, the choice of pooling ratio is not very important.
Table A.4: Results of graph classification using GraphConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods PROTEINS NCI1 NCI109 ENZYMES BZR IMDB-M COLORS-3
Sparse Pooling, pooling ratio = 0.1
TopKPool 70.48$\pm$3.99 73.44$\pm$1.69 58.26$\pm$2.22 31.11$\pm$5.26 81.94$\pm$1.58 40.71$\pm$3.21 22.59$\pm$7.61
SAGPool 71.31$\pm$3.40 73.00$\pm$1.20 64.48$\pm$2.16 26.44$\pm$4.63 84.52$\pm$3.16 37.96$\pm$2.98 29.17$\pm$2.48
ASAPool 70.36$\pm$4.66 76.47$\pm$2.13 68.19$\pm$2.97 21.11$\pm$6.67 85.48$\pm$2.70 41.96$\pm$2.48 40.23$\pm$1.90
PANPool 68.33$\pm$1.09 77.86$\pm$0.92 65.00$\pm$1.22 30.89$\pm$4.73 86.45$\pm$3.32 39.20$\pm$2.34 21.57$\pm$0.75
COPool 70.36$\pm$1.48 75.59$\pm$0.96 66.97$\pm$2.26 39.11$\pm$4.63 81.61$\pm$1.29 35.11$\pm$2.55 28.25$\pm$2.91
CGIPool 67.86$\pm$3.63 76.21$\pm$1.71 64.84$\pm$2.83 40.00$\pm$5.26 84.52$\pm$2.99 44.98$\pm$2.49 28.27$\pm$7.25
KMISPool 75.60$\pm$1.13 76.34$\pm$0.77 73.29$\pm$1.31 35.11$\pm$9.34 85.48$\pm$5.40 38.76$\pm$1.76 57.35$\pm$2.29
GSAPool 71.67$\pm$3.03 74.42$\pm$2.36 66.19$\pm$3.58 27.56$\pm$1.91 83.87$\pm$1.02 40.71$\pm$1.72 28.32$\pm$6.78
HGPSLPool 75.95$\pm$1.10 77.21$\pm$0.85 68.03$\pm$1.82 32.00$\pm$3.74 86.77$\pm$2.77 33.69$\pm$0.81 34.80$\pm$1.43
Sparse Pooling, pooling ratio = 0.3
TopKPool 70.48$\pm$3.99 70.99$\pm$0.78 65.55$\pm$0.80 28.22$\pm$2.86 83.23$\pm$1.64 41.24$\pm$3.95 28.76$\pm$5.69
SAGPool 70.83$\pm$3.30 72.03$\pm$1.95 70.26$\pm$2.07 27.33$\pm$4.01 83.55$\pm$2.14 41.60$\pm$2.15 35.58$\pm$3.54
ASAPool 74.40$\pm$3.43 76.82$\pm$0.41 72.52$\pm$1.40 26.00$\pm$9.10 82.58$\pm$2.37 37.96$\pm$4.56 50.46$\pm$1.99
PANPool 68.33$\pm$1.52 77.70$\pm$1.92 68.23$\pm$1.79 32.44$\pm$3.25 82.90$\pm$2.99 40.89$\pm$2.71 32.55$\pm$0.99
COPool 70.71$\pm$0.69 74.26$\pm$2.05 71.16$\pm$2.07 36.22$\pm$8.96 83.23$\pm$2.62 39.02$\pm$3.40 33.16$\pm$2.87
CGIPool 68.69$\pm$4.27 76.11$\pm$1.53 68.55$\pm$2.33 29.56$\pm$6.31 80.00$\pm$2.19 38.04$\pm$3.27 39.25$\pm$5.92
KMISPool 77.02$\pm$1.53 76.95$\pm$0.65 73.48$\pm$1.42 32.22$\pm$8.40 84.84$\pm$4.96 39.02$\pm$2.49 57.35$\pm$2.29
GSAPool 70.71$\pm$2.80 74.23$\pm$1.61 70.65$\pm$1.03 28.22$\pm$7.88 84.19$\pm$2.58 38.76$\pm$2.27 35.18$\pm$4.46
HGPSLPool 75.36$\pm$1.23 77.99$\pm$1.32 71.61$\pm$1.00 30.00$\pm$7.34 84.84$\pm$3.62 38.76$\pm$2.17 50.32$\pm$1.38
Sparse Pooling, pooling ratio = 0.5
TopKPool 73.10$\pm$1.09 70.57$\pm$1.56 69.68$\pm$3.15 39.33$\pm$5.10 82.26$\pm$4.45 39.82$\pm$1.11 35.06$\pm$5.96
SAGPool 72.62$\pm$3.22 71.96$\pm$1.88 71.94$\pm$2.19 33.33$\pm$1.99 83.55$\pm$2.37 39.64$\pm$2.15 41.83$\pm$4.26
ASAPool 71.90$\pm$5.10 76.11$\pm$1.35 72.52$\pm$1.40 22.67$\pm$9.85 83.55$\pm$3.13 35.38$\pm$5.10 59.79$\pm$2.37
PANPool 73.81$\pm$1.88 77.83$\pm$1.61 70.35$\pm$1.82 41.11$\pm$4.61 86.13$\pm$2.19 39.02$\pm$3.57 41.29$\pm$1.03
COPool 73.21$\pm$1.92 74.75$\pm$2.40 72.87$\pm$2.17 38.22$\pm$2.06 82.90$\pm$1.29 39.29$\pm$2.36 38.43$\pm$2.52
CGIPool 74.29$\pm$1.15 76.73$\pm$2.37 71.90$\pm$0.89 28.67$\pm$5.51 84.84$\pm$5.26 37.16$\pm$3.33 41.32$\pm$4.06
KMISPool 75.95$\pm$0.97 76.30$\pm$1.06 74.03$\pm$0.54 33.11$\pm$8.36 84.84$\pm$3.47 38.76$\pm$2.32 57.37$\pm$1.94
GSAPool 74.05$\pm$2.98 74.04$\pm$1.98 72.16$\pm$2.73 33.11$\pm$2.57 82.90$\pm$2.62 41.33$\pm$0.80 38.22$\pm$6.40
HGPSLPool 74.64$\pm$1.58 77.41$\pm$1.39 73.61$\pm$1.08 29.11$\pm$3.94 85.16$\pm$2.58 41.60$\pm$1.70 62.30$\pm$2.55
Sparse Pooling, pooling ratio = 0.7
TopKPool 73.81$\pm$1.25 74.88$\pm$1.04 72.94$\pm$1.04 37.78$\pm$4.61 81.29$\pm$3.16 40.44$\pm$2.77 42.77$\pm$4.11
SAGPool 72.86$\pm$1.90 74.55$\pm$1.41 73.19$\pm$2.33 41.11$\pm$5.07 84.84$\pm$3.32 39.82$\pm$2.97 41.83$\pm$4.26
ASAPool 74.05$\pm$2.25 79.06$\pm$0.68 74.87$\pm$1.26 18.00$\pm$1.30 85.48$\pm$3.68 34.40$\pm$4.44 59.79$\pm$2.37
PANPool 74.64$\pm$1.44 77.05$\pm$0.39 74.35$\pm$0.87 42.00$\pm$6.02 85.16$\pm$3.13 41.33$\pm$1.76 41.29$\pm$1.03
COPool 74.17$\pm$1.23 75.20$\pm$0.90 73.97$\pm$2.58 39.56$\pm$5.60 81.29$\pm$2.19 39.64$\pm$4.14 38.43$\pm$2.52
CGIPool 75.83$\pm$2.51 78.38$\pm$1.70 74.71$\pm$0.97 28.67$\pm$5.51 84.84$\pm$5.26 37.16$\pm$3.33 41.32$\pm$4.06
KMISPool 76.67$\pm$1.02 76.66$\pm$0.42 73.06$\pm$1.45 34.00$\pm$6.76 85.16$\pm$2.37 37.78$\pm$5.36 41.32$\pm$4.06
GSAPool 72.14$\pm$1.52 75.49$\pm$1.14 72.77$\pm$1.39 38.00$\pm$5.77 83.55$\pm$2.37 41.87$\pm$3.17 38.22$\pm$6.40
HGPSLPool 75.00$\pm$2.23 78.28$\pm$1.00 73.84$\pm$0.55 24.67$\pm$4.52 86.45$\pm$1.94 41.33$\pm$0.97 62.30$\pm$2.55
Sparse Pooling, pooling ratio = 0.9
TopKPool 73.21$\pm$1.92 75.82$\pm$1.45 73.19$\pm$1.73 39.56$\pm$3.27 82.58$\pm$3.13 41.16$\pm$1.58 47.87$\pm$2.36
SAGPool 72.74$\pm$0.95 76.21$\pm$1.96 74.45$\pm$0.56 39.33$\pm$2.59 81.61$\pm$2.19 41.24$\pm$3.48 49.66$\pm$2.52
ASAPool 73.69$\pm$1.48 79.09$\pm$0.45 77.16$\pm$1.42 19.33$\pm$1.51 86.77$\pm$1.88 34.84$\pm$4.11 72.74$\pm$1.21
PANPool 74.88$\pm$1.02 77.99$\pm$0.72 76.06$\pm$1.05 36.44$\pm$4.73 83.55$\pm$1.21 39.11$\pm$1.93 64.61$\pm$2.22
COPool 74.52$\pm$1.38 77.08$\pm$2.42 75.03$\pm$1.02 36.44$\pm$2.85 81.94$\pm$1.21 35.73$\pm$2.31 43.67$\pm$1.68
CGIPool 72.62$\pm$1.92 79.29$\pm$1.19 75.29$\pm$0.98 31.33$\pm$4.89 82.90$\pm$0.79 34.84$\pm$3.38 44.35$\pm$1.25
KMISPool 74.29$\pm$3.14 75.85$\pm$0.93 74.03$\pm$1.30 31.11$\pm$8.66 84.84$\pm$4.85 38.84$\pm$1.91 57.35$\pm$2.29
GSAPool 72.38$\pm$1.83 76.79$\pm$1.70 73.42$\pm$0.76 33.33$\pm$2.33 83.23$\pm$2.99 40.62$\pm$3.26 52.43$\pm$2.50
HGPSLPool 75.60$\pm$1.46 77.21$\pm$1.53 74.84$\pm$0.58 27.11$\pm$5.10 86.45$\pm$3.62 38.22$\pm$3.85 87.12$\pm$1.82
Dense Pooling
AsymCheegerCutPool 75.12$\pm$1.05 73.87$\pm$0.89 72.74$\pm$1.03 31.78$\pm$3.19 83.55$\pm$2.77 43.02$\pm$3.46 OOT
DiffPool 75.00$\pm$0.77 78.38$\pm$0.43 79.81$\pm$0.96 35.33$\pm$6.38 84.19$\pm$1.88 47.73$\pm$1.72 OOT
MincutPool 73.54$\pm$5.56 78.90$\pm$0.99 77.84$\pm$1.78 38.00$\pm$2.15 80.65$\pm$2.70 39.02$\pm$5.07 OOT
DMoNPool 75.49$\pm$1.94 74.46$\pm$0.79 72.71$\pm$0.55 41.33$\pm$3.81 82.58$\pm$1.88 51.20$\pm$2.44 OOT
HoscPool 73.05$\pm$2.06 78.61$\pm$0.69 77.71$\pm$1.84 41.11$\pm$4.39 79.68$\pm$6.00 52.80$\pm$2.54 OOT
JustBalancePool 74.51$\pm$2.84 78.61$\pm$0.69 77.71$\pm$1.84 38.00$\pm$4.35 79.68$\pm$6.00 52.09$\pm$1.42 OOT
Table A.5: Results of graph classification using GCNConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods PROTEINS NCI1 NCI109 ENZYMES BZR IMDB-M COLORS-3
Sparse Pooling, pooling ratio = 0.1
TopKPool 69.52$\pm$1.09 58.83$\pm$2.47 60.52$\pm$4.30 29.33$\pm$6.46 83.23$\pm$0.79 33.60$\pm$3.69 18.31$\pm$4.39
SAGPool 70.36$\pm$4.48 63.66$\pm$3.17 63.19$\pm$1.34 26.22$\pm$5.65 84.19$\pm$1.21 39.20$\pm$4.52 32.24$\pm$1.06
ASAPool 76.90$\pm$2.01 64.64$\pm$1.96 65.00$\pm$1.56 25.56$\pm$8.04 81.94$\pm$2.77 32.53$\pm$2.17 34.76$\pm$0.77
PANPool 64.88$\pm$3.63 65.90$\pm$1.09 66.48$\pm$1.21 39.56$\pm$3.49 84.19$\pm$1.21 31.47$\pm$2.00 23.34$\pm$1.36
COPool 67.98$\pm$2.39 61.33$\pm$1.99 60.87$\pm$2.57 36.22$\pm$4.80 83.23$\pm$0.79 40.53$\pm$0.76 31.21$\pm$1.54
CGIPool 70.83$\pm$3.30 66.74$\pm$4.68 66.19$\pm$3.61 35.11$\pm$3.62 85.48$\pm$1.44 40.98$\pm$1.74 36.62$\pm$2.99
KMISPool 75.95$\pm$5.36 74.26$\pm$1.32 70.81$\pm$1.43 23.56$\pm$2.15 86.13$\pm$1.29 30.84$\pm$1.42 57.56$\pm$1.53
GSAPool 68.81$\pm$1.71 60.68$\pm$1.97 63.39$\pm$3.30 30.22$\pm$4.30 80.32$\pm$2.58 33.16$\pm$4.49 23.05$\pm$6.52
HGPSLPool 71.43$\pm$0.75 64.21$\pm$1.73 64.87$\pm$1.65 31.33$\pm$2.15 84.19$\pm$2.58 31.11$\pm$1.35 22.84$\pm$0.74
Sparse Pooling, pooling ratio = 0.3
TopKPool 69.88$\pm$1.10 64.89$\pm$1.20 63.13$\pm$2.24 35.56$\pm$7.73 84.19$\pm$1.58 32.98$\pm$3.38 30.84$\pm$6.31
SAGPool 70.95$\pm$1.52 69.69$\pm$1.66 67.94$\pm$1.58 29.78$\pm$6.34 83.23$\pm$3.76 40.80$\pm$2.77 29.59$\pm$4.17
ASAPool 75.12$\pm$1.90 69.79$\pm$0.77 66.87$\pm$0.84 23.56$\pm$6.02 82.58$\pm$1.88 31.82$\pm$1.90 45.95$\pm$2.82
PANPool 67.74$\pm$1.15 69.92$\pm$0.90 66.16$\pm$1.25 34.22$\pm$2.37 81.94$\pm$1.21 33.51$\pm$1.98 34.99$\pm$2.01
COPool 72.38$\pm$3.03 69.50$\pm$1.27 65.68$\pm$1.64 35.56$\pm$4.50 81.61$\pm$2.62 38.13$\pm$3.49 37.35$\pm$2.24
CGIPool 73.57$\pm$2.28 73.84$\pm$2.73 69.32$\pm$2.52 34.44$\pm$2.98 85.16$\pm$2.37 39.29$\pm$3.11 41.76$\pm$2.47
KMISPool 76.90$\pm$6.38 74.94$\pm$0.36 71.48$\pm$1.18 20.89$\pm$3.68 85.16$\pm$2.37 31.73$\pm$1.48 57.55$\pm$2.67
GSAPool 68.33$\pm$2.14 67.59$\pm$2.22 67.94$\pm$2.34 31.56$\pm$5.65 84.52$\pm$2.19 30.76$\pm$2.15 36.50$\pm$4.39
HGPSLPool 69.40$\pm$1.90 69.37$\pm$1.08 67.71$\pm$1.72 34.22$\pm$2.47 83.87$\pm$1.44 32.62$\pm$1.60 37.51$\pm$1.02
Sparse Pooling, pooling ratio = 0.5
TopKPool 71.19$\pm$2.25 66.61$\pm$1.40 67.71$\pm$2.79 34.67$\pm$7.32 84.52$\pm$1.29 34.04$\pm$1.28 38.33$\pm$5.81
SAGPool 73.57$\pm$1.39 70.50$\pm$1.74 69.19$\pm$2.32 33.56$\pm$2.57 84.52$\pm$1.29 37.24$\pm$1.87 39.72$\pm$2.19
ASAPool 75.71$\pm$1.27 73.29$\pm$1.07 69.26$\pm$1.18 22.00$\pm$4.58 83.23$\pm$2.81 34.76$\pm$2.07 52.72$\pm$1.83
PANPool 70.60$\pm$1.67 73.78$\pm$1.12 70.81$\pm$1.33 31.56$\pm$2.69 83.23$\pm$3.32 34.58$\pm$5.16 43.31$\pm$1.03
COPool 72.62$\pm$2.41 70.89$\pm$1.21 68.61$\pm$1.66 31.33$\pm$4.41 82.90$\pm$2.41 36.44$\pm$1.26 39.77$\pm$2.49
CGIPool 73.57$\pm$1.49 75.72$\pm$1.65 72.90$\pm$0.64 34.00$\pm$3.89 84.19$\pm$2.58 39.82$\pm$1.28 43.66$\pm$3.08
KMISPool 72.26$\pm$5.07 76.40$\pm$0.73 71.65$\pm$1.05 18.44$\pm$3.49 85.48$\pm$2.70 31.56$\pm$1.80 57.88$\pm$1.55
GSAPool 72.14$\pm$1.09 68.65$\pm$1.93 67.84$\pm$1.71 33.33$\pm$4.04 83.55$\pm$1.58 33.33$\pm$2.37 46.14$\pm$3.13
HGPSLPool 71.67$\pm$2.08 71.83$\pm$1.05 69.26$\pm$1.46 30.44$\pm$4.53 83.87$\pm$1.44 31.20$\pm$0.18 50.30$\pm$0.55
Sparse Pooling, pooling ratio = 0.7
TopKPool 70.36$\pm$0.69 68.85$\pm$2.84 68.45$\pm$2.03 33.11$\pm$1.91 84.19$\pm$2.58 32.00$\pm$1.66 45.43$\pm$7.52
SAGPool 68.69$\pm$1.58 72.84$\pm$1.07 70.48$\pm$1.28 34.44$\pm$3.37 83.87$\pm$2.70 38.22$\pm$2.12 45.90$\pm$1.52
ASAPool 75.00$\pm$1.96 72.67$\pm$0.94 70.45$\pm$0.84 25.56$\pm$6.32 86.13$\pm$2.41 31.11$\pm$2.05 58.02$\pm$2.00
PANPool 69.52$\pm$1.43 71.83$\pm$0.87 69.84$\pm$1.07 36.22$\pm$4.69 83.87$\pm$0.00 32.27$\pm$0.96 52.69$\pm$2.90
COPool 73.21$\pm$1.06 72.61$\pm$0.97 70.65$\pm$1.25 40.89$\pm$6.06 84.19$\pm$2.77 40.80$\pm$1.42 41.81$\pm$2.80
CGIPool 72.26$\pm$1.04 73.39$\pm$0.84 71.84$\pm$2.04 36.00$\pm$4.75 82.58$\pm$1.58 40.62$\pm$2.04 44.51$\pm$3.79
KMISPool 76.79$\pm$5.66 74.98$\pm$0.99 71.16$\pm$1.58 23.33$\pm$2.98 85.48$\pm$1.02 31.56$\pm$0.56 58.68$\pm$1.55
GSAPool 71.90$\pm$3.11 70.53$\pm$1.18 68.32$\pm$1.04 36.89$\pm$1.09 83.87$\pm$0.00 32.27$\pm$1.90 49.71$\pm$8.30
HGPSLPool 72.02$\pm$1.73 71.64$\pm$0.72 68.97$\pm$1.43 33.78$\pm$3.03 84.52$\pm$1.29 32.62$\pm$0.92 57.67$\pm$2.07
Sparse Pooling, pooling ratio = 0.9
TopKPool 70.83$\pm$1.25 70.34$\pm$1.80 69.65$\pm$1.61 38.22$\pm$2.78 82.90$\pm$1.29 32.44$\pm$1.32 54.48$\pm$4.48
SAGPool 74.64$\pm$1.53 73.10$\pm$1.21 71.29$\pm$0.82 34.67$\pm$4.52 82.90$\pm$1.29 40.80$\pm$1.39 53.18$\pm$1.71
ASAPool 73.69$\pm$1.48 73.48$\pm$1.03 70.06$\pm$1.06 24.44$\pm$7.44 85.16$\pm$2.37 30.93$\pm$0.87 58.50$\pm$0.68
PANPool 70.48$\pm$0.81 72.77$\pm$1.09 70.84$\pm$1.23 40.67$\pm$4.48 82.90$\pm$1.29 32.98$\pm$1.98 61.86$\pm$1.15
COPool 75.24$\pm$2.46 74.10$\pm$1.06 71.35$\pm$1.05 39.56$\pm$4.19 80.32$\pm$2.14 39.56$\pm$2.47 45.03$\pm$1.19
CGIPool 70.83$\pm$1.46 75.49$\pm$1.11 73.81$\pm$0.42 38.44$\pm$1.51 83.23$\pm$3.62 41.07$\pm$2.35 48.91$\pm$4.14
KMISPool 76.79$\pm$5.73 74.46$\pm$0.98 70.06$\pm$1.66 21.56$\pm$1.33 85.48$\pm$2.28 30.04$\pm$1.81 58.82$\pm$1.21
GSAPool 70.95$\pm$0.69 71.12$\pm$1.33 70.65$\pm$1.45 40.89$\pm$5.04 85.16$\pm$1.88 32.98$\pm$0.99 57.77$\pm$7.03
HGPSLPool 70.95$\pm$1.66 72.22$\pm$0.42 70.35$\pm$1.31 31.11$\pm$3.98 83.55$\pm$0.65 33.07$\pm$1.83 73.46$\pm$1.70
Dense Pooling
AsymCheegerCutPool 77.07$\pm$0.91 73.61$\pm$0.57 73.42$\pm$0.89 45.33$\pm$3.25 73.87$\pm$2.58 43.02$\pm$3.46 OOT
DiffPool 75.24$\pm$0.62 74.98$\pm$1.37 73.84$\pm$1.99 43.78$\pm$6.19 78.06$\pm$4.52 47.73$\pm$1.72 OOT
MincutPool 76.22$\pm$1.22 68.72$\pm$6.67 73.35$\pm$1.28 38.89$\pm$8.78 81.61$\pm$1.29 39.02$\pm$5.07 OOT
DMoNPool 76.71$\pm$0.24 72.32$\pm$0.39 73.42$\pm$0.48 42.67$\pm$5.29 76.77$\pm$4.03 51.20$\pm$2.44 OOT
HoscPool 73.90$\pm$0.60 74.98$\pm$1.34 73.94$\pm$0.77 40.89$\pm$3.25 78.06$\pm$3.62 52.80$\pm$2.54 OOT
JustBalancePool 74.39$\pm$1.39 74.98$\pm$0.85 74.26$\pm$0.48 42.44$\pm$2.76 78.06$\pm$3.62 52.09$\pm$1.42 OOT
Table A.6: Results of node classification using GraphConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods Cora Citeseer Pubmed Cornell Texas Wisconsin Github
Sparse Pooling, pooling ratio = 0.1
TopKPool 86.70$\pm$0.90 74.85$\pm$0.96 86.52$\pm$0.67 81.82$\pm$3.45 74.91$\pm$7.66 73.68$\pm$5.20 82.73$\pm$1.10
SAGPool 87.00$\pm$1.10 74.89$\pm$0.63 85.28$\pm$1.02 79.64$\pm$2.41 74.91$\pm$6.94 72.89$\pm$4.13 83.65$\pm$0.45
ASAPool 84.70$\pm$1.90 74.53$\pm$0.99 70.14$\pm$2.74 82.18$\pm$3.71 77.45$\pm$5.22 73.16$\pm$4.29 OOM
PANPool 81.90$\pm$1.10 73.95$\pm$1.13 72.47$\pm$2.24 78.18$\pm$3.25 76.0$\pm$7.66 73.95$\pm$3.05 OOM
COPool 85.70$\pm$0.40 74.71$\pm$0.99 78.63$\pm$1.81 78.91$\pm$1.45 74.18$\pm$7.75 72.37$\pm$3.22 OOM
CGIPool 86.50$\pm$0.90 74.53$\pm$0.80 81.96$\pm$4.51 80.00$\pm$4.45 75.27$\pm$5.22 72.11$\pm$3.05 79.85$\pm$4.62
KMISPool 77.00$\pm$3.00 74.57$\pm$1.10 64.21$\pm$4.07 80.73$\pm$3.17 75.27$\pm$5.70 73.16$\pm$4.60 83.43$\pm$0.39
GSAPool 87.00$\pm$0.60 74.33$\pm$1.10 85.88$\pm$0.67 80.00$\pm$4.45 76.0$\pm$8.24 73.42$\pm$3.47 83.60$\pm$0.16
HGPSLPool 29.40$\pm$2.40 37.12$\pm$3.31 45.04$\pm$1.56 77.45$\pm$4.96 76.00$\pm$7.58 70.26$\pm$5.30 OOM
Sparse Pooling, pooling ratio = 0.3
TopKPool 86.99$\pm$1.01 75.28$\pm$1.24 86.35$\pm$0.94 80.36$\pm$4.05 75.27$\pm$8.50 73.68$\pm$4.32 82.93$\pm$0.78
SAGPool 87.23$\pm$0.99 74.63$\pm$1.32 85.16$\pm$1.10 81.09$\pm$3.74 75.64$\pm$7.68 72.37$\pm$3.72 83.24$\pm$0.52
ASAPool 77.96$\pm$5.96 74.17$\pm$0.81 72.09$\pm$1.99 79.64$\pm$2.91 71.27$\pm$6.84 67.37$\pm$5.73 OOM
PANPool 44.21$\pm$2.12 64.34$\pm$2.51 68.23$\pm$1.79 78.91$\pm$2.47 74.91$\pm$6.34 73.42$\pm$2.26 OOM
COPool 84.45$\pm$0.96 74.47$\pm$1.08 77.52$\pm$1.76 77.82$\pm$3.13 72.00$\pm$8.65 70.79$\pm$4.19 OOM
CGIPool 86.42$\pm$0.82 74.27$\pm$0.69 82.96$\pm$5.22 81.82$\pm$4.30 76.00$\pm$6.44 72.11$\pm$2.68 79.92$\pm$4.70
KMISPool 77.32$\pm$2.42 74.55$\pm$1.24 65.45$\pm$3.46 81.09$\pm$3.74 75.64$\pm$5.59 72.89$\pm$4.53 83.60$\pm$0.37
GSAPool 87.06$\pm$0.61 74.31$\pm$1.11 85.58$\pm$1.14 80.36$\pm$3.71 76.36$\pm$4.74 72.37$\pm$3.99 82.39$\pm$1.44
HGPSLPool 29.47$\pm$1.21 28.51$\pm$1.59 44.10$\pm$0.91 58.55$\pm$3.71 62.18$\pm$4.66 51.32$\pm$3.99 OOM
Sparse Pooling, pooling ratio = 0.5
TopKPool 86.74$\pm$0.68 75.12$\pm$1.23 86.34$\pm$0.97 81.45$\pm$5.06 72.73$\pm$8.76 76.05$\pm$4.28 82.14$\pm$2.23
SAGPool 86.86$\pm$1.27 74.51$\pm$1.48 85.44$\pm$1.24 80.36$\pm$4.36 73.45$\pm$5.34 73.42$\pm$2.93 79.74$\pm$3.23
ASAPool 70.97$\pm$7.63 73.11$\pm$1.55 69.37$\pm$4.84 72.00$\pm$6.76 65.82$\pm$4.93 62.37$\pm$6.64 OOM
PANPool 47.87$\pm$2.66 55.4$\pm$2.66 67.86$\pm$1.86 78.55$\pm$2.91 76.00$\pm$8.32 71.32$\pm$5.22 OOM
COPool 83.86$\pm$0.94 74.43$\pm$1.09 77.86$\pm$1.88 76.36$\pm$3.45 69.82$\pm$8.02 67.89$\pm$3.59 OOM
CGIPool 86.49$\pm$0.90 74.05$\pm$0.93 82.80$\pm$3.99 82.55$\pm$4.96 74.91$\pm$7.58 72.63$\pm$3.85 81.28$\pm$3.46
KMISPool 76.92$\pm$2.52 74.85$\pm$1.32 64.51$\pm$4.13 80.73$\pm$2.47 76.00$\pm$5.06 73.42$\pm$3.94 83.46$\pm$0.21
GSAPool 86.96$\pm$0.83 74.51$\pm$1.00 85.90$\pm$1.21 81.45$\pm$3.88 78.55$\pm$5.56 73.95$\pm$2.93 80.18$\pm$2.73
HGPSLPool 30.68$\pm$1.26 27.63$\pm$0.88 OOM 55.27$\pm$5.82 58.55$\pm$4.80 50.53$\pm$2.29 OOM
Sparse Pooling, pooling ratio = 0.7
TopKPool 86.86$\pm$0.93 74.93$\pm$1.34 86.09$\pm$0.99 81.09$\pm$4.08 76.00$\pm$6.65 76.58$\pm$5.09 82.71$\pm$1.00
SAGPool 86.17$\pm$1.50 74.31$\pm$1.03 85.16$\pm$1.17 79.27$\pm$4.54 76.36$\pm$4.74 73.42$\pm$3.67 80.78$\pm$1.57
ASAPool 69.69$\pm$6.18 71.51$\pm$3.63 69.34$\pm$2.61 64.36$\pm$5.34 65.82$\pm$3.13 61.84$\pm$3.90 OOM
PANPool 51.56$\pm$2.50 53.45$\pm$0.74 68.01$\pm$2.01 85.16$\pm$3.13 74.55$\pm$6.8 68.42$\pm$5.71 OOM
COPool 84.72$\pm$1.68 74.23$\pm$1.03 78.84$\pm$1.49 75.64$\pm$4.69 68.00$\pm$6.86 62.89$\pm$3.76 OOM
CGIPool 85.46$\pm$2.59 73.97$\pm$1.05 81.07$\pm$2.62 80.73$\pm$3.92 76.36$\pm$5.14 72.11$\pm$3.85 80.36$\pm$3.34
KMISPool 76.09$\pm$3.12 74.85$\pm$1.45 64.89$\pm$3.73 80.73$\pm$4.08 75.64$\pm$5.47 73.95$\pm$4.11 83.41$\pm$0.30
GSAPool 86.45$\pm$0.82 74.33$\pm$1.16 86.10$\pm$1.05 81.09$\pm$3.37 77.45$\pm$6.46 73.95$\pm$2.11 80.40$\pm$4.29
HGPSLPool 30.43$\pm$0.99 27.53$\pm$0.73 OOM 53.82$\pm$2.95 59.64$\pm$3.71 45.53$\pm$2.44 OOM
Sparse Pooling, pooling ratio = 0.9
TopKPool 86.69$\pm$0.92 74.69$\pm$1.03 85.49$\pm$0.67 81.09$\pm$3.92 77.82$\pm$9.01 75.00$\pm$3.43 82.97$\pm$0.58
SAGPool 85.36$\pm$2.02 74.15$\pm$1.29 83.60$\pm$0.70 80.00$\pm$3.25 77.09$\pm$5.34 72.37$\pm$4.71 79.64$\pm$1.98
ASAPool 63.44$\pm$7.05 65.63$\pm$2.88 OOM 61.82$\pm$3.45 62.91$\pm$4.39 59.47$\pm$5.85 OOM
PANPool 52.25$\pm$2.38 52.15$\pm$2.00 70.64$\pm$3.47 79.27$\pm$2.95 71.64$\pm$5.22 67.37$\pm$5.09 OOM
COPool 85.29$\pm$1.34 74.25$\pm$0.75 79.32$\pm$1.72 76.73$\pm$2.12 68.36$\pm$6.57 65.79$\pm$4.71 OOM
CGIPool 86.49$\pm$0.87 73.05$\pm$1.92 79.63$\pm$1.74 81.45$\pm$3.71 75.27$\pm$7.59 73.95$\pm$3.67 81.21$\pm$3.57
KMISPool 76.43$\pm$2.34 74.73$\pm$1.23 64.93$\pm$3.64 80.36$\pm$3.53 76.36$\pm$4.88 72.89$\pm$4.75 83.33$\pm$0.31
GSAPool 84.94$\pm$1.63 74.49$\pm$0.95 83.72$\pm$1.42 82.18$\pm$3.53 77.45$\pm$5.09 73.16$\pm$3.96 79.66$\pm$2.51
HGPSLPool 30.60$\pm$1.28 27.87$\pm$0.52 OOM 57.82$\pm$3.13 64.0$\pm$3.53 51.58$\pm$1.53 OOM
Table A.7: Results of node classification using GCNConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods Cora Citeseer Pubmed Cornell Texas Wisconsin Github
Sparse Pooling, pooling ratio = 0.1
TopKPool 88.91$\pm$0.93 77.56$\pm$0.85 85.73$\pm$0.34 48.00$\pm$0.89 54.18$\pm$4.8 50.79$\pm$3.49 86.95$\pm$0.20
SAGPool 89.18$\pm$0.65 77.56$\pm$0.81 85.83$\pm$0.23 48.73$\pm$3.13 55.64$\pm$2.95 51.32$\pm$3.53 86.83$\pm$0.20
ASAPool 89.03$\pm$0.84 77.76$\pm$1.01 85.74$\pm$0.18 49.82$\pm$1.45 54.55$\pm$4.74 50.53$\pm$3.96 OOM
PANPool 89.05$\pm$0.96 77.20$\pm$0.98 85.88$\pm$0.11 49.82$\pm$2.18 56.00$\pm$4.80 50.53$\pm$3.18 OOM
COPool 89.00$\pm$0.70 77.24$\pm$0.90 85.27$\pm$0.27 47.64$\pm$2.12 54.91$\pm$5.32 51.84$\pm$0.64 86.68$\pm$0.20
CGIPool 89.15$\pm$0.84 77.40$\pm$0.81 85.80$\pm$0.13 49.45$\pm$1.78 54.55$\pm$4.15 50.79$\pm$3.78 86.88$\pm$0.22
KMISPool 89.05$\pm$0.93 77.60$\pm$0.81 85.92$\pm$0.07 48.73$\pm$1.36 54.18$\pm$3.88 52.89$\pm$2.11 86.88$\pm$0.24
GSAPool 89.05$\pm$0.77 77.16$\pm$0.92 85.70$\pm$0.19 49.09$\pm$1.15 54.18$\pm$3.88 51.58$\pm$3.85 87.12$\pm$0.09
HGPSLPool 89.00$\pm$0.64 77.84$\pm$0.67 86.00$\pm$0.16 48.73$\pm$2.12 54.91$\pm$4.21 51.58$\pm$1.75 OOM
Sparse Pooling, pooling ratio = 0.3
TopKPool 88.63$\pm$1.05 77.12$\pm$1.16 85.53$\pm$0.35 49.09$\pm$2.57 54.18$\pm$4.80 51.32$\pm$3.00 86.89$\pm$0.27
SAGPool 88.81$\pm$0.73 77.22$\pm$0.94 85.68$\pm$0.54 49.82$\pm$2.95 55.64$\pm$2.95 51.32$\pm$3.53 86.85$\pm$0.21
ASAPool 89.10$\pm$0.86 77.40$\pm$1.13 OOM 79.64$\pm$2.91 54.55$\pm$5.14 50.53$\pm$3.87 OOM
PANPool 88.98$\pm$0.70 77.16$\pm$0.99 85.64$\pm$0.15 49.82$\pm$2.47 56.36$\pm$5.14 50.00$\pm$2.04 OOM
COPool 88.73$\pm$0.84 77.26$\pm$0.89 84.70$\pm$0.24 48.73$\pm$1.78 56.00$\pm$5.80 51.05$\pm$1.93 86.61$\pm$0.29
CGIPool 89.15$\pm$0.84 77.14$\pm$0.80 85.92$\pm$0.66 49.82$\pm$1.85 54.55$\pm$4.15 51.05$\pm$3.85 86.84$\pm$0.12
KMISPool 89.08$\pm$0.86 77.59$\pm$0.65 85.93$\pm$0.06 48.73$\pm$1.36 54.18$\pm$3.88 52.89$\pm$2.11 86.86$\pm$0.25
GSAPool 88.91$\pm$1.20 77.12$\pm$1.16 86.00$\pm$0.42 49.09$\pm$1.15 54.18$\pm$3.88 51.58$\pm$3.85 86.95$\pm$0.28
HGPSLPool 89.03$\pm$0.84 77.80$\pm$0.71 OOM 48.36$\pm$1.85 55.27$\pm$3.37 51.58$\pm$2.26 OOM
Sparse Pooling, pooling ratio = 0.5
TopKPool 88.81$\pm$1.07 77.08$\pm$1.10 85.60$\pm$0.06 47.30$\pm$1.80 54.18$\pm$4.80 50.53$\pm$3.39 86.94$\pm$0.31
SAGPool 88.98$\pm$0.67 77.10$\pm$0.82 86.07$\pm$0.70 81.09$\pm$3.74 55.64$\pm$2.72 51.32$\pm$3.53 86.88$\pm$0.25
ASAPool 88.86$\pm$0.96 77.18$\pm$1.02 OOM 72.00$\pm$6.76 54.18$\pm$3.88 50.79$\pm$3.49 OOM
PANPool 88.95$\pm$0.73 77.02$\pm$0.90 OOM 78.55$\pm$2.91 56.36$\pm$5.14 49.74$\pm$1.93 OOM
COPool 88.56$\pm$1.02 76.78$\pm$0.98 OOM 76.36$\pm$3.45 56.36$\pm$5.98 52.37$\pm$2.68 86.57$\pm$0.29
CGIPool 89.15$\pm$0.84 77.10$\pm$0.83 85.80$\pm$0.78 81.82$\pm$4.30 54.55$\pm$4.15 50.79$\pm$3.49 86.86$\pm$0.08
KMISPool 89.15$\pm$0.93 77.58$\pm$0.62 85.88$\pm$0.14 81.09$\pm$3.74 54.18$\pm$3.88 52.89$\pm$2.11 86.89$\pm$0.29
GSAPool 88.98$\pm$1.13 76.96$\pm$0.93 86.04$\pm$0.53 80.36$\pm$3.71 54.18$\pm$3.88 51.32$\pm$3.81 86.94$\pm$0.27
HGPSLPool 89.08$\pm$0.90 77.80$\pm$0.96 OOM 58.55$\pm$3.71 55.27$\pm$4.39 50.53$\pm$1.78 OOM
Sparse Pooling, pooling ratio = 0.7
TopKPool 88.44$\pm$1.09 77.08$\pm$0.98 86.04$\pm$0.41 47.30$\pm$1.80 49.09$\pm$1.15 51.05$\pm$3.85 86.76$\pm$0.33
SAGPool 88.91$\pm$0.62 77.08$\pm$0.80 85.94$\pm$0.76 81.09$\pm$3.74 50.55$\pm$1.36 51.05$\pm$3.47 86.99$\pm$0.22
ASAPool 88.76$\pm$0.72 76.74$\pm$1.05 OOM 79.64$\pm$2.91 50.55$\pm$1.36 49.74$\pm$4.11 OOM
PANPool 88.59$\pm$1.01 76.98$\pm$1.03 OOM 78.91$\pm$2.47 50.91$\pm$1.15 49.74$\pm$1.93 OOM
COPool 88.27$\pm$1.09 76.72$\pm$1.07 OOM 77.82$\pm$3.13 50.18$\pm$3.74 51.58$\pm$2.93 86.57$\pm$0.28
CGIPool 89.15$\pm$0.84 77.02$\pm$0.65 85.50$\pm$0.69 81.82$\pm$4.30 49.82$\pm$1.45 50.26$\pm$3.37 86.77$\pm$0.14
KMISPool 89.15$\pm$0.91 77.52$\pm$0.65 85.86$\pm$0.10 81.09$\pm$3.74 50.55$\pm$1.36 52.89$\pm$2.11 86.89$\pm$0.26
GSAPool 88.66$\pm$0.98 77.1$\pm$0.72 86.21$\pm$0.73 80.36$\pm$3.71 49.82$\pm$2.72 51.84$\pm$4.29 86.83$\pm$0.19
HGPSLPool 89.03$\pm$0.87 77.74$\pm$0.85 OOM 58.55$\pm$3.71 49.82$\pm$2.47 50.26$\pm$2.26 OOM
Sparse Pooling, pooling ratio = 0.9
TopKPool 88.44$\pm$0.91 77.08$\pm$0.75 86.13$\pm$0.34 47.30$\pm$1.80 48.73$\pm$2.12 51.58$\pm$3.05 86.80$\pm$0.25
SAGPool 88.83$\pm$0.97 77.08$\pm$0.81 85.89$\pm$0.63 81.09$\pm$3.74 50.55$\pm$1.36 50.79$\pm$3.49 86.87$\pm$0.17
ASAPool 88.98$\pm$0.74 76.76$\pm$0.87 OOM 79.64$\pm$2.91 50.55$\pm$1.36 50.00$\pm$4.32 OOM
PANPool 88.76$\pm$1.11 76.70$\pm$0.92 OOM 78.91$\pm$2.47 51.27$\pm$2.41 49.47$\pm$1.97 OOM
COPool 88.54$\pm$1.09 76.72$\pm$1.41 OOM 77.82$\pm$3.13 49.45$\pm$3.13 51.84$\pm$2.58 86.54$\pm$0.28
CGIPool 89.15$\pm$0.84 77.10$\pm$0.54 84.95$\pm$0.39 81.82$\pm$4.30 50.55$\pm$2.12 50.79$\pm$3.78 86.65$\pm$0.25
KMISPool 89.13$\pm$0.90 77.64$\pm$0.58 85.87$\pm$0.09 81.09$\pm$3.74 50.18$\pm$0.89 52.89$\pm$2.11 86.88$\pm$0.28
GSAPool 88.93$\pm$0.88 76.94$\pm$0.74 86.07$\pm$0.62 80.36$\pm$3.71 49.82$\pm$2.72 51.58$\pm$4.19 86.75$\pm$0.24
HGPSLPool 89.08$\pm$0.83 77.70$\pm$1.01 OOM 58.55$\pm$3.71 49.45$\pm$2.12 50.53$\pm$2.44 OOM