### v2

# A Comprehensive Graph Pooling Benchmark: Effectiveness, Robustness and Generalizability

Abstract
Graph pooling has gained attention for its ability to obtain effective node and graph representations for various downstream tasks. Despite the recent surge in graph pooling approaches, there is a lack of standardized experimental settings and fair benchmarks to evaluate their performance. To address this issue, we have constructed a comprehensive benchmark that includes 15 graph pooling methods and 21 different graph datasets. This benchmark systematically assesses the performance of graph pooling methods in three dimensions, i.e., effectiveness, robustness, and generalizability. We first evaluate the performance of these graph pooling approaches across different tasks including graph classification, graph regression and node classification. Then, we investigate their performance under potential noise attacks and out-of-distribution shifts in real-world scenarios. We also involve detailed efficiency analysis and parameter analysis. Extensive experiments validate the strong capability and applicability of graph pooling approaches in various scenarios, which can provide valuable insights and guidance for deep geometric learning research. The source code of our benchmark is available at https://github.com/goose315/Graph_Pooling_Benchmark.
Keywords Graph Pooling; Benchmark; Graph Neural Networks; Graph Machine Learning
1 Introduction
Recently, graph neural networks (GNNs) have garnered significant attention with extensive benchmarks [1, 2, 3] due to their remarkable ability to process graph-structured data across various domains including social networks [4, 5, 6, 7], rumor detection [8, 9], biological networks [10, 11], recommender systems [12] and community detection [13, 14]. Graph pooling approaches play a crucial role in GNNs by enabling the hierarchical reduction of graph representations, which is essential for capturing multi-scale structures and long-range dependencies [15, 16, 17]. They can preserve crucial topological semantics and relationships, which have shown effective for tasks including graph classification, node clustering, and graph generation [15, 18, 19, 2]. In addition, by aggregating nodes and edges, graph pooling can also simplify large-scale graphs, facilitating the application of GNNs in real-world problems [20, 21, 22, 23, 24]. Therefore, understanding and enhancing graph pooling approaches is the key to increasing GNN performance, driving the progress of deep geometric learning.
In literature, existing graph pooling approaches [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] can be roughly divided into two categories [40, 15], i.e., sparse pooling [25, 26, 27, 28, 29, 30, 31, 32, 33] and dense pooling approaches [34, 35, 36, 37, 38, 39] based on the number of nodes after pooling. Sparse pooling approaches generally keep the number of nodes as the constant cardinality, i.e., $O(1)$ while dense pooling approaches typically have the number of nodes proportional to the number of total nodes after pooling [41, 40]. Even though graph pooling research is becoming increasingly popular, there is still no standardized benchmark that allows for an impartial and consistent comparison of various graph pooling methods. Furthermore, due to the diversity and complexity of graph datasets, numerous experimental settings have been used in previous studies, such as varied proportions of training data and train/validation/test splits [8, 34, 17, 42]. As a result, a comprehensive and publicly available benchmark of graph pooling approaches is highly expected that can facilitate the evaluation and comparison of different approaches, ensuring the reproducibility of results and further advancing the area of graph machine learning.
Towards this end, we present a comprehensive graph pooling benchmark, which includes 15 graph pooling methods and 21 datasets across different graph machine learning problems. In particular, we extensively investigate graph pooling approaches across three key perspectives, i.e., effectiveness, robustness, and generalizability. To begin, we provide a fair and thorough effectiveness comparison of existing graph pooling approaches across graph classification, graph regression and node classification. Then, we evaluate the robustness of graph pooling approaches under both noise attacks on graph structures and node attributes. In addition, we study the generalizability of different approaches under out-of-distribution shifts from both size and density levels. Finally, we include efficiency comparison, parameter analysis and backbone analysis for completeness.
From extensive experimental results, we have four observations as follows: (1) Dense pooling approaches generally outperform pooling on graph classification and graph regression. (2) Different graph pooling approaches have a limited impact on node classification. (3) Feature masking influences the performance of graph pooling more seriously compared with structure perturbation and dense pooling is more robust to noise attack. (4) Most of graph pooling approaches suffer from serious performance decrement from distribution shifts.
The main contributions of this paper are as follows:
  • Comprehensive Benchmark. We present the first comprehensive graph pooling benchmark, which incorporates 15 state-of-the-art graph pooling approaches and 21 diverse datasets across graph classification, graph regression and node classification.
  • Extensive Analysis. To investigate the pros and cons of graph pooling approaches, we thoroughly evaluate current approaches from three perspectives, i.e., effectiveness, robustness, and generalizability, which can serve as guidance for researchers in different applications.
  • Open-source Material. We have made our benchmark of all these graph pooling approaches publicly available and reproducible, and believe our benchmark can benefit researchers in both graph machine learning and interdisciplinary fields.
Figure 1: Overview of our hierarchical backbone for graph classification and graph regression.
2 Preliminaries
Notations. Consider a graph $G$ characterized by a vertex set $V$ and an edge set $E$. The features associated with each vertex are represented by the matrix $\bm{X}\in\mathbb{R}^{|V|\times d}$, where $|V|$ denotes the number of vertices, and $d$ signifies the dimensionality of the attribute vectors. The adjacency relationships within the graph are encapsulated by the adjacency matrix $\bm{A}\in\{0,1\}^{|V|\times|V|}$, where an entry $\bm{A}[i,j]=1$ indicates the presence of an edge between vertex $v_{i}$ and vertex $v_{j}$; otherwise, $\bm{A}[i,j]=0$.
Graph Pooling [41, 43, 40]. The aim of graph pooling is to reduce the spatial size of feature maps while preserving essential semantics, which thereby decreases computational complexity and memory usage. In this work, we focus on hierarchical pooling approaches [21]. Let POOL denote a graph pooling function which maps $G$ to a graph $G^{\prime}=(V^{\prime},E^{\prime})$ with the reduced size:
${G}^{\prime}=\texttt{POOL}({G}),$ (1)
where $|V^{\prime}|<|V|$. the process has two main principal components, i.e., reduction, and connection [43]. In particular, the reduction component aims to generate pooled nodes and their attributes in ${G}^{\prime}$ while the connection. The connection component computes the edges $E^{\prime}$ among the $V^{\prime}$ nodes.
Figure 2: Overview of the graph U-Net framework for node classification.
Graph Classification and Regression [25, 44, 43]. The two primary graph-level tasks are graph regression and graph classification. Here, a graph dataset $\mathcal{G}$ is provided as a set of graph-label pairs $(G_{i},y_{i})$, where $y_{i}$ denotes the label for graph $G_{i}$. The objective is to train a powerful discriminative model $f$ that predicts the correct label $y_{i}$ given an input graph $G_{i}$. In graph classification, $y_{i}$ are categorical labels ${1,\cdots,K}$ with $K$ as the number of classes, while in graph regression, $y_{i}$ are continuous values. A well-trained graph classification model should output labels that closely match the true labels, and similarly, a graph regression model should predict values that are nearly identical to the ground truth values. In these tasks, graph pooling always accompanies graph convolutional operators. In formulation, the basic updating rule is written as follows:
$\bm{H}^{(l+1)}=\sigma(\tilde{\bm{D}}^{-\frac{1}{2}}\tilde{\bm{A}}\tilde{D}^{-\frac{1}{2}}\bm{H}^{(l)}\bm{W}^{(l)}),$ (2)
where $\bm{H}^{(l)}$ denotes the node feature matrix at layer $l$, $\bm{W}^{(l)}$ denotes the weight matrix at the crossponding layer, $\tilde{\bm{A}}=\bm{A}+\bm{I}$ is the adjacency matrix $\bm{A}$ plus the identity matrix $\bm{I}$, $\tilde{\bm{D}}$ is the degree matrix of $\tilde{\bm{A}}$, and $\sigma$ is a nonlinear activation function [45]. The pooling layers can be formulated as:
$\bm{H}^{(pool)}=\texttt{POOL}(\bm{H}^{(L)}),$ (3)
where $\bm{H}^{(pool)}$ is the node feature matrix after pooling We iteratively conduct graph convolution and graph pooling operators and adopt a readout function to output the graph representation for downstream tasks. The overview of the basic hierarchical backbone can be found in Figure 1.
Node Classification [45, 46, 47]. The aim of node classification is to assign semantic labels to nodes in a graph according to their attributes and relationships with different nodes. Each dataset involves a graph $G$, consisting of nodes $v_{i}$ and their corresponding labels $y_{i}$. $|V|$ is divided into a labeled set $V^{l}$ and a unlabeled set $V^{u}$. We are required to train a graph neural network model that can predict the missing labels of nodes in $V^{u}$ using the attributes of other nodes. As shown in Figure 2, U-Net framework [48] is widely used to incorporate pooling operations for node classification. In the encoder part, U-Net progressively applies pooling and graph convolution to downsample the graphs and extract multi-scale features. The decoder part of U-Net utilizes upsampling and graph convolution to gradually upsample the low-resolution feature maps back to the original graph size. Residual connections are employed to directly transfer the feature maps from the encoder to the decoder, facilitating the preservation of fine-grained semantics during upsampling [48, 49, 50].
3 Graph Pooling Benchmark
In this section, we present an overview of our graph pooling benchmark, which includes extensive state-of-the-art approaches, graph datasets, and evaluation protocols.
3.1 Graph Pooling Approaches
Our benchmark contains 15 state-of-the-art graph pooling approaches, which can be divided into two categories, i.e., sparse pooling and dense pooling based on the number of nodes after pooling. Table 1 summarizes the experimental details in their corresponding papers.
Sparse Pooling. Sparse pooling approaches typically maintain the number of nodes as a constant cardinality $O(1)$ after graph pooling [43, 40]. These approaches usually utilize a score function to evaluate the importance of each node and keep the nodes with high scores and are easy to implement using sparse matrix calculation. Therefore, they require significantly fewer computational resources, especially when it comes to large-scale graphs. In our benchmark, we select 9 sparse pooling approaches including TopKPool [25], SAGPool [26], ASAPool [27], PANPool [28], COPool [29], CGIPool [30], KMISPool [31], GSAPool [32], and HGPSLPool [33].
Dense Pooling. Dense pooling approaches usually have the number of nodes after pooling proportional to the original node numbers, i.e., $O(|V|)$. These approaches typically adopt graph clustering to learn a new coarsened graph, which can provide a comprehensive semantic view by summarizing similar nodes. One limitation of these approaches is the potential high complexity. In this work, we select 6 dense pooling approaches including AsymCheegerCutPool [34], DiffPool [35], MincutPool [36], DMoNPool [37], HoscPool [38], and JustBalancePool [39].
3.2 Datasets
To systematically evaluate different graph pooling methods, we integrate 21 datasets from different domains for three types of tasks. For graph classification, we select eight publicly available datasets from TUDataset [51], including three molecules datasets, i.e., BZR [52], NCI1 [53], and NCI109 [53], four bioinformatics datasets, i.e., PROTEINS, PROTEINS_full [54], D&D [55], and ENZYMES [56], one social network dataset, i.e., IMDB-MULTI [57], and one synthetic dataset, i.e., COLORS-3 [58]. For graph regression, we choose five datasets from previous works [10] including QM8, BACE, ESOL, FreeSolv, and Lipophilicity. For node classification, we utilize three citation network datasets, i.e., Cora, Citeseer, Pubmed [59] and three website networks, i.e., Cornell, Texas, and Wisconsin [60] and the GitHub dataset [61]. More information can be found in Appendix.
Table 1: Overview of experimental details of graph pooling research. These papers utilize different settings, which validates the necessity of building a comprehensive and fair graph pooling benchmark.
Methods Datasets Tasks
Sparse Pooling
TopKPool NIPS’19 MNIST, COLLAB, PROTEINS, D&D Graph Classification
SAGPool ICML’19 D&D, PROTEINS, NCI1, NCI109, FRANKENSTEIN Graph Classification
ASAPool AAAI’20 D&D, PROTEINS, NCI1, NCI109, FRANKENSTEIN Graph Classification
PANPool NIPS’20 PROTEINS, PROTEINS_FULL, NCI1, AIDS, MUTAGENCITY Graph Classification
COPool ECMLPKDD’22 BZR, AIDS, NCI1, NCI109, PROTEINS, QM7, IMDB-M Graph Classification, Graph Regression
CGIPool SIGIR’22 NCI1, NCI109, Mutagenicity, IMDB-B, IMDB-M, COLLAB PROTEINS Graph Classification
KMISPool AAAI’23 D&D, REDDIT-B, REDDIT-5K, REDDIT-12K, Github Graph Classification, Node Classification
GSAPool WWW’20 D&D, NCI1, NCI109, Mutagencity Graph Classification
HGPSLPool Arxiv’19 D&D, PROTEINS, NCI1, NCI109, ENZYMES, Mutagencity Graph Classification
Dense Pooling
AsymCheegerCutPool ICML’23 Cora, Citeseer, Pubmed, DBLP Node Classification
DiffPool NIPS’18 D&D, PROTEINS, COLLAB, ENZYMES, REDDIT-MULTI Graph Classification
MincutPool ICML’20 D&D, PROTEINS, COLLAB, REDDIT-B, Mutagenicity, QM9 Graph Classification, Graph Regression
DMoNPool JMLR’23 Cora, Citeseer Pubmed, Coauthor Node Classification
HoscPool CIKM’22 Cora, Citeseer Pubmed, Coauthor, DBLP, Email-EU Node Classification
JustBalancePool Arxiv’22 Cora, Citeseer, Pubmed, DBLP Node Classification
3.3 Evaluation Protocols
Our benchmark evaluation encompasses three key aspects, i.e., effectiveness, robustness, and generalizability. Firstly, we conduct a performance comparison of graph pooling approaches across three tasks including graph classification, graph regression, and node classification. For graph and node classification tasks, we employ accuracy as the evaluation metric. For graph regression, we use root mean square error (RMSE) for ESOL, FreeSolv, and Lipophilicity [10]. Following previous research [42], we use the area under the receiver operating characteristic (AUROC) curve to evaluate BACE, and Mean Absolute Error (MAE) for QM8. Secondly, our benchmark evaluates the robustness of graph pooling approaches across two views, i.e., structural robustness and feature robustness [62]. In particular, we add and drop edges of graphs to study structural robustness and mask node features to investigate feature robustness. Thirdly, we employ size-based and density-based distribution shifts to evaluate the generalizability of different pooling methods under real-world scenarios [63]. In addition to these three views, we conduct a further analysis of these graph pooling approaches including the comparison of efficiency, and different backbone parameter choices.
4 Experiment
Table 2: Results of graph classification for different graph pooling methods. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods PROTEINS NCI1 NCI109 ENZYMES BZR IMDB-M COLORS-3 Avg. Rank
Sparse Pooling
TopKPool 72.74$\pm$2.36 69.08$\pm$2.43 67.65$\pm$2.11 38.22$\pm$7.59 83.23$\pm$2.41 41.21$\pm$0.34 43.88$\pm$5.39 62.02 12.50
SAGPool 75.71$\pm$1.90 71.77$\pm$2.31 71.29$\pm$1.98 38.89$\pm$5.35 82.90$\pm$2.19 45.43$\pm$2.08 45.27$\pm$2.89 64.33 8.50
ASAPool 73.45$\pm$2.18 74.81$\pm$1.73 71.97$\pm$0.62 39.33$\pm$8.68 86.45$\pm$2.99 45.72$\pm$1.63 42.79$\pm$2.02 65.29 6.83
PANPool 73.69$\pm$1.38 74.75$\pm$1.45 70.03$\pm$1.01 39.56$\pm$2.86 85.16$\pm$1.21 38.29$\pm$0.31 32.00$\pm$3.68 63.58 8.83
COPool 73.69$\pm$4.54 71.25$\pm$1.24 69.90$\pm$3.45 42.89$\pm$3.49 80.00$\pm$3.47 48.68$\pm$0.75 38.44$\pm$3.56 64.40 8.33
CGIPool 72.26$\pm$1.79 76.24$\pm$2.04 71.81$\pm$1.41 40.00$\pm$5.26 84.84$\pm$2.81 50.36$\pm$0.38 37.78$\pm$5.35 65.92 7.50
KMISPool 72.98$\pm$1.10 79.06$\pm$0.82 72.23$\pm$0.67 20.00$\pm$5.71 87.10$\pm$3.38 40.37$\pm$0.38 20.67$\pm$3.76 61.96 8.17
GSAPool 72.50$\pm$1.90 69.40$\pm$1.55 68.65$\pm$1.81 35.78$\pm$3.74 83.87$\pm$0.00 41.35$\pm$1.28 38.44$\pm$4.19 61.93 12.17
HGPSLPool 73.45$\pm$2.49 75.24$\pm$0.89 68.74$\pm$0.45 43.56$\pm$2.93 88.06$\pm$1.64 39.45$\pm$0.28 41.33$\pm$1.91 64.75 7.83
Dense Pooling
AsymCheegerCutPool 76.59$\pm$1.47 74.75$\pm$0.12 72.13$\pm$0.69 43.78$\pm$5.52 84.52$\pm$2.41 49.12$\pm$0.43 OOT 66.82 5.00
DiffPool 73.05$\pm$1.98 78.35$\pm$0.85 77.00$\pm$1.81 46.00$\pm$3.62 75.16$\pm$1.64 48.48$\pm$0.33 OOT 66.34 6.00
MincutPool 73.41$\pm$2.74 77.60$\pm$1.51 76.58$\pm$1.45 42.89$\pm$2.69 81.61$\pm$3.90 42.44$\pm$0.20 OOT 65.76 6.50
DMoNPool 76.71$\pm$1.18 71.67$\pm$0.58 72.32$\pm$1.48 48.89$\pm$3.78 75.81$\pm$1.77 41.14$\pm$0.34 OOT 64.42 7.50
HoscPool 73.17$\pm$2.00 77.44$\pm$1.79 73.94$\pm$1.53 46.67$\pm$3.78 78.39$\pm$3.16 42.34$\pm$0.58 OOT 65.33 6.67
JustBalancePool 73.17$\pm$2.00 77.63$\pm$1.85 73.94$\pm$1.53 43.78$\pm$3.76 78.39$\pm$3.16 43.08$\pm$0.42 OOT 65.00 6.33
4.1 Experimental Settings
All graph pooling methods in our benchmark are implemented by PyTorch [64]. Graph convolutional networks serve as the default encoders for all algorithms unless otherwise specified. The experimental setup includes a Linux server equipped with NVIDIA L4 and NVIDIA A100 GPUs, as well as an Intel Xeon Gold 6354 CPU. The software stack comprises PyTorch version 1.11.0, PyTorch-geometric version 2.1.0 [65], and Python version 3.9.16.
4.2 Effectiveness Analysis
Performance on Graph Classification. To begin, we investigate the performance of different graph pooling approaches on graph classification. The results of compared approaches on seven popular datasets are recorded in Table 2. From the results, we have the following observations. Firstly, in general, dense pooling methods outperform sparse pooling methods. However, dense pooling methods encounter memory bottlenecks when processing large datasets with tens of thousands of graphs, such as COLORS-3. Secondly, two dense pooling approaches, AsymCheegerCutPool and DiffPool are among the top out of all the compared methods, achieving optimal or near-optimal results on five datasets. The potential reasons include that AsymCheegerCutPool minimizes the graph total variation of the cluster assignments, and DiffPool improves the quality of membership assignments by using a link prediction auxiliary objective [35, 34]. Thirdly, among the sparse pooling methods, ASAPool, CGIPool, and HGPSLPool demonstrate the best overall performance. To be specific, CGIPool achieves the best results on IMDB-M, while HGPSLPool achieves the best results on BZR.
Table 3: Results of graph regression for different pooling methods.
Methods QM8 BACE ESOL FreeSolv Lipophilicity Avg. Rank
Sparse Pooling
TopKPool 0.026 $\pm$ 0.000 0.79 $\pm$ 0.01 1.14 $\pm$ 0.06 1.91 $\pm$ 0.30 0.99 $\pm$ 0.09 1.35 10.67
SAGPool 0.024 $\pm$ 0.000 0.78 $\pm$ 0.01 1.49 $\pm$ 0.28 1.87 $\pm$ 0.06 1.00 $\pm$ 0.09 1.45 12.33
ASAPool 0.025 $\pm$ 0.002 0.77 $\pm$ 0.04 1.06 $\pm$ 0.13 2.52 $\pm$ 0.38 0.95 $\pm$ 0.09 1.51 10.67
PANPool 0.024 $\pm$ 0.000 0.75 $\pm$ 0.02 1.31 $\pm$ 0.05 3.42 $\pm$ 0.32 1.11 $\pm$ 0.02 1.95 14.67
COPool 0.024 $\pm$ 0.002 0.81 $\pm$ 1.56 1.21$\pm$ 0.10 2.53 $\pm$ 0.26 0.94 $\pm$ 0.03 1.56 11.67
CGIPool 0.020 $\pm$ 0.001 0.83 $\pm$ 0.01 1.09 $\pm$ 0.19 1.78 $\pm$ 0.17 0.91 $\pm$ 0.05 1.26 8.00
KMISPool 0.018 $\pm$ 0.001 0.81 $\pm$ 0.01 0.80 $\pm$ 0.03 1.56 $\pm$ 0.65 0.77 $\pm$ 0.01 1.04 5.67
GSAPool 0.025 $\pm$ 0.001 0.81 $\pm$ 0.02 1.22 $\pm$ 0.10 1.95 $\pm$ 0.32 0.99 $\pm$ 0.08 1.39 11.67
HGPSLPool 0.019 $\pm$ 0.000 0.81 $\pm$ 0.02 1.02 $\pm$ 0.14 2.00 $\pm$ 0.31 0.93 $\pm$ 0.03 1.32 9.00
Dense Pooling
AsymCheegerCutPool 0.031 $\pm$ 0.005 0.84 $\pm$ 0.01 0.99 $\pm$ 0.12 2.00 $\pm$ 0.18 0.95 $\pm$ 0.11 1.31 9.33
DiffPool 0.037 $\pm$ 0.010 0.84 $\pm$ 0.02 0.81 $\pm$ 0.05 1.20 $\pm$ 0.09 0.73 $\pm$ 0.03 0.91 4.67
MincutPool 0.020 $\pm$ 0.001 0.85 $\pm$ 0.02 0.76 $\pm$ 0.02 1.19 $\pm$ 0.18 0.73 $\pm$ 0.02 0.89 3.33
DMoNPool 0.021 $\pm$ 0.001 0.85 $\pm$ 0.02 0.68 $\pm$ 0.02 1.16 $\pm$ 0.15 0.69 $\pm$ 0.02 0.84 1.33
HoscPool 0.019 $\pm$ 0.002 0.84 $\pm$ 0.01 0.76 $\pm$ 0.02 1.14 $\pm$ 0.13 0.72 $\pm$ 0.02 0.87 2.33
JustBalancePool 0.022 $\pm$ 0.004 0.85 $\pm$ 0.02 0.74 $\pm$ 0.03 1.26 $\pm$ 0.16 0.70 $\pm$ 0.01 0.90 3.00
Performance on Graph Regression. We further explore the performance of different pooling methods through graph-level regression tasks. As shown in Table 3, we can observe the following: Firstly, dense pooling methods significantly outperform sparse pooling methods. DMoNPool, HoscPool, and JustBalancePool are the best-performing pooling methods, with DMoNPool achieving optimal or near-optimal results on four datasets, and JustBalancePool achieving optimal or near-optimal results on three datasets. Secondly, KMISPool performs well among sparse pooling methods, outperforming AsymCheegerCutPool in dense pooling methods.
Performance on Node Classification. Table 4 presents the performance of various sparse pooling methods in node classification tasks. We can observe that: Firstly, different pooling approaches have consistent performance across these datasets. The potential reason is that graph convolution is the core of semantics exploration on the node level. Secondly, TopKPool demonstrates the best performance among the methods used, potentially because it can adapt to the irregularity of graph data and extract information from the k most important nodes [25]. GSAPool achieves the best or second-best accuracy in 5 out of 7 datasets, as previously mentioned studies have indicated. Before discarding less important nodes, GSAPool aggregates the features of nodes, ensuring that the pooled nodes contain sufficient effective graph information [32]. This aggregation strategy ensures that the selected nodes not only contain their own features but also integrate information from neighboring nodes. Thirdly, the scalability of ASAPool, PANPool, and HGPSLPool still requires improvement, as they are unable to complete training on larger datasets such as GitHub.
Table 4: Results of node classification for different pooling methods.
Methods Cora Citeseer Pubmed Cornell Texas Wisconsin Github Avg. Rank
TopKPool 89.67$\pm$0.18 77.58$\pm$0.13 87.90$\pm$0.15 79.99$\pm$0.83 82.95$\pm$1.66 82.23$\pm$1.13 87.12$\pm$0.05 82.48 2.80
SAGPool 89.57$\pm$0.18 77.37$\pm$0.12 87.85$\pm$0.16 80.10$\pm$1.58 82.41$\pm$1.06 81.99$\pm$1.07 86.87$\pm$0.04 82.29 4.60
ASAPool 89.70$\pm$0.14 77.57$\pm$0.24 87.59$\pm$0.15 78.89$\pm$1.27 82.52$\pm$0.58 82.55$\pm$0.96 OOM 82.25 3.80
PANPool 89.36$\pm$0.26 77.45$\pm$0.20 87.60$\pm$0.09 79.33$\pm$0.86 81.63$\pm$0.74 81.99$\pm$1.51 OOM 81.95 6.00
COPool 89.29$\pm$0.25 77.36$\pm$0.23 87.52$\pm$0.23 78.25$\pm$0.65 82.97$\pm$1.26 82.23$\pm$0.66 86.25$\pm$0.45 82.02 5.80
CGIPool 89.03$\pm$0.43 77.31$\pm$0.26 87.76$\pm$0.21 79.88$\pm$2.14 81.54$\pm$1.14 81.19$\pm$1.11 84.09$\pm$1.31 81.79 7.80
KMISPool 89.75$\pm$0.17 77.69$\pm$0.27 87.80$\pm$0.06 79.12$\pm$0.91 82.08$\pm$0.98 81.12$\pm$1.19 87.09$\pm$0.04 81.95 4.80
GSAPool 89.61$\pm$0.16 77.61$\pm$0.22 87.89$\pm$0.04 80.01$\pm$0.56 83.50$\pm$1.23 81.75$\pm$0.68 87.19$\pm$0.04 82.50 3.20
HGPSLPool 89.63$\pm$0.24 77.50$\pm$0.18 OOM 79.24$\pm$0.76 82.08$\pm$0.63 81.67$\pm$1.14 OOM 82.02 5.60
Table 5: Results of graph classification under random noise attack for different pooling methods.
Dataset Ptb Method TopKPool SAGPool ASAPool CoPool KMISPool DiffPool MincutPool JustBalancePool
PROTEINS RAW 72.74$\pm$2.36 75.71$\pm$1.90 73.45$\pm$2.18 73.69$\pm$4.54 72.98$\pm$1.10 73.05$\pm$1.98 73.41$\pm$2.74 73.17$\pm$2.00
ADD 66.71$\pm$1.99 70.49$\pm$2.06 64.15$\pm$1.41 63.54$\pm$1.98 59.88$\pm$2.26 71.59$\pm$1.42 69.63$\pm$1.98 70.12$\pm$0.67
DROP 72.50$\pm$2.21 69.09$\pm$1.44 68.17$\pm$1.89 68.41$\pm$2.45 63.72$\pm$4.11 70.37$\pm$3.06 68.84$\pm$4.58 73.05$\pm$1.51
MASK 65.49$\pm$4.24 62.87$\pm$2.89 63.35$\pm$2.89 64.88$\pm$4.71 61.34$\pm$0.56 69.63$\pm$3.91 68.72$\pm$2.33 69.39$\pm$3.39
NCI1 RAW 69.08$\pm$2.43 71.77$\pm$2.31 74.81$\pm$1.73 71.25$\pm$1.24 79.06$\pm$0.82 78.35$\pm$0.85 77.60$\pm$1.51 77.63$\pm$1.85
ADD 62.14$\pm$2.19 65.71$\pm$1.03 63.76$\pm$2.18 63.37$\pm$1.11 62.85$\pm$0.54 70.47$\pm$2.08 66.52$\pm$1.94 70.11$\pm$1.32
DROP 72.51$\pm$0.30 71.11$\pm$0.79 72.99$\pm$0.81 72.46$\pm$0.64 74.66$\pm$0.27 69.66$\pm$0.95 72.65$\pm$0.56 72.63$\pm$0.50
MASK 56.40$\pm$3.65 64.76$\pm$1.61 66.09$\pm$0.76 64.08$\pm$1.37 65.09$\pm$0.91 66.84$\pm$0.95 63.82$\pm$3.06 67.52$\pm$2.48
NCI109 RAW 67.65$\pm$2.11 71.29$\pm$1.9 71.97$\pm$0.62 69.90$\pm$3.45 72.23$\pm$0.67 77.00$\pm$1.81 76.58$\pm$1.45 73.94$\pm$1.53
ADD 66.81$\pm$1.44 56.23$\pm$1.63 60.32$\pm$2.06 63.87$\pm$2.39 62.48$\pm$1.14 64.32$\pm$1.60 65.52$\pm$2.24 69.00$\pm$1.84
DROP 61.00$\pm$1.52 64.39$\pm$1.45 63.39$\pm$1.55 62.06$\pm$1.90 61.29$\pm$1.39 64.26$\pm$2.58 62.42$\pm$1.66 67.48$\pm$1.66
MASK 59.45$\pm$2.88 62.81$\pm$1.61 65.16$\pm$0.94 61.94$\pm$1.27 64.35$\pm$0.72 64.00$\pm$1.74 63.23$\pm$0.95 65.77$\pm$1.96
ENZYMES RAW 38.22$\pm$7.59 38.89$\pm$5.35 39.33$\pm$8.68 42.89$\pm$3.49 20.00$\pm$5.71 46.00$\pm$3.62 42.89$\pm$2.69 43.78$\pm$3.76
ADD 40.44$\pm$2.29 37.56$\pm$0.83 20.67$\pm$1.81 38.89$\pm$2.22 35.56$\pm$4.27 35.33$\pm$5.32 30.67$\pm$3.11 33.33$\pm$3.14
DROP 44.00$\pm$2.29 39.33$\pm$3.11 24.44$\pm$0.60 42.00$\pm$1.47 14.22$\pm$1.78 30.22$\pm$3.81 24.44$\pm$4.61 31.56$\pm$5.56
MASK 35.78$\pm$4.06 33.33$\pm$1.86 16.44$\pm$3.47 30.67$\pm$3.56 16.67$\pm$0.00 23.11$\pm$2.57 22.00$\pm$6.02 23.11$\pm$3.68
BZR RAW 83.23$\pm$2.41 82.90$\pm$2.19 86.45$\pm$2.99 80.00$\pm$3.47 87.10$\pm$3.38 75.16$\pm$1.64 81.61$\pm$3.90 78.39$\pm$3.16
ADD 75.81$\pm$1.77 80.97$\pm$1.21 88.39$\pm$1.88 74.52$\pm$4.38 69.35$\pm$3.38 80.32$\pm$1.58 81.94$\pm$2.37 82.26$\pm$2.04
DROP 76.45$\pm$0.79 71.61$\pm$3.32 74.19$\pm$0.00 74.19$\pm$1.44 79.35$\pm$0.65 72.90$\pm$1.21 76.45$\pm$2.99 76.77$\pm$2.19
MASK 70.00$\pm$2.99 78.39$\pm$3.16 64.52$\pm$8.22 78.06$\pm$3.32 71.94$\pm$0.79 80.32$\pm$2.96 78.71$\pm$4.00 81.29$\pm$3.62
Figure 3: Performance of different approaches w.r.t. different rates of random noise.
4.3 Robustness Analysis
The compared performance for three types of random noise on eight graph pooling methods on the PROTEINS, NCI1, NCI109, ENZYMES, and BZR datasets are shown in Table LABEL:noise. With a probability of 50%, edges of the graph are randomly removed or added, and node features are randomly masked with the same likelihood. From the results, we have the following observations. Firstly, masking node features would degrade the performance the most, which demonstrates that node attributes are still the most crucial for graph representation learning. Secondly, the performance ranking of different methods is consistent under different noise attacks but not consistent across different datasets. This indicates that we can choose the optimal graph pooling approaches by validating the performance under any type of noise attack in practice. Thirdly, overall, dense pooling methods exhibit stronger resistance to random noise attacks compared to sparse pooling methods. For instance, JustBalancePool achieves optimal or near-optimal results on three datasets, and DiffPool also performs exceptionally well on most datasets. In addition, sparse pooling methods significantly outperform dense pooling methods on the multi-classification dataset ENZYMES, as evidenced by the performance of TopKPool and COPool. This indicate that the choice of pooling method can be tailored based on different classification tasks.
As depicted in Figure 3, the model’s performance generally declines as the noise intensity increases. It is observed that at the same level of noise, the accuracy decrease is more evident on NCI109, while it is relatively minor on ENZYMES. The accuracy degradation due to edge addition noise is less than that caused by node masking noise and edge deletion noise. Among the three types of noise, although the accuracy of nearly all methods decreases with fluctuations, JustbalancePool and DiffPool exhibit the strongest robustness. For the NCI1 dataset, MincutPool shows significant variance in terms of accuracy.
Table 6: Results of graph classification under distribution shifts. Size and density denote two popular types of shifts across training and test datasets.
Method D&D PROTEINS NCI1 NCI109
Size Density Size Density Size Density Size Density
Sparse Pooling
TopKPool 56.27$\pm$8.82 51.36$\pm$4.27 49.29$\pm$6.74 43.69$\pm$6.26 28.59$\pm$7.73 45.88$\pm$5.69 24.18$\pm$2.05 53.00$\pm$5.00
SAGPool 50.51$\pm$9.52 52.97$\pm$2.67 39.11$\pm$6.90 49.19$\pm$6.30 27.86$\pm$4.15 52.95$\pm$3.06 25.89$\pm$2.91 56.91$\pm$2.59
ASAPool 62.88$\pm$6.87 52.88$\pm$5.99 56.07$\pm$11.66 47.39$\pm$7.12 23.93$\pm$2.22 67.02$\pm$2.25 23.47$\pm$1.48 63.54$\pm$1.70
PANPool 56.1$\pm$5.73 51.86$\pm$5.88 43.39$\pm$14.91 33.69$\pm$15.52 20.24$\pm$0.57 55.46$\pm$2.18 23.56$\pm$2.64 55.18$\pm$2.77
COPool 55.93$\pm$9.94 49.49$\pm$4.88 49.29$\pm$8.21 52.34$\pm$6.70 24.47$\pm$1.34 53.63$\pm$3.40 26.98$\pm$2.92 52.98$\pm$2.62
CGIPool 63.56$\pm$3.83 52.37$\pm$3.81 49.29$\pm$17.31 44.95$\pm$11.85 28.83$\pm$5.88 58.61$\pm$0.88 27.55$\pm$2.45 55.78$\pm$3.12
KMISPool 75.08$\pm$4.37 57.37$\pm$3.68 84.11$\pm$4.70 61.35$\pm$10.65 69.61$\pm$3.09 61.75$\pm$2.76 65.37$\pm$9.00 62.62$\pm$1.67
GSAPool 53.56$\pm$9.34 50.00$\pm$4.30 42.68$\pm$7.54 44.32$\pm$9.66 28.5$\pm$2.51 53.03$\pm$3.76 26.46$\pm$6.55 54.68$\pm$1.57
HGPSLPool 54.41$\pm$5.71 52.97$\pm$2.31 50.71$\pm$6.32 53.42$\pm$3.22 28.01$\pm$0 56.44$\pm$2.16 23.94$\pm$2.18 58.24$\pm$1.39
Dense Pooling
AsymCheegerCutPool OOM OOM 75.71$\pm$3.85 67.21$\pm$6.76 23.83$\pm$5.14 72.18$\pm$0.90 55.25$\pm$24.14 70.35$\pm$0.46
DiffPool 59.83$\pm$7.63 48.77$\pm$14.98 70.00$\pm$5.43 76.04$\pm$5.21 30.73$\pm$6.32 71.35$\pm$1.51 35.82$\pm$21.10 71.43$\pm$1.54
MincutPool 63.39$\pm$10.32 60.17$\pm$22.57 65.00$\pm$9.01 63.15$\pm$8.51 21.21$\pm$2.50 66.12$\pm$4.96 24.70$\pm$3.32 70.04$\pm$1.58
DMoNPool 60.17$\pm$4.12 39.72$\pm$31.15 66.79$\pm$2.22 64.59$\pm$6.78 22.67$\pm$3.09 69.65$\pm$1.72 41.05$\pm$21.66 68.83$\pm$1.49
HoscPool 59.83$\pm$5.53 71.23$\pm$11.4 67.86$\pm$6.63 70.99$\pm$4.19 33.64$\pm$11.34 69.60$\pm$1.77 29.88$\pm$6.60 69.96$\pm$1.75
JustBalancePool 59.83$\pm$5.53 71.23$\pm$11.4 69.64$\pm$3.04 69.10$\pm$3.22 25.39$\pm$3.27 69.60$\pm$1.77 38.48$\pm$19.09 68.63$\pm$1.69
Figure 4: Performance vs Training time and space of different approaches.
4.4 Generalizability Analysis
Table 6 presents the performance of all 15 graph pooling methods under out-of-distribution shifts. For the datasets D&D, PROTEINS, NCI1, and NCI109, two types of distribution shifts are implemented. The first type is based on the number of nodes, where the smallest 50% of graphs by node count are used as the training set, and the largest 20% as the test set, with the remaining graphs serving as the validation set [66, 67]. Following the same criteria, the second type of out-of-distribution shifts are generated based on graph density [67]. For further details, please refer to the Appendix.
From Table 6, we have the following observations. Firstly, the majority of graph pooling approaches suffer from serious performance decreases when it comes to out-of-distribution shifts. For example, most of the approaches have the accuracy below 35% when it comes to size-level distribution shifts. Secondly, for the size-based out-of-distribution shift, KMISPool exhibits superior performance compared with the other approaches in all cases. The underlying reason is its ability to preserve key topological features in the coarsened graph [31]. Thirdly, for the density-based out-of-distribution shift, DiffPool demonstrates optimal performance. The potential reason is that it adopts a learnable graph clustering process to adaptively explore hierarchical semantics [35, 43].
4.5 Further Analysis
Efficiency Comparison. In this part, we conduct an efficiency analysis of graph pooling methods on the ENZYMES and NCI1 datasets. We calculate the time of different algorithms needed to complete 200 epochs of training with the 256 batch size. For space efficiency, we compute the GPU memory utilization during the training process. The compared results are shown in Figure 4. From the results, dense pooling methods such as DiffPool, MincutPool, and JustBalancePool have significantly higher time and space costs. Notably, although KMISPool performs poorly on the multi-classification dataset ENZYMES, it has one of the lowest time and space costs across all three datasets. Conversely, COPool has the highest GPU memory usage among the eight compared methods.
Backbone Analysis. Figure 5 presents the performance of four pooling methods based on GCNConv [68], TAGConv [69], SAGEConv [70], and GraphConv [71] on four datasets, NCI1, NCI109, PROTEINS, and PROTEINS_full. We observe that TAGConv achieves the best results across all datasets and all four methods except for PROTEINS, while SAGEConv consistently performs the worst on NCI1, NCI109, and PROTEINS_full datasets.
Figure 5: Performance w.r.t. graph convolution backbones for different pooling methods.
Figure 6: Performance w.r.t. different pooling ratios for four pooling methods.
Parameter Analysis. Figure 6 shows the performance of four pooling methods on different datasets. From the results, we observe that as the pooling rate increases from 0.1 to 0.9, the performance would increase before saturation in most datasets. The performance is relatively stable when the ratio is around $0.7$. However, the performance on PROTEINS is limited and keeps declining when the ratio is over $0.3$. The potential reason is the lack of node attributes in PROTEINS, which validates the significance of node attributes in graph classification.
5 Conclusion
In this paper, we construct the first graph pooling benchmark that includes 15 state-of-the-art approaches and 21 different graph datasets across graph classification, graph regression and node classification. This benchmark systematically analyzes the effectiveness, robustness, and generalizability of graph pooling methods. We also make our benchmark publically available to advance the fields of graph machine learning and applications. One limitation of our benchmark is the lack of more complicated settings under label scarcity. In future works, we would extend our graph pooling benchmark to more realistic settings such as semi-supervised learning and few-shot learning.
Appendix A Related Work
A.1 Graph Classification and Graph Regression
Graphs provide an effective tool to represent interaction among different objects [72]. Graph classification [73] is a fundamental graph machine learning problem, which aims to classify each graph into its corresponding category. The majority of current works adopt message passing mechanisms [72], where each node receives information from its neighbors in a recursive manner. Then, a graph readout function is adopted to summarize all node representations into a graph-level representation for downstream classification. Graph classification has extensive applications in various domains such as molecular property prediction [74] and protein function analysis [75]. Graph regression [76] is close to graph classification which maps graph-level data into continuous vectors. Researchers usually utilize graph regression to formulate molecular property predictions [77]. Graph pooling has been an important topic in graph-level tasks [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], which generally utilize a hierarchical way to refine the graph structures [40, 41]. In this work, we generally study the performance of graph pooling on graph-level tasks and validate the effectiveness of graph pooling approaches in most cases.
A.2 Node Classification
Node classification aims to classify each node in a graph based on its attributes and relationship with the other nodes [78, 79]. Node classification has various applications in the real world, including social network analysis [80], knowledge graphs [81], bioinformatics [82] and online commerce services [83]. Graph neural networks have been widely utilized to solve the problem by learning semantics information across nodes by neighborhood propagation. Since graph pooling would reduce the number of nodes, recent works utilize a U-Net architecture [84], which involves down-sampling and up-sampling with residue connections. In this work, we systematically evaluate the performance of graph pooling on node classification and observe that graph pooling has limited improvement compared with basic graph neural network architectures.
Appendix B Details of Selected Pooling Methods
TopKPool [25]. TopKPool utilizes the attention mechanism to learn the scores of different nodes and then selects the nodes with top scores, which can learn important local portions from original graphs.
SAGPool [26]. SAGPool utilizes a different graph neural network to learn importance scores, which can guide the pooling process effectively.
ASAPool [27]. ASAPool considers the neighboring subgraphs to represent nodes and then adopts the attention mechanism to generate subgraph representations. The importance nodes are selected by a graph neural networks with local extremum information.
PANPool [28]. PANPool constructs the maximal entropy transition (MET) matrix based on graph Laplacian, which can generate importance scores for different nodes.
COPool [29]. COPool learns pooled representations from the complimentary edge and node views. The edge view comes from high-order semantics information while the node view stems from importance scores from the cut proximity matrix.
CGIPool [30]. CGIPool incorporates the mutual information optimization into graph pooling, which can enhance the graph-level relationships between the original graph and the pooled graph.
KMISPool [31]. KMISPool incorporates the Maximal k-Independent Sets (k-MIS) into graph pooling, which can detect the importance nodes in the graph with topological preserved.
GSAPool [32]. GSAPool integrates both structural and attribute information in different components. These scores from different components are then fused to guide the graph pooling process.
HGPSLPool [33]. HGPSLPool not only utilizes graph pooling to determine important nodes from the original graph, but also leverages graph structure learning to explore the topological information in the pooled graph.
AsymCheegerCutPool [34]. AsymCheegerCutPool conducts graph clustering to generate the assignment of each node according to graph total variation (GTV). Each cluster is aggregated during graph pooling in a hierarchical manner.
DiffPool [35]. DiffPool introduces a learnable soft assignment of each node during graph clustering, and then maps each cluster into the coarsened nodes in the pooling graph.
MincutPool [36]. MincutPool relaxes the classic normalized mincut problem into a continuous fashion, and then optimizes a graph neural network to achieve this. The graph clustering results are adopted to guide the graph pooling process.
DMoNPool [37]. DMoNPool introduces an objective based on modularity for graph clustering and then adds a regularization term to avoid trivial solutions during optimization. Similarly, graph clustering results are leveraged for graph pooling.
HoscPool [38]. HoscPool combines higher-order relationships in the graph with graph pooling based on motif conductance. It minimizes a relaxed motif spectral clustering objective and involves multiple motifs to learn hierarchical semantics.
JustBalancePool [39]. JustBalancePool consists of two components. On the one hand, it aims to reduces the local quadratic variation during graph clustering. On the other hand, it involves a balanced term to reduce the risk of degenerate solutions.
Appendix C Additional Experimental Details
C.1 Graph Classification
The classification model comprises three primary components: SAGEConv layers, pooling methods, and a global average pooling layer. The hidden and output channels for this model are both set to 64. Initially, the data passes through three SAGEConv layers with ReLU activation functions, followed by two pooling layers as well as a global average pooling layer. The embedding output from this global layer undergoes further processing through a linear layer with ReLU activation, with dimensions (64, 32), followed by another linear layer without any activation function but with dimensions (32, number of classes). The final output can be available after applying softmax to the embedding output. Hyperparameter tuning is not conducted due to limitations in computational resources and time. All models use the Adam optimizer with a learning rate of 0.001 and are trained for 200 epochs by minimizing the negative log-likelihood loss function. The whole data is divided into training, validation, and test sets (70%, 15%, and 15% ratios), and each trial is repeated five times with different random seeds.
C.2 Graph Regression
We use a backbone network inspired by MESPool [85] for graph regression. The model mainly consists of three GINConv layers with ReLU activation functions and BatchNorm, along with two pooling layers, followed by a global average pooling layer. All channels (both hidden and output) are set to 64. The embedding output from the global average pooling layer passes through another linear layer with ReLU activation, having dimensions (64, 32). Hyperparameter tuning is not conducted due to limitations in computational resources and time. All models use the Adam optimizer with a learning rate of 0.001 and are trained for 200 epochs by minimizing the negative log-likelihood loss function. The data is split uniformly into training, validation, and test sets in ratios of 80%, 10%, and 10% respectively. Each experiment is carried out five times using different random seeds.
C.3 Node Classification
For node classification, we utilize a U-net architecture, which we divide into a downsampling convolutional part and an upsampling convolutional part [86]. The downsampling convolutional section includes two SAGEConv layers with ReLU activation functions, with pooling applied between these layers. In the upsampling convolutional section, we use the indices saved during pooling for upsampling, restoring features to their pre-pooling size. The upsampled features are then fused with the corresponding residual features from the downsampling path, either through summation or concatenation. Finally, the fused features are processed and activated through a SAGEConv layer.
Due to constraints in computational resources and time, we do not perform hyperparameter tuning. All models employ the Adam optimizer with a learning rate set to 0.001 and are trained for 200 epochs using cross-entropy loss. All data are processed using a 5-fold cross-validation and are run on five different seeds.
Appendix D Detailed Description of Datasets
D.1 Graph Classification
Table 7 provides descriptive statistics of the selected datasets, which indicates that our chosen datasets encompass graph data of varying scales and features. This diversity establishes a robust foundation for our benchmarking research. The following are detailed descriptions of these datasets:
PROTEINS represents protein structures; nodes denote secondary structure elements (SSEs) and the edges indicate the relationships between these SSEs that are in close proximity. The primary goal is to classify each protein into different classes, i.e., a non-enzyme or an enzyme based on their amino acid sequences and topological characteristics. The dataset covers diverse protein structures [54].
PROTEINS_full is an extended version of PROTEINS, which includes the node attributes as well. The node attribute corresponds to different properties of SSEs including the polarity and the polarizability [54].
NCI1 is a collection of chemical compound graphs. Originating from the National Cancer Institute (NCI) database, each graph sample is a compound in which nodes represent atoms and edges represent the bonds between them. The dataset is binary-class labeled, indicating biological activity via compounds’ anti-cancer activity against specific cell lines [53].
NCI109 is also a collection of chemical compound graphs derived from the NCI database. Similarly, each node in the graph denotes an atom and each edge denotes a bond. The two classes in NCI109 are about compounds’ ability to inhibit or interact with the specified cancer cell line [53].
ENZYMES is developed to classify each enzyme into one of six EC (Enzyme Commission) top-level classes. Each graph at hand describes a protein in a way that nodes represent secondary structure elements (SSEs) while edges show the relationship between these SSEs either in space or sequence [56].
BZR consists of molecular graphs that represent chemical compounds. Each graph in this dataset is a chemical compound where nodes are atoms and edges are the chemical bonds between them. The classification task is focused on classifying compounds based on their biological activity with respect to binding with benzodiazepine receptors. In other words, the binary label of each graph can indicate binding or non-binding status to the benzodiazepine receptor [52].
IMDB-M is a collection of social network graphs derived from the Internet Movie Database (IMDB). Each graph is about a collaboration network from movies whereby nodes stand for actors or actresses and edges indicate that the two actors appeared in the same movie. The dataset comprises three classes reflecting the movie genres [57].
COLORS-3 is a collection of artificially generated graphs. The dataset’s primary target is to classify these graphs based on their color distribution patterns and structural properties [58].
D&D is a dataset of protein graphs. Each graph in this dataset represents a protein, with nodes corresponding to amino acids and edges corresponding to the spatial or sequential proximity between these amino acids. The primary objective of the D&D dataset is to classify proteins into one of two categories, i.e., enzymes or non-enzymes [55].
Table 7: Summary statistics of datasets for graph classification.
Datasets Graphs Classes Avg. Nodes Avg. Edges Node Attr.
PROTEINS 1,113 2 39.06 72.82 + (1)
PROTEINS_full 1,113 2 39.06 72.82 + (29)
NCI1 4,110 2 29.87 32.30 -
NCI109 4,127 2 29.68 32.13 -
ENZYMES 600 6 32.63 62.14 + (18)
BZR 405 2 35.75 38.36 + (3)
IMDB-M 1,500 3 13.00 65.94 -
COLORS-3 10,500 11 61.31 91.03 + (4)
D&D 1,178 2 284.32 715.66 -
D.2 Graph Regression
Table 8 provides an overview of the selected datasets in terms of their task numbers, compounds and other features, including splits, and metrics. A more detailed description is provided below.
QM8 is a benchmark dataset in computational chemistry, which is developed to facilitate the development and evaluation of machine learning approaches for quantum mechanical property prediction. It contains approximately 21,786 molecular structures, each characterized by their calculated properties using quantum chemistry methods, specifically focusing on electronic spectra [10, 87].
BACE is a collection of biochemical data used to evaluate computational methods for drug discovery. The dataset includes a total of 1,522 compounds, each annotated with their binding affinities, as well as molecular descriptors and fingerprints to facilitate the development and assessment of machine learning models [10, 88].
ESOL is a prominent resource in cheminformatics, which is leveraged to evaluate machine learning models on the prediction of aqueous solubility of small molecules. The dataset, derived from the work of Delaney, consists of a diverse range of chemical compounds with experimentally determined solubility values expressed in logS, where S is the solubility in mols per liter. It includes 1128 compounds, serving as a benchmark dataset for solubility prediction tasks [89, 10].
FreeSolv is a dataset containing hydration-free energies for small molecules in an aqueous solution. It comprises data for a wide range of organic molecules, providing both experimental values and calculated predictions based on molecular dynamics simulations [90, 10].
Lipophilicity is primarily utilized for studying and evaluating molecular lipophilicity. This dataset comprises 4,200 compounds sourced from the ChEMBL database, with experimentally measured partition coefficient (logD) values that reflect the distribution behavior of compounds in a water-octanol system [91, 10].
Table 8: Details of datasets for graph regression.
Datasets Data type Tasks Compounds Split Metrics
QM8 SMILES, 3D coordinates 12 21,786 Random MAE
BACE SMILES 1 1,522 Scaffold AUROC
ESOL SMILES 1 1,128 Random RMSE
FreeSolv SMILES 1 643 Random RMSE
Lipophilicity SMILES 1 4,200 Random RMSE
D.3 Node Classification
Table 9 presents descriptive statistics of the seven datasets used for node classification. It is evident that there is a significant variance in the scale of the selected datasets and each of them show different characteristics. Further background information and details about these datasets are provided below.
Cora comprises a collection of 2,708 scientific publications classified into seven distinct categories. Each publication in the dataset is represented as a node in a citation network, where edges indicate citation relationships between papers [59].
CiteSeer is a widely used citation network dataset. It comprises scientific publications categorized into six classes, with each publication represented by a 3,327-dimensional binary vector recording the presence or absence of specific words [59].
PubMed consists of scientific publications from the PubMed database, categorized into three classes based on their Medical Subject Headings (MeSH) terms. Each node is represented by a sparse bag-of-words vector derived from the content of the corresponding publication [59].
Table 9: Summary statistics of datasets for node classification.
Datasets Number of Nodes Number of Edges Number of Features Number of Classes
Cora 2,708 10,556 1,433 7
CiteSeer 3,327 9,104 3,703 6
PubMed 19,717 88,648 500 3
Cornell 183 298 1,703 5
Texas 183 325 1,703 5
Wisconsin 251 515 1,703 5
Github 37,700 578,006 0 2
Cornell, Texas, and Wisconsin consist of nodes that represent web pages and edges which denote hyperlinks between these pages. Each node has a class label which typically denotes the topic of the web page; this allows tasks including node classification and link prediction to be performed. The datasets have different sizes. In particular, Cornell and Texas have 183 nodes while Wisconsin has 251 nodes. These networks are relatively small and sparse when compared to other larger citation datasets [60].
Github includes rich node attributes representing the diverse features of developers, such as their interests, skills, and contributions to various repositories. Additionally, the edges within the network capture the interactions and collaborations among developers, which result in a multi-faceted graph structure [61].
D.4 Out-of-distribution shifts
Size shifts. For the selected datasets including D&D, PROTEINS, NCI1, and NCI109, we utilize the data provided by the previous work [66]. In this setup, the graphs with the smallest 50% of nodes are used as the training set, those with the largest 20% of nodes are used as the test set, and the remaining graphs are used as the validation set. The data can be downloaded from https://www.dropbox.com/s/38eg3twe4dd1hbt/data.zip.
Density shifts. For the selected datasets, we divide the datasets based on graph density. In particular, the 50% of graphs with the lowest density are used as the training set, the 20% with the highest density are used as the test set, and the remaining graphs are used as the validation set. After applying density shifts, the following densities are observed: for D&D, the training set density is 0.0274, the validation set density is 0.0567, and the test set density is 0.1245; for PROTEINS, the training set density is 0.1709, the validation set density is 0.4183, and the test set density is 1.0752; for NCI1, the training set density is 0.1229, the validation set density is 0.1976, and the test set density is 0.2922; for NCI109, the training set density is 0.1248, the validation set density is 0.2000, and the test set density is 0.2916.
Appendix E Additional Experiments
Table 10 presents the results of fixing the backbone as GraphConv and setting the pooling ratios of the sparse pooling method to 0.1, 0.3, 0.5, 0.7, and 0.9, respectively. The pooling ratios for dense pooling remain unchanged. As shown in Table 10, with the increase in pooling ratio, the accuracy exhibits a fluctuating upward trend. Among the sparse pooling methods, HGPSLPool and KMISPool consistently demonstrate superior performance. However, changes in the pooling ratio also result in varying stability across different methods. Furthermore, compared to binary classification tasks, the accuracy improvement with the increasing pooling ratio is more evident in multi-class classification tasks. This indicate that researchers should consider using larger pooling ratios when conducting studies on multi-class classification.
Table 11 presents the results of using GCNConv as the backbone model while varying the pooling ratio. It can be observed that the overall accuracy of GCNConv is lower than that of GraphConv. Generally, as the pooling ratio increases, classification accuracy improves and dense pooling methods are more powerful than sparse pooling methods at the cost of increased computation time. With the backbone of GCNConv, CGIPool and KMISPool perform best among the sparse pooling methods. Additionally, the impact of the pooling ratio on multi-class classification accuracy remains significantly greater than its impact on binary classification accuracy.
Table 12 presents the results of varying the pooling ratio (0.1, 0.3, 0.5, 0.7, 0.9) in node classification tasks using GraphConv as the backbone model. Overall, the change in pooling ratio has a minimal impact on the accuracy of node classification tasks. For most methods, accuracy does not improve with an increased pooling ratio but rather fluctuates. The two best-performing methods overall are TopKPool and GSAPool. Additionally, higher pooling ratios result in increased computational costs for some methods, such as HGPSLPool and ASAPool. This indicates that when performing node classification, graph pooling methods should be used more cautiously, and excessive concern over pooling ratios may be unnecessary.
Table 13 presents the results of node classification with GCNConv as the backbone model, varying the pooling ratio from 0.1, 0.3, 0.5, 0.7, to 0.9. It can be observed that, except for HGPSLPool, the performance of most methods with GCNConv is inferior to that with GraphConv. The performance of different pooling methods varies across different pooling ratios, making it difficult to identify a single pooling method that consistently outperforms others at all pooling ratios. Nevertheless, CGIPool and KMISPool still relatively demonstrate the best overall performance. Additionally, increasing the pooling ratio does not consistently improve node classification tasks. Lastly, GCNConv has higher memory consumption compared to GraphConv, particularly for ASAPool, PANPool, COPool, and HGPSLPool.
In summary, based on the aforementioned experiments, GraphConv is recommended as the backbone model over GCNConv for both graph classification and node classification tasks due to its higher accuracy and lower computational resource consumption. In terms of graph classification tasks, a larger pooling ratio generally improves results, especially for multi-class classification tasks. In contrast, the choice of pooling ratio is not very important for node classification tasks.
Table 10: Results of graph classification using GraphConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods PROTEINS NCI1 NCI109 ENZYMES BZR IMDB-M COLORS-3
Sparse Pooling, pooling ratio = 0.1
TopKPool 70.48$\pm$3.99 73.44$\pm$1.69 58.26$\pm$2.22 31.11$\pm$5.26 81.94$\pm$1.58 40.71$\pm$3.21 22.59$\pm$7.61
SAGPool 71.31$\pm$3.40 73.00$\pm$1.20 64.48$\pm$2.16 26.44$\pm$4.63 84.52$\pm$3.16 37.96$\pm$2.98 29.17$\pm$2.48
ASAPool 70.36$\pm$4.66 76.47$\pm$2.13 68.19$\pm$2.97 21.11$\pm$6.67 85.48$\pm$2.70 41.96$\pm$2.48 40.23$\pm$1.90
PANPool 68.33$\pm$1.09 77.86$\pm$0.92 65.00$\pm$1.22 30.89$\pm$4.73 86.45$\pm$3.32 39.20$\pm$2.34 21.57$\pm$0.75
COPool 70.36$\pm$1.48 75.59$\pm$0.96 66.97$\pm$2.26 39.11$\pm$4.63 81.61$\pm$1.29 35.11$\pm$2.55 28.25$\pm$2.91
CGIPool 67.86$\pm$3.63 76.21$\pm$1.71 64.84$\pm$2.83 40.00$\pm$5.26 84.52$\pm$2.99 44.98$\pm$2.49 28.27$\pm$7.25
KMISPool 75.60$\pm$1.13 76.34$\pm$0.77 73.29$\pm$1.31 35.11$\pm$9.34 85.48$\pm$5.40 38.76$\pm$1.76 57.35$\pm$2.29
GSAPool 71.67$\pm$3.03 74.42$\pm$2.36 66.19$\pm$3.58 27.56$\pm$1.91 83.87$\pm$1.02 40.71$\pm$1.72 28.32$\pm$6.78
HGPSLPool 75.95$\pm$1.10 77.21$\pm$0.85 68.03$\pm$1.82 32.00$\pm$3.74 86.77$\pm$2.77 33.69$\pm$0.81 34.80$\pm$1.43
Sparse Pooling, pooling ratio = 0.3
TopKPool 70.48$\pm$3.99 70.99$\pm$0.78 65.55$\pm$0.80 28.22$\pm$2.86 83.23$\pm$1.64 41.24$\pm$3.95 28.76$\pm$5.69
SAGPool 70.83$\pm$3.30 72.03$\pm$1.95 70.26$\pm$2.07 27.33$\pm$4.01 83.55$\pm$2.14 41.60$\pm$2.15 35.58$\pm$3.54
ASAPool 74.40$\pm$3.43 76.82$\pm$0.41 72.52$\pm$1.40 26.00$\pm$9.10 82.58$\pm$2.37 37.96$\pm$4.56 50.46$\pm$1.99
PANPool 68.33$\pm$1.52 77.70$\pm$1.92 68.23$\pm$1.79 32.44$\pm$3.25 82.90$\pm$2.99 40.89$\pm$2.71 32.55$\pm$0.99
COPool 70.71$\pm$0.69 74.26$\pm$2.05 71.16$\pm$2.07 36.22$\pm$8.96 83.23$\pm$2.62 39.02$\pm$3.40 33.16$\pm$2.87
CGIPool 68.69$\pm$4.27 76.11$\pm$1.53 68.55$\pm$2.33 29.56$\pm$6.31 80.00$\pm$2.19 38.04$\pm$3.27 39.25$\pm$5.92
KMISPool 77.02$\pm$1.53 76.95$\pm$0.65 73.48$\pm$1.42 32.22$\pm$8.40 84.84$\pm$4.96 39.02$\pm$2.49 57.35$\pm$2.29
GSAPool 70.71$\pm$2.80 74.23$\pm$1.61 70.65$\pm$1.03 28.22$\pm$7.88 84.19$\pm$2.58 38.76$\pm$2.27 35.18$\pm$4.46
HGPSLPool 75.36$\pm$1.23 77.99$\pm$1.32 71.61$\pm$1.00 30.00$\pm$7.34 84.84$\pm$3.62 38.76$\pm$2.17 50.32$\pm$1.38
Sparse Pooling, pooling ratio = 0.5
TopKPool 73.10$\pm$1.09 70.57$\pm$1.56 69.68$\pm$3.15 39.33$\pm$5.10 82.26$\pm$4.45 39.82$\pm$1.11 35.06$\pm$5.96
SAGPool 72.62$\pm$3.22 71.96$\pm$1.88 71.94$\pm$2.19 33.33$\pm$1.99 83.55$\pm$2.37 39.64$\pm$2.15 41.83$\pm$4.26
ASAPool 71.90$\pm$5.10 76.11$\pm$1.35 72.52$\pm$1.40 22.67$\pm$9.85 83.55$\pm$3.13 35.38$\pm$5.10 59.79$\pm$2.37
PANPool 73.81$\pm$1.88 77.83$\pm$1.61 70.35$\pm$1.82 41.11$\pm$4.61 86.13$\pm$2.19 39.02$\pm$3.57 41.29$\pm$1.03
COPool 73.21$\pm$1.92 74.75$\pm$2.40 72.87$\pm$2.17 38.22$\pm$2.06 82.90$\pm$1.29 39.29$\pm$2.36 38.43$\pm$2.52
CGIPool 74.29$\pm$1.15 76.73$\pm$2.37 71.90$\pm$0.89 28.67$\pm$5.51 84.84$\pm$5.26 37.16$\pm$3.33 41.32$\pm$4.06
KMISPool 75.95$\pm$0.97 76.30$\pm$1.06 74.03$\pm$0.54 33.11$\pm$8.36 84.84$\pm$3.47 38.76$\pm$2.32 57.37$\pm$1.94
GSAPool 74.05$\pm$2.98 74.04$\pm$1.98 72.16$\pm$2.73 33.11$\pm$2.57 82.90$\pm$2.62 41.33$\pm$0.80 38.22$\pm$6.40
HGPSLPool 74.64$\pm$1.58 77.41$\pm$1.39 73.61$\pm$1.08 29.11$\pm$3.94 85.16$\pm$2.58 41.60$\pm$1.70 62.30$\pm$2.55
Sparse Pooling, pooling ratio = 0.7
TopKPool 73.81$\pm$1.25 74.88$\pm$1.04 72.94$\pm$1.04 37.78$\pm$4.61 81.29$\pm$3.16 40.44$\pm$2.77 42.77$\pm$4.11
SAGPool 72.86$\pm$1.90 74.55$\pm$1.41 73.19$\pm$2.33 41.11$\pm$5.07 84.84$\pm$3.32 39.82$\pm$2.97 41.83$\pm$4.26
ASAPool 74.05$\pm$2.25 79.06$\pm$0.68 74.87$\pm$1.26 18.00$\pm$1.30 85.48$\pm$3.68 34.40$\pm$4.44 59.79$\pm$2.37
PANPool 74.64$\pm$1.44 77.05$\pm$0.39 74.35$\pm$0.87 42.00$\pm$6.02 85.16$\pm$3.13 41.33$\pm$1.76 41.29$\pm$1.03
COPool 74.17$\pm$1.23 75.20$\pm$0.90 73.97$\pm$2.58 39.56$\pm$5.60 81.29$\pm$2.19 39.64$\pm$4.14 38.43$\pm$2.52
CGIPool 75.83$\pm$2.51 78.38$\pm$1.70 74.71$\pm$0.97 28.67$\pm$5.51 84.84$\pm$5.26 37.16$\pm$3.33 41.32$\pm$4.06
KMISPool 76.67$\pm$1.02 76.66$\pm$0.42 73.06$\pm$1.45 34.00$\pm$6.76 85.16$\pm$2.37 37.78$\pm$5.36 41.32$\pm$4.06
GSAPool 72.14$\pm$1.52 75.49$\pm$1.14 72.77$\pm$1.39 38.00$\pm$5.77 83.55$\pm$2.37 41.87$\pm$3.17 38.22$\pm$6.40
HGPSLPool 75.00$\pm$2.23 78.28$\pm$1.00 73.84$\pm$0.55 24.67$\pm$4.52 86.45$\pm$1.94 41.33$\pm$0.97 62.30$\pm$2.55
Sparse Pooling, pooling ratio = 0.9
TopKPool 73.21$\pm$1.92 75.82$\pm$1.45 73.19$\pm$1.73 39.56$\pm$3.27 82.58$\pm$3.13 41.16$\pm$1.58 47.87$\pm$2.36
SAGPool 72.74$\pm$0.95 76.21$\pm$1.96 74.45$\pm$0.56 39.33$\pm$2.59 81.61$\pm$2.19 41.24$\pm$3.48 49.66$\pm$2.52
ASAPool 73.69$\pm$1.48 79.09$\pm$0.45 77.16$\pm$1.42 19.33$\pm$1.51 86.77$\pm$1.88 34.84$\pm$4.11 72.74$\pm$1.21
PANPool 74.88$\pm$1.02 77.99$\pm$0.72 76.06$\pm$1.05 36.44$\pm$4.73 83.55$\pm$1.21 39.11$\pm$1.93 64.61$\pm$2.22
COPool 74.52$\pm$1.38 77.08$\pm$2.42 75.03$\pm$1.02 36.44$\pm$2.85 81.94$\pm$1.21 35.73$\pm$2.31 43.67$\pm$1.68
CGIPool 72.62$\pm$1.92 79.29$\pm$1.19 75.29$\pm$0.98 31.33$\pm$4.89 82.90$\pm$0.79 34.84$\pm$3.38 44.35$\pm$1.25
KMISPool 74.29$\pm$3.14 75.85$\pm$0.93 74.03$\pm$1.30 31.11$\pm$8.66 84.84$\pm$4.85 38.84$\pm$1.91 57.35$\pm$2.29
GSAPool 72.38$\pm$1.83 76.79$\pm$1.70 73.42$\pm$0.76 33.33$\pm$2.33 83.23$\pm$2.99 40.62$\pm$3.26 52.43$\pm$2.50
HGPSLPool 75.60$\pm$1.46 77.21$\pm$1.53 74.84$\pm$0.58 27.11$\pm$5.10 86.45$\pm$3.62 38.22$\pm$3.85 87.12$\pm$1.82
Dense Pooling
AsymCheegerCutPool 75.12$\pm$1.05 73.87$\pm$0.89 72.74$\pm$1.03 31.78$\pm$3.19 83.55$\pm$2.77 43.02$\pm$3.46 OOT
DiffPool 75.00$\pm$0.77 78.38$\pm$0.43 79.81$\pm$0.96 35.33$\pm$6.38 84.19$\pm$1.88 47.73$\pm$1.72 OOT
MincutPool 73.54$\pm$5.56 78.90$\pm$0.99 77.84$\pm$1.78 38.00$\pm$2.15 80.65$\pm$2.70 39.02$\pm$5.07 OOT
DMoNPool 75.49$\pm$1.94 74.46$\pm$0.79 72.71$\pm$0.55 41.33$\pm$3.81 82.58$\pm$1.88 51.20$\pm$2.44 OOT
HoscPool 73.05$\pm$2.06 78.61$\pm$0.69 77.71$\pm$1.84 41.11$\pm$4.39 79.68$\pm$6.00 52.80$\pm$2.54 OOT
JustBalancePool 74.51$\pm$2.84 78.61$\pm$0.69 77.71$\pm$1.84 38.00$\pm$4.35 79.68$\pm$6.00 52.09$\pm$1.42 OOT
Table 11: Results of graph classification using GCNConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods PROTEINS NCI1 NCI109 ENZYMES BZR IMDB-M COLORS-3
Sparse Pooling, pooling ratio = 0.1
TopKPool 69.52$\pm$1.09 58.83$\pm$2.47 60.52$\pm$4.30 29.33$\pm$6.46 83.23$\pm$0.79 33.60$\pm$3.69 18.31$\pm$4.39
SAGPool 70.36$\pm$4.48 63.66$\pm$3.17 63.19$\pm$1.34 26.22$\pm$5.65 84.19$\pm$1.21 39.20$\pm$4.52 32.24$\pm$1.06
ASAPool 76.90$\pm$2.01 64.64$\pm$1.96 65.00$\pm$1.56 25.56$\pm$8.04 81.94$\pm$2.77 32.53$\pm$2.17 34.76$\pm$0.77
PANPool 64.88$\pm$3.63 65.90$\pm$1.09 66.48$\pm$1.21 39.56$\pm$3.49 84.19$\pm$1.21 31.47$\pm$2.00 23.34$\pm$1.36
COPool 67.98$\pm$2.39 61.33$\pm$1.99 60.87$\pm$2.57 36.22$\pm$4.80 83.23$\pm$0.79 40.53$\pm$0.76 31.21$\pm$1.54
CGIPool 70.83$\pm$3.30 66.74$\pm$4.68 66.19$\pm$3.61 35.11$\pm$3.62 85.48$\pm$1.44 40.98$\pm$1.74 36.62$\pm$2.99
KMISPool 75.95$\pm$5.36 74.26$\pm$1.32 70.81$\pm$1.43 23.56$\pm$2.15 86.13$\pm$1.29 30.84$\pm$1.42 57.56$\pm$1.53
GSAPool 68.81$\pm$1.71 60.68$\pm$1.97 63.39$\pm$3.30 30.22$\pm$4.30 80.32$\pm$2.58 33.16$\pm$4.49 23.05$\pm$6.52
HGPSLPool 71.43$\pm$0.75 64.21$\pm$1.73 64.87$\pm$1.65 31.33$\pm$2.15 84.19$\pm$2.58 31.11$\pm$1.35 22.84$\pm$0.74
Sparse Pooling, pooling ratio = 0.3
TopKPool 69.88$\pm$1.10 64.89$\pm$1.20 63.13$\pm$2.24 35.56$\pm$7.73 84.19$\pm$1.58 32.98$\pm$3.38 30.84$\pm$6.31
SAGPool 70.95$\pm$1.52 69.69$\pm$1.66 67.94$\pm$1.58 29.78$\pm$6.34 83.23$\pm$3.76 40.80$\pm$2.77 29.59$\pm$4.17
ASAPool 75.12$\pm$1.90 69.79$\pm$0.77 66.87$\pm$0.84 23.56$\pm$6.02 82.58$\pm$1.88 31.82$\pm$1.90 45.95$\pm$2.82
PANPool 67.74$\pm$1.15 69.92$\pm$0.90 66.16$\pm$1.25 34.22$\pm$2.37 81.94$\pm$1.21 33.51$\pm$1.98 34.99$\pm$2.01
COPool 72.38$\pm$3.03 69.50$\pm$1.27 65.68$\pm$1.64 35.56$\pm$4.50 81.61$\pm$2.62 38.13$\pm$3.49 37.35$\pm$2.24
CGIPool 73.57$\pm$2.28 73.84$\pm$2.73 69.32$\pm$2.52 34.44$\pm$2.98 85.16$\pm$2.37 39.29$\pm$3.11 41.76$\pm$2.47
KMISPool 76.90$\pm$6.38 74.94$\pm$0.36 71.48$\pm$1.18 20.89$\pm$3.68 85.16$\pm$2.37 31.73$\pm$1.48 57.55$\pm$2.67
GSAPool 68.33$\pm$2.14 67.59$\pm$2.22 67.94$\pm$2.34 31.56$\pm$5.65 84.52$\pm$2.19 30.76$\pm$2.15 36.50$\pm$4.39
HGPSLPool 69.40$\pm$1.90 69.37$\pm$1.08 67.71$\pm$1.72 34.22$\pm$2.47 83.87$\pm$1.44 32.62$\pm$1.60 37.51$\pm$1.02
Sparse Pooling, pooling ratio = 0.5
TopKPool 71.19$\pm$2.25 66.61$\pm$1.40 67.71$\pm$2.79 34.67$\pm$7.32 84.52$\pm$1.29 34.04$\pm$1.28 38.33$\pm$5.81
SAGPool 73.57$\pm$1.39 70.50$\pm$1.74 69.19$\pm$2.32 33.56$\pm$2.57 84.52$\pm$1.29 37.24$\pm$1.87 39.72$\pm$2.19
ASAPool 75.71$\pm$1.27 73.29$\pm$1.07 69.26$\pm$1.18 22.00$\pm$4.58 83.23$\pm$2.81 34.76$\pm$2.07 52.72$\pm$1.83
PANPool 70.60$\pm$1.67 73.78$\pm$1.12 70.81$\pm$1.33 31.56$\pm$2.69 83.23$\pm$3.32 34.58$\pm$5.16 43.31$\pm$1.03
COPool 72.62$\pm$2.41 70.89$\pm$1.21 68.61$\pm$1.66 31.33$\pm$4.41 82.90$\pm$2.41 36.44$\pm$1.26 39.77$\pm$2.49
CGIPool 73.57$\pm$1.49 75.72$\pm$1.65 72.90$\pm$0.64 34.00$\pm$3.89 84.19$\pm$2.58 39.82$\pm$1.28 43.66$\pm$3.08
KMISPool 72.26$\pm$5.07 76.40$\pm$0.73 71.65$\pm$1.05 18.44$\pm$3.49 85.48$\pm$2.70 31.56$\pm$1.80 57.88$\pm$1.55
GSAPool 72.14$\pm$1.09 68.65$\pm$1.93 67.84$\pm$1.71 33.33$\pm$4.04 83.55$\pm$1.58 33.33$\pm$2.37 46.14$\pm$3.13
HGPSLPool 71.67$\pm$2.08 71.83$\pm$1.05 69.26$\pm$1.46 30.44$\pm$4.53 83.87$\pm$1.44 31.20$\pm$0.18 50.30$\pm$0.55
Sparse Pooling, pooling ratio = 0.7
TopKPool 70.36$\pm$0.69 68.85$\pm$2.84 68.45$\pm$2.03 33.11$\pm$1.91 84.19$\pm$2.58 32.00$\pm$1.66 45.43$\pm$7.52
SAGPool 68.69$\pm$1.58 72.84$\pm$1.07 70.48$\pm$1.28 34.44$\pm$3.37 83.87$\pm$2.70 38.22$\pm$2.12 45.90$\pm$1.52
ASAPool 75.00$\pm$1.96 72.67$\pm$0.94 70.45$\pm$0.84 25.56$\pm$6.32 86.13$\pm$2.41 31.11$\pm$2.05 58.02$\pm$2.00
PANPool 69.52$\pm$1.43 71.83$\pm$0.87 69.84$\pm$1.07 36.22$\pm$4.69 83.87$\pm$0.00 32.27$\pm$0.96 52.69$\pm$2.90
COPool 73.21$\pm$1.06 72.61$\pm$0.97 70.65$\pm$1.25 40.89$\pm$6.06 84.19$\pm$2.77 40.80$\pm$1.42 41.81$\pm$2.80
CGIPool 72.26$\pm$1.04 73.39$\pm$0.84 71.84$\pm$2.04 36.00$\pm$4.75 82.58$\pm$1.58 40.62$\pm$2.04 44.51$\pm$3.79
KMISPool 76.79$\pm$5.66 74.98$\pm$0.99 71.16$\pm$1.58 23.33$\pm$2.98 85.48$\pm$1.02 31.56$\pm$0.56 58.68$\pm$1.55
GSAPool 71.90$\pm$3.11 70.53$\pm$1.18 68.32$\pm$1.04 36.89$\pm$1.09 83.87$\pm$0.00 32.27$\pm$1.90 49.71$\pm$8.30
HGPSLPool 72.02$\pm$1.73 71.64$\pm$0.72 68.97$\pm$1.43 33.78$\pm$3.03 84.52$\pm$1.29 32.62$\pm$0.92 57.67$\pm$2.07
Sparse Pooling, pooling ratio = 0.9
TopKPool 70.83$\pm$1.25 70.34$\pm$1.80 69.65$\pm$1.61 38.22$\pm$2.78 82.90$\pm$1.29 32.44$\pm$1.32 54.48$\pm$4.48
SAGPool 74.64$\pm$1.53 73.10$\pm$1.21 71.29$\pm$0.82 34.67$\pm$4.52 82.90$\pm$1.29 40.80$\pm$1.39 53.18$\pm$1.71
ASAPool 73.69$\pm$1.48 73.48$\pm$1.03 70.06$\pm$1.06 24.44$\pm$7.44 85.16$\pm$2.37 30.93$\pm$0.87 58.50$\pm$0.68
PANPool 70.48$\pm$0.81 72.77$\pm$1.09 70.84$\pm$1.23 40.67$\pm$4.48 82.90$\pm$1.29 32.98$\pm$1.98 61.86$\pm$1.15
COPool 75.24$\pm$2.46 74.10$\pm$1.06 71.35$\pm$1.05 39.56$\pm$4.19 80.32$\pm$2.14 39.56$\pm$2.47 45.03$\pm$1.19
CGIPool 70.83$\pm$1.46 75.49$\pm$1.11 73.81$\pm$0.42 38.44$\pm$1.51 83.23$\pm$3.62 41.07$\pm$2.35 48.91$\pm$4.14
KMISPool 76.79$\pm$5.73 74.46$\pm$0.98 70.06$\pm$1.66 21.56$\pm$1.33 85.48$\pm$2.28 30.04$\pm$1.81 58.82$\pm$1.21
GSAPool 70.95$\pm$0.69 71.12$\pm$1.33 70.65$\pm$1.45 40.89$\pm$5.04 85.16$\pm$1.88 32.98$\pm$0.99 57.77$\pm$7.03
HGPSLPool 70.95$\pm$1.66 72.22$\pm$0.42 70.35$\pm$1.31 31.11$\pm$3.98 83.55$\pm$0.65 33.07$\pm$1.83 73.46$\pm$1.70
Dense Pooling
AsymCheegerCutPool 77.07$\pm$0.91 73.61$\pm$0.57 73.42$\pm$0.89 45.33$\pm$3.25 73.87$\pm$2.58 43.02$\pm$3.46 OOT
DiffPool 75.24$\pm$0.62 74.98$\pm$1.37 73.84$\pm$1.99 43.78$\pm$6.19 78.06$\pm$4.52 47.73$\pm$1.72 OOT
MincutPool 76.22$\pm$1.22 68.72$\pm$6.67 73.35$\pm$1.28 38.89$\pm$8.78 81.61$\pm$1.29 39.02$\pm$5.07 OOT
DMoNPool 76.71$\pm$0.24 72.32$\pm$0.39 73.42$\pm$0.48 42.67$\pm$5.29 76.77$\pm$4.03 51.20$\pm$2.44 OOT
HoscPool 73.90$\pm$0.60 74.98$\pm$1.34 73.94$\pm$0.77 40.89$\pm$3.25 78.06$\pm$3.62 52.80$\pm$2.54 OOT
JustBalancePool 74.39$\pm$1.39 74.98$\pm$0.85 74.26$\pm$0.48 42.44$\pm$2.76 78.06$\pm$3.62 52.09$\pm$1.42 OOT
Table 12: Results of node classification using GraphConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods Cora Citeseer Pubmed Cornell Texas Wisconsin Github
Sparse Pooling, pooling ratio = 0.1
TopKPool 86.70$\pm$0.90 74.85$\pm$0.96 86.52$\pm$0.67 81.82$\pm$3.45 74.91$\pm$7.66 73.68$\pm$5.20 82.73$\pm$1.10
SAGPool 87.00$\pm$1.10 74.89$\pm$0.63 85.28$\pm$1.02 79.64$\pm$2.41 74.91$\pm$6.94 72.89$\pm$4.13 83.65$\pm$0.45
ASAPool 84.70$\pm$1.90 74.53$\pm$0.99 70.14$\pm$2.74 82.18$\pm$3.71 77.45$\pm$5.22 73.16$\pm$4.29 OOM
PANPool 81.90$\pm$1.10 73.95$\pm$1.13 72.47$\pm$2.24 78.18$\pm$3.25 76.0$\pm$7.66 73.95$\pm$3.05 OOM
COPool 85.70$\pm$0.40 74.71$\pm$0.99 78.63$\pm$1.81 78.91$\pm$1.45 74.18$\pm$7.75 72.37$\pm$3.22 OOM
CGIPool 86.50$\pm$0.90 74.53$\pm$0.80 81.96$\pm$4.51 80.00$\pm$4.45 75.27$\pm$5.22 72.11$\pm$3.05 79.85$\pm$4.62
KMISPool 77.00$\pm$3.00 74.57$\pm$1.10 64.21$\pm$4.07 80.73$\pm$3.17 75.27$\pm$5.70 73.16$\pm$4.60 83.43$\pm$0.39
GSAPool 87.00$\pm$0.60 74.33$\pm$1.10 85.88$\pm$0.67 80.00$\pm$4.45 76.0$\pm$8.24 73.42$\pm$3.47 83.60$\pm$0.16
HGPSLPool 29.40$\pm$2.40 37.12$\pm$3.31 45.04$\pm$1.56 77.45$\pm$4.96 76.00$\pm$7.58 70.26$\pm$5.30 OOM
Sparse Pooling, pooling ratio = 0.3
TopKPool 86.99$\pm$1.01 75.28$\pm$1.24 86.35$\pm$0.94 80.36$\pm$4.05 75.27$\pm$8.50 73.68$\pm$4.32 82.93$\pm$0.78
SAGPool 87.23$\pm$0.99 74.63$\pm$1.32 85.16$\pm$1.10 81.09$\pm$3.74 75.64$\pm$7.68 72.37$\pm$3.72 83.24$\pm$0.52
ASAPool 77.96$\pm$5.96 74.17$\pm$0.81 72.09$\pm$1.99 79.64$\pm$2.91 71.27$\pm$6.84 67.37$\pm$5.73 OOM
PANPool 44.21$\pm$2.12 64.34$\pm$2.51 68.23$\pm$1.79 78.91$\pm$2.47 74.91$\pm$6.34 73.42$\pm$2.26 OOM
COPool 84.45$\pm$0.96 74.47$\pm$1.08 77.52$\pm$1.76 77.82$\pm$3.13 72.00$\pm$8.65 70.79$\pm$4.19 OOM
CGIPool 86.42$\pm$0.82 74.27$\pm$0.69 82.96$\pm$5.22 81.82$\pm$4.30 76.00$\pm$6.44 72.11$\pm$2.68 79.92$\pm$4.70
KMISPool 77.32$\pm$2.42 74.55$\pm$1.24 65.45$\pm$3.46 81.09$\pm$3.74 75.64$\pm$5.59 72.89$\pm$4.53 83.60$\pm$0.37
GSAPool 87.06$\pm$0.61 74.31$\pm$1.11 85.58$\pm$1.14 80.36$\pm$3.71 76.36$\pm$4.74 72.37$\pm$3.99 82.39$\pm$1.44
HGPSLPool 29.47$\pm$1.21 28.51$\pm$1.59 44.10$\pm$0.91 58.55$\pm$3.71 62.18$\pm$4.66 51.32$\pm$3.99 OOM
Sparse Pooling, pooling ratio = 0.5
TopKPool 86.74$\pm$0.68 75.12$\pm$1.23 86.34$\pm$0.97 81.45$\pm$5.06 72.73$\pm$8.76 76.05$\pm$4.28 82.14$\pm$2.23
SAGPool 86.86$\pm$1.27 74.51$\pm$1.48 85.44$\pm$1.24 80.36$\pm$4.36 73.45$\pm$5.34 73.42$\pm$2.93 79.74$\pm$3.23
ASAPool 70.97$\pm$7.63 73.11$\pm$1.55 69.37$\pm$4.84 72.00$\pm$6.76 65.82$\pm$4.93 62.37$\pm$6.64 OOM
PANPool 47.87$\pm$2.66 55.4$\pm$2.66 67.86$\pm$1.86 78.55$\pm$2.91 76.00$\pm$8.32 71.32$\pm$5.22 OOM
COPool 83.86$\pm$0.94 74.43$\pm$1.09 77.86$\pm$1.88 76.36$\pm$3.45 69.82$\pm$8.02 67.89$\pm$3.59 OOM
CGIPool 86.49$\pm$0.90 74.05$\pm$0.93 82.80$\pm$3.99 82.55$\pm$4.96 74.91$\pm$7.58 72.63$\pm$3.85 81.28$\pm$3.46
KMISPool 76.92$\pm$2.52 74.85$\pm$1.32 64.51$\pm$4.13 80.73$\pm$2.47 76.00$\pm$5.06 73.42$\pm$3.94 83.46$\pm$0.21
GSAPool 86.96$\pm$0.83 74.51$\pm$1.00 85.90$\pm$1.21 81.45$\pm$3.88 78.55$\pm$5.56 73.95$\pm$2.93 80.18$\pm$2.73
HGPSLPool 30.68$\pm$1.26 27.63$\pm$0.88 OOM 55.27$\pm$5.82 58.55$\pm$4.80 50.53$\pm$2.29 OOM
Sparse Pooling, pooling ratio = 0.7
TopKPool 86.86$\pm$0.93 74.93$\pm$1.34 86.09$\pm$0.99 81.09$\pm$4.08 76.00$\pm$6.65 76.58$\pm$5.09 82.71$\pm$1.00
SAGPool 86.17$\pm$1.50 74.31$\pm$1.03 85.16$\pm$1.17 79.27$\pm$4.54 76.36$\pm$4.74 73.42$\pm$3.67 80.78$\pm$1.57
ASAPool 69.69$\pm$6.18 71.51$\pm$3.63 69.34$\pm$2.61 64.36$\pm$5.34 65.82$\pm$3.13 61.84$\pm$3.90 OOM
PANPool 51.56$\pm$2.50 53.45$\pm$0.74 68.01$\pm$2.01 85.16$\pm$3.13 74.55$\pm$6.8 68.42$\pm$5.71 OOM
COPool 84.72$\pm$1.68 74.23$\pm$1.03 78.84$\pm$1.49 75.64$\pm$4.69 68.00$\pm$6.86 62.89$\pm$3.76 OOM
CGIPool 85.46$\pm$2.59 73.97$\pm$1.05 81.07$\pm$2.62 80.73$\pm$3.92 76.36$\pm$5.14 72.11$\pm$3.85 80.36$\pm$3.34
KMISPool 76.09$\pm$3.12 74.85$\pm$1.45 64.89$\pm$3.73 80.73$\pm$4.08 75.64$\pm$5.47 73.95$\pm$4.11 83.41$\pm$0.30
GSAPool 86.45$\pm$0.82 74.33$\pm$1.16 86.10$\pm$1.05 81.09$\pm$3.37 77.45$\pm$6.46 73.95$\pm$2.11 80.40$\pm$4.29
HGPSLPool 30.43$\pm$0.99 27.53$\pm$0.73 OOM 53.82$\pm$2.95 59.64$\pm$3.71 45.53$\pm$2.44 OOM
Sparse Pooling, pooling ratio = 0.9
TopKPool 86.69$\pm$0.92 74.69$\pm$1.03 85.49$\pm$0.67 81.09$\pm$3.92 77.82$\pm$9.01 75.00$\pm$3.43 82.97$\pm$0.58
SAGPool 85.36$\pm$2.02 74.15$\pm$1.29 83.60$\pm$0.70 80.00$\pm$3.25 77.09$\pm$5.34 72.37$\pm$4.71 79.64$\pm$1.98
ASAPool 63.44$\pm$7.05 65.63$\pm$2.88 OOM 61.82$\pm$3.45 62.91$\pm$4.39 59.47$\pm$5.85 OOM
PANPool 52.25$\pm$2.38 52.15$\pm$2.00 70.64$\pm$3.47 79.27$\pm$2.95 71.64$\pm$5.22 67.37$\pm$5.09 OOM
COPool 85.29$\pm$1.34 74.25$\pm$0.75 79.32$\pm$1.72 76.73$\pm$2.12 68.36$\pm$6.57 65.79$\pm$4.71 OOM
CGIPool 86.49$\pm$0.87 73.05$\pm$1.92 79.63$\pm$1.74 81.45$\pm$3.71 75.27$\pm$7.59 73.95$\pm$3.67 81.21$\pm$3.57
KMISPool 76.43$\pm$2.34 74.73$\pm$1.23 64.93$\pm$3.64 80.36$\pm$3.53 76.36$\pm$4.88 72.89$\pm$4.75 83.33$\pm$0.31
GSAPool 84.94$\pm$1.63 74.49$\pm$0.95 83.72$\pm$1.42 82.18$\pm$3.53 77.45$\pm$5.09 73.16$\pm$3.96 79.66$\pm$2.51
HGPSLPool 30.60$\pm$1.28 27.87$\pm$0.52 OOM 57.82$\pm$3.13 64.0$\pm$3.53 51.58$\pm$1.53 OOM
Table 13: Results of node classification using GCNConv as the backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
Methods Cora Citeseer Pubmed Cornell Texas Wisconsin Github
Sparse Pooling, pooling ratio = 0.1
TopKPool 88.91$\pm$0.93 77.56$\pm$0.85 85.73$\pm$0.34 48.00$\pm$0.89 54.18$\pm$4.8 50.79$\pm$3.49 86.95$\pm$0.20
SAGPool 89.18$\pm$0.65 77.56$\pm$0.81 85.83$\pm$0.23 48.73$\pm$3.13 55.64$\pm$2.95 51.32$\pm$3.53 86.83$\pm$0.20
ASAPool 89.03$\pm$0.84 77.76$\pm$1.01 85.74$\pm$0.18 49.82$\pm$1.45 54.55$\pm$4.74 50.53$\pm$3.96 OOM
PANPool 89.05$\pm$0.96 77.20$\pm$0.98 85.88$\pm$0.11 49.82$\pm$2.18 56.00$\pm$4.80 50.53$\pm$3.18 OOM
COPool 89.00$\pm$0.70 77.24$\pm$0.90 85.27$\pm$0.27 47.64$\pm$2.12 54.91$\pm$5.32 51.84$\pm$0.64 86.68$\pm$0.20
CGIPool 89.15$\pm$0.84 77.40$\pm$0.81 85.80$\pm$0.13 49.45$\pm$1.78 54.55$\pm$4.15 50.79$\pm$3.78 86.88$\pm$0.22
KMISPool 89.05$\pm$0.93 77.60$\pm$0.81 85.92$\pm$0.07 48.73$\pm$1.36 54.18$\pm$3.88 52.89$\pm$2.11 86.88$\pm$0.24
GSAPool 89.05$\pm$0.77 77.16$\pm$0.92 85.70$\pm$0.19 49.09$\pm$1.15 54.18$\pm$3.88 51.58$\pm$3.85 87.12$\pm$0.09
HGPSLPool 89.00$\pm$0.64 77.84$\pm$0.67 86.00$\pm$0.16 48.73$\pm$2.12 54.91$\pm$4.21 51.58$\pm$1.75 OOM
Sparse Pooling, pooling ratio = 0.3
TopKPool 88.63$\pm$1.05 77.12$\pm$1.16 85.53$\pm$0.35 49.09$\pm$2.57 54.18$\pm$4.80 51.32$\pm$3.00 86.89$\pm$0.27
SAGPool 88.81$\pm$0.73 77.22$\pm$0.94 85.68$\pm$0.54 49.82$\pm$2.95 55.64$\pm$2.95 51.32$\pm$3.53 86.85$\pm$0.21
ASAPool 89.10$\pm$0.86 77.40$\pm$1.13 OOM 79.64$\pm$2.91 54.55$\pm$5.14 50.53$\pm$3.87 OOM
PANPool 88.98$\pm$0.70 77.16$\pm$0.99 85.64$\pm$0.15 49.82$\pm$2.47 56.36$\pm$5.14 50.00$\pm$2.04 OOM
COPool 88.73$\pm$0.84 77.26$\pm$0.89 84.70$\pm$0.24 48.73$\pm$1.78 56.00$\pm$5.80 51.05$\pm$1.93 86.61$\pm$0.29
CGIPool 89.15$\pm$0.84 77.14$\pm$0.80 85.92$\pm$0.66 49.82$\pm$1.85 54.55$\pm$4.15 51.05$\pm$3.85 86.84$\pm$0.12
KMISPool 89.08$\pm$0.86 77.59$\pm$0.65 85.93$\pm$0.06 48.73$\pm$1.36 54.18$\pm$3.88 52.89$\pm$2.11 86.86$\pm$0.25
GSAPool 88.91$\pm$1.20 77.12$\pm$1.16 86.00$\pm$0.42 49.09$\pm$1.15 54.18$\pm$3.88 51.58$\pm$3.85 86.95$\pm$0.28
HGPSLPool 89.03$\pm$0.84 77.80$\pm$0.71 OOM 48.36$\pm$1.85 55.27$\pm$3.37 51.58$\pm$2.26 OOM
Sparse Pooling, pooling ratio = 0.5
TopKPool 88.81$\pm$1.07 77.08$\pm$1.10 85.60$\pm$0.06 47.30$\pm$1.80 54.18$\pm$4.80 50.53$\pm$3.39 86.94$\pm$0.31
SAGPool 88.98$\pm$0.67 77.10$\pm$0.82 86.07$\pm$0.70 81.09$\pm$3.74 55.64$\pm$2.72 51.32$\pm$3.53 86.88$\pm$0.25
ASAPool 88.86$\pm$0.96 77.18$\pm$1.02 OOM 72.00$\pm$6.76 54.18$\pm$3.88 50.79$\pm$3.49 OOM
PANPool 88.95$\pm$0.73 77.02$\pm$0.90 OOM 78.55$\pm$2.91 56.36$\pm$5.14 49.74$\pm$1.93 OOM
COPool 88.56$\pm$1.02 76.78$\pm$0.98 OOM 76.36$\pm$3.45 56.36$\pm$5.98 52.37$\pm$2.68 86.57$\pm$0.29
CGIPool 89.15$\pm$0.84 77.10$\pm$0.83 85.80$\pm$0.78 81.82$\pm$4.30 54.55$\pm$4.15 50.79$\pm$3.49 86.86$\pm$0.08
KMISPool 89.15$\pm$0.93 77.58$\pm$0.62 85.88$\pm$0.14 81.09$\pm$3.74 54.18$\pm$3.88 52.89$\pm$2.11 86.89$\pm$0.29
GSAPool 88.98$\pm$1.13 76.96$\pm$0.93 86.04$\pm$0.53 80.36$\pm$3.71 54.18$\pm$3.88 51.32$\pm$3.81 86.94$\pm$0.27
HGPSLPool 89.08$\pm$0.90 77.80$\pm$0.96 OOM 58.55$\pm$3.71 55.27$\pm$4.39 50.53$\pm$1.78 OOM
Sparse Pooling, pooling ratio = 0.7
TopKPool 88.44$\pm$1.09 77.08$\pm$0.98 86.04$\pm$0.41 47.30$\pm$1.80 49.09$\pm$1.15 51.05$\pm$3.85 86.76$\pm$0.33
SAGPool 88.91$\pm$0.62 77.08$\pm$0.80 85.94$\pm$0.76 81.09$\pm$3.74 50.55$\pm$1.36 51.05$\pm$3.47 86.99$\pm$0.22
ASAPool 88.76$\pm$0.72 76.74$\pm$1.05 OOM 79.64$\pm$2.91 50.55$\pm$1.36 49.74$\pm$4.11 OOM
PANPool 88.59$\pm$1.01 76.98$\pm$1.03 OOM 78.91$\pm$2.47 50.91$\pm$1.15 49.74$\pm$1.93 OOM
COPool 88.27$\pm$1.09 76.72$\pm$1.07 OOM 77.82$\pm$3.13 50.18$\pm$3.74 51.58$\pm$2.93 86.57$\pm$0.28
CGIPool 89.15$\pm$0.84 77.02$\pm$0.65 85.50$\pm$0.69 81.82$\pm$4.30 49.82$\pm$1.45 50.26$\pm$3.37 86.77$\pm$0.14
KMISPool 89.15$\pm$0.91 77.52$\pm$0.65 85.86$\pm$0.10 81.09$\pm$3.74 50.55$\pm$1.36 52.89$\pm$2.11 86.89$\pm$0.26
GSAPool 88.66$\pm$0.98 77.1$\pm$0.72 86.21$\pm$0.73 80.36$\pm$3.71 49.82$\pm$2.72 51.84$\pm$4.29 86.83$\pm$0.19
HGPSLPool 89.03$\pm$0.87 77.74$\pm$0.85 OOM 58.55$\pm$3.71 49.82$\pm$2.47 50.26$\pm$2.26 OOM
Sparse Pooling, pooling ratio = 0.9
TopKPool 88.44$\pm$0.91 77.08$\pm$0.75 86.13$\pm$0.34 47.30$\pm$1.80 48.73$\pm$2.12 51.58$\pm$3.05 86.80$\pm$0.25
SAGPool 88.83$\pm$0.97 77.08$\pm$0.81 85.89$\pm$0.63 81.09$\pm$3.74 50.55$\pm$1.36 50.79$\pm$3.49 86.87$\pm$0.17
ASAPool 88.98$\pm$0.74 76.76$\pm$0.87 OOM 79.64$\pm$2.91 50.55$\pm$1.36 50.00$\pm$4.32 OOM
PANPool 88.76$\pm$1.11 76.70$\pm$0.92 OOM 78.91$\pm$2.47 51.27$\pm$2.41 49.47$\pm$1.97 OOM
COPool 88.54$\pm$1.09 76.72$\pm$1.41 OOM 77.82$\pm$3.13 49.45$\pm$3.13 51.84$\pm$2.58 86.54$\pm$0.28
CGIPool 89.15$\pm$0.84 77.10$\pm$0.54 84.95$\pm$0.39 81.82$\pm$4.30 50.55$\pm$2.12 50.79$\pm$3.78 86.65$\pm$0.25
KMISPool 89.13$\pm$0.90 77.64$\pm$0.58 85.87$\pm$0.09 81.09$\pm$3.74 50.18$\pm$0.89 52.89$\pm$2.11 86.88$\pm$0.28
GSAPool 88.93$\pm$0.88 76.94$\pm$0.74 86.07$\pm$0.62 80.36$\pm$3.71 49.82$\pm$2.72 51.58$\pm$4.19 86.75$\pm$0.24
HGPSLPool 89.08$\pm$0.83 77.70$\pm$1.01 OOM 58.55$\pm$3.71 49.45$\pm$2.12 50.53$\pm$2.44 OOM